### v1

# A Reanalysis of the FDA’s Benefit-Risk Assessment of Moderna’s mRNA-1273 COVID Vaccine: For 18–25-Year-Old Males, Risks Exceeded Benefits Relative to Hospitalizations

Abstract: The Food and Drug Administration (FDA) conducted a benefit-risk assessment for Moderna’s COVID vaccine mRNA-1273 prior to its full approval (announced 1/31/2022). The FDA focused its assessment on males of ages 18–64 years because the agency’s risk analysis was limited to vaccine-attributable myocarditis/pericarditis (VAM/P) given the excess risk among males. The FDA analysis concluded that vaccine benefits clearly outweighed risks, even for 18–25-year-old males (those at highest VAM/P risk). In this paper, we reanalyze the FDA benefit-risk assessment using information available through the third week of January 2022 and focusing on 18–25-year-old males. We use the FDA’s framework but extend its model by accounting for protection derived from prior COVID infection, finer age-stratification in COVID-hospitalization rates, and incidental hospitalizations (those of patients who test positive for COVID but are being treated for something else). We also use more realistic projections of Omicron-infection rates and more accurate rates of VAM/P. With hospitalizations as the principal endpoint of the analysis (those prevented by vaccination vs. those caused by VAM/P), we find a net harm: vaccine risks outweighed benefits for 18–25-year-old males, except in scenarios projecting implausibly high Omicron-infection prevalence. Comparing the projections of FDA’s “most likely scenario” to ours, we find that the FDA’s benefit-risk ratio, 43.33, is over 60 times more favorable than ours, 0.67. Our assessment suggests that mRNA-1273 vaccination of 18–25-year-old males generated between 16% and 63% more hospitalizations from vaccine-attributable myocarditis/pericarditis alone compared to COVID hospitalizations prevented (over a five-month period of vaccine protection assumed by the FDA). The preceding assessment is based on assumptions drawn from data available before 1/22/22. Moreover, these assumptions are validated by subsequently available data.
1. Introduction
Because vaccines can be given to large populations of healthy persons and can be introduced as mandatory, robust benefit-risk analyses must be conducted to ensure that vaccines have highly favorable benefit-risk profiles [1].
The Food and Drug Administration (FDA) issued emergency-use authorizations for mRNA COVID-19 vaccines from Pfizer (BNT162b2) and Moderna (mRNA-1273) in December 2020 [2]. Once the mRNA COVID-19 vaccines were given to large populations, it became evident that the vaccines could cause myocarditis and pericarditis (hereafter myo/pericarditis), inflammation of the heart and of the membrane surrounding the heart, respectively, with studies indicating the highest risk among males younger than 40 years ([3, 4, 5, 6, 7]).
The FDA has published two benefit-risk analyses of primary series mRNA vaccination, the first for BNT162b2 [8] and the second for mRNA-1273 [9]. The FDA presented its analyses through a number of scenarios involving different modeling assumptions to account for uncertainties associated with pandemic dynamics as well as with the benefits and risks of vaccination. To assess vaccine benefits, the FDA estimated vaccine-preventable COVID-19-cases, hospitalizations, intensive care unit (ICU) admissions, and deaths [8, 9]. To assess vaccine risks, the FDA estimated vaccine-attributable myo/pericarditis (VAM/P) cases, hospitalizations, ICU admissions, and deaths [8, 9]. The FDA concluded that the benefits of the mRNA vaccines outweighed the risks of VAM/P even for males in its highest-risk age-groups: 16–17 years for BNT162b2 [8] and 18–25 years for mRNA-1273 [9].
However, both of the FDA’s benefit-risk assessments omitted important inputs in the analysis of the hospitalizations-prevented benefit of vaccination as well as in the analysis of VAM/P risk. We focus mainly on the FDA’s benefit-risk analysis for Moderna’s mRNA-1273 to simplify our exposition—there are significant differences in modeling assumptions in the FDA’s Pfizer and Moderna assessments. Also, VAM/P risk is substantially greater for mRNA-1273 than for BNT162b2 [3, 5, 6, 10].
For the FDA’s assessment of mRNA-1273, benefits of vaccination were overestimated. First, the FDA did not account for the protection conferred by prior infection; and, at the time of the FDA’s Moderna analysis (mid-January 2022), it is likely that over 70% of 18–29 year-olds had been infected by COVID-19—see Supplement 1 for details. Second, the FDA did not account for incidental COVID-19 hospitalizations. Third, the FDA overestimated benefits for young males by assuming the COVID-hospitalization rate for 18–25 year-olds is equivalent to that for 26–35 year-olds and 36–45 year-olds. Finally, the FDA assumed in its most likely scenario that case and hospitalization rates would remain constant at peak 2021 Omicron levels for 5 months.
By contrast, the FDA underestimated VAM/P risks for mRNA-1273. The agency omitted dose-one mRNA-1273 VAM/P events [9, Section 2.3.1] (but included dose-one events for its BNT162b2 assessment [8, Section 2.3.1]). Moreover, the FDA’s analysis is inconsistent with evidence from the CDC that VAM/P risk for mRNA-1273 was over twice that for BNT162b2 [10, p. 25]. Finally, the FDA relied on internal FDA databases that yielded mRNA-1273 VAM/P rates substantially lower than those found in major international studies [3, 5, 6] as well as a U.S. study [4].
In this study, our primary aim is to extend and improve the FDA’s benefit-risk model for mRNA-1273 vaccination by accounting for prior-infection protection, incidental COVID-19 hospitalizations, and age-specific infection-hospitalization rates, as well as for a more accurate assessment of VAM/P risk. We apply our model using data available to the FDA at the time of its benefit-risk analysis for mRNA-1273, which we take to be the 3rd week of January 2022 for reasons described Section 2.1 below. We also analyze benefits and risks of vaccination for specific subpopulations such as males 18–25 having prior-infection protection and no comorbidities. Our modeling not only addresses the shortcomings in the FDA’s assessment of the mRNA-1273 vaccine but also presents a framework for future benefit-risk assessments of vaccination.
2. The FDA’s mRNA-1273 Benefit-Risk Assessment
In [9], the FDA describes its benefit-risk modeling of Moderna’s mRNA-1273 (Spikevax) for males ages 18 to 64 years. Outcomes are provided via six scenarios reflecting uncertainties in three major model inputs: COVID-19 incidence, vaccine effectiveness against cases (VE) as well as hospitalization (VEH), and vaccine-attributable myo/pericarditis (VAM/P) rate. At the time of the FDA’s mRNA-1273 analysis, the Centers for Disease Control and Prevention (CDC) was recommending a 5-month interval between primary-series completion and a booster dose. Consequently, the FDA evaluated benefits and risks over a 5-month period, which we take to be 1/1/22 – 5/31/22. The FDA assumed that all vaccine effectiveness estimates remained constant over this 5-month evaluation period. The FDA’s major modeling assumptions are summarized in Table 1 below.
Table 1. The FDA’s mRNA-1273 model-input assumptions for 18–25-year-old males [9, Table 1].
Scenario COVID-19 Incidence Vaccine Effectiveness VAM/P Rate Assumptions Per Million 2nd Vaccinations*
$1$ Highest 2021 Incidence, 12/25–12/31 Omicron Dominant; VE = 30%, VEH=72% 128 cases, 110 hospitalizations
2 Average 2021 Incidence Same as Scenario 1 Same as Scenario 1
3 Lowest 2021 Incidence, 6/5   Same as Scenario 1 Same as Scenario 1
4 Same as Scenario 1 Delta Dominant; VE = 80%, VEH=90% Same as Scenario 1
5 Same as Scenario 1 Same as Scenario 1 68 cases, 58 hospitalizations
6 Same as Scenario 1 Same as Scenario 1 241 cases, 207 hospitalizations
*From [9, Table 2].
Using the modeling assumptions described in the first row of the Table 1, with incidence data drawn from [9, Table 3], the FDA concluded for Scenario 1, its “most likely scenario,” that vaccinating one million males ages 18–25 years with two doses would prevent, over 5 months, 82,484 COVID-19 cases, 4,766 COVID-19 hospitalizations, 1,144 ICU admissions, and 51 deaths due to COVID-19, while risking 128 VAM/P cases, 110 VAM/P hospitalizations, 0 ICU admissions, and 0 deaths due to VAM/P [9, Table 2]. However, the FDA’s model overestimates benefits and underestimates risks.
2.1. The Time of the FDA’s mRNA-1273 Analysis
Although the FDA does not define in [9] the time of its benefit-risk analysis for mRNA-1273, a natural choice is the 3rd week of January 2022, which concluded 10 days prior to the FDA’s announcement of the full approval of mRNA-1273 on 1/31/2022 [12]. This choice is natural because it is consistent with the FDA’s time-of-analysis definition for its BNT162b2 assessment: the 2nd week of August 2021 [8, Section 2.1], which concluded nine days prior to the FDA’s announcement of the full approval of BNT162b2 on 8/23/2021 [11]. Consequently, our reanalysis of the FDA’s benefit-risk assessment [9] for the general population of 18–25-year-old males is based on assumptions derived from information available before 1/22/2022. Subsequent information validating our assumptions will often be provided but only via footnotes. Starting in Section 4.2, we no longer restrict ourselves to basing modeling assumptions only on information available before 1/22/2022; in this section, our intent is to demonstrate how our model might be modified to incorporate a comorbidity (obesity) as a risk modifier.
Starting in Section 4.2, where we explore how vaccination’s benefit-risk picture changes with respect to Body Mass Index (BMI), we no longer restrict ourselves to basing modeling assumptions only on information available before 1/22/2022.
2.2. The FDA’s mRNA-1273 Benefits Assessment
The FDA’s computation in [9] of the cases-prevented and hospitalizations-prevented benefits of vaccination is straightforward. For each scenario, CDC data is used to anticipate over the 5-month evaluation period a certain number of COVID cases $C$ and hospitalizations $H$ per million unvaccinated males, and the corresponding evaluation-period benefits per 1 million vaccinations are estimated to be
[…]
where $VE$ and $VEH$ are, respectively, the effectiveness of full mRNA -1273 vaccination against cases and against hospitalization; the FDA assumed that for Omicron $VE=0.30$ and $VEH=0.72$, as noted in Table  1 above.
For Scenario 1, its most likely scenario, and two others, the FDA assumed that peak 2021 Omicron-case rates (late December 2021) would be sustained for 5 months, resulting in an exceptionally high projection of total COVID-19 cases among the unvaccinated; for example, $C=274,947$ for 18–25-year-old males. Similarly, for its most likely scenario (and two others), the FDA assumed that the CDC’s reported late-December 2021 Omicron hospitalization rate for unvaccinated 18–45-year-old males would be sustained for 5 months at approximately 43.54 cases per million daily [9, Table 3] and assumed that the same rate would apply to the younger 18–25 bracket, arriving at $H=6,619$ hospitalizations. Thus, for its most likely scenario (and two others), the FDA predicted that full mRNA-1273 vaccination of 1 million 18–25-year-old males would prevent $274,947\cdot 0.30\approx 82,484$ COVID cases and $6619\cdot 0.72\approx 4,766$ COVID hospitalizations [9, Table 2].
To obtain its projections of ICU admissions prevented by vaccination, the FDA used CDC data to estimate age-based percentages of males hospitalized with COVID going to ICU [9, Table 3]. According to Table 3 of [9], 24% of 18–25-year-old males admitted with COVID require ICU care; thus, the FDA predicted that, under its most likely scenario, fully vaccinating 1 million males in this age group with mRNA-1273 would prevent approximately 1,144 ICU admissions (24% of 4,766)[9, Table 2]. The FDA’s projections of deaths prevented by vaccination utilized the same assumptions as hospitalizations prevented. Specifically, for its most likely scenario, the FDA assumed the CDC’s reported late-December 2021 Omicron death rate for unvaccinated males 18–25 would be sustained for 5 months resulting in a total of about 71 deaths per million and assumed vaccine effectiveness against death to be the same as VEH, estimating $71\cdot 0.72\approx 51$ COVID deaths prevented per million vaccinations.
The following factors contributed to the FDA’s overestimation of vaccine benefits:
$\bullet$ The FDA did not account for the protection conferred by prior infection. The FDA’s not accounting for prior-infection protection in its mRNA-1273 assessment is difficult to justify. A study of SARS-CoV-2–specific immune memory found “Substantial immune memory is generated after COVID-19,” with about 95% of subjects retaining memory at  6 months post infection [15]. A meta-analysis based on studies published December 2020 through August 2021 demonstrated that protection from prior COVID infection “is, at least, equivalent to the protection afforded by complete vaccination of COVID-naïve populations” [16]. The CDC acknowledged in a science brief released 10/29/21 that “The immunity provided by vaccine and prior infection are both high” [17], while another CDC study found that during the Delta wave unvaccinated persons with prior COVID infection had significantly greater protection against infection and hospitalization than fully vaccinated, COVID-naïve persons [18]. Moreover, by the end of December 2021, the FDA acknowledged, “The omicron variant has significantly more mutations than previous SARS-CoV-2 variants, particularly in its S-gene, the gene that encodes the virus’s spike protein”[19], suggesting that prior infection, which entails exposure to all SARS-CoV-2 proteins (such as the nucleoprotein), would likely yield greater protection against Omicron than would mRNA-1273 vaccination, which entails exposure only to the spike protein of the ancestral strain of SARS-CoV-2. The summary of findings of a study published 7/20/21 concludes, “CD4+ T cell responses equally target several SARS-CoV-2 proteins, whereas the CD8+ T cell responses preferentially target the nucleoprotein, highlighting the potential importance of including the nucleoprotein in future vaccines. Taken together, these results suggest that broad and effective immunity may persist long-term in recovered COVID-19 patients[20]. [^1] Consistent with the immunological considerations discussed in this paragraph, a study from Qatar found prior infection provides 52.2% effective protection against symptomatic Omicron infection and 91% effective protection against “severe, critical, or fatal” Omicron infection; whereas, the study’s corresponding effectiveness estimates for 2 doses of mRNA-1273 are 2.2% and 66.3%, respectively [21, Figure 2C, 2D] (study period 12/23/21 – 2/21/22). In Section 7 of Supplement 1, we explain how the meta-analyses and systematic reviews [22, 23, 24, 25] provide convincing evidence that prior infection provided greater protection against Omicron during the evaluation period than did two-dose vaccination among the COVID-naïve.
Additional reasons why the FDA’s not accounting for prior-infection protection is difficult to justify include (i) according to the CDC, 54.9% of those 18–49 years old had been infected by COVID-19 by the end of September 2021 [26]; (ii) a South African study [27] released 12/14/21 suggested overall effectiveness of prior-infection protection against Omicron to be about 47% (see Supplement 1), while data from England [28] released 12/22/21 suggests that prior infection among the unvaccinated provides significant protection from Omicron hospitalization (see Supplement 1); and (iii) the FDA adopted a scenario-based structure for its benefit-risk assessment to account for “the impact of uncertainty associated with pandemic dynamics” [9, Abstract], yet the agency did not consider prior-infection protection in any of its six scenarios.
$\bullet$ The FDA did not account for incidental COVID-19 hospitalizations. At the time of its mRNA-1273 assessment, it was well known that incidental-hospitalization rates were substantial. The day after CDC director R. Walensky [29] stated during a 1/9/22 interview “[U]p to 40 percent of the patients who are coming in with COVID-19 are coming in …with something else,” [^2] We searched unsuccessfully for published CDC data on incidental COVID-hospitalization rates. Figure 2 of a large U.S. study suggests an incidental rate for 0–18 year-olds typically well above 50% during the study period 3/1/2020–1/31/22 and for those 19–64 an incidental rate above 40% for 1/1/2022–1/31/2022 [30]. Canada’s province of Ontario started publishing incidental rates, with rates for January 8–15 exceeding 45% [31]. There is reason to believe incidental rates are especially high for the young—see Supplement 1.
$\bullet$ The FDA overestimated the proportion of projected COVID cases requiring hospitalization. One way to gauge the degree of implausibility of the FDA’s evaluation-period hospitalization projections is to consider the corresponding case-hospitalization rates:
Table 2. Case-hospitalization-rate projections for unvaccinated 18–25-year-old males over the evaluation period 1/1/22 – 5/31/22 for the Omicron-based scenarios of [9, Table 2].*
Scenario Projected Cases Projected Hospitalizations Projected Case-Hospitalization Rate
$1$   274,947   6,619 $\frac{6619}{274,947}\approx 2.41\%$
2 89,017 2,900 $\frac{2900}{89,017}\approx 3.26\%$
3 13,010 882 $\frac{882}{13,010}\approx 6.78\%$
$5$ 274,947 6,619 $\frac{6619}{274,947}\approx 2.41\%$
$6$ 274,947 6,619 $\frac{6619}{274,947}\approx 2.41\%$
*Case and hospitalization projections are obtained from Table 2 of [9] by dividing, respectively, cases-prevented estimates by 0.30 and hospitalizations-prevented estimates by 0.72.
The FDA’s projected case-hospitalization rates are all over four times the Omicron case-hospitalization rate ($0.58\%\approx 56/9,585$) for the unvaccinated of all ages based on Omicron cases in England tested between 1 December 2021 and 14 December 2021, inclusive [28, Table 3]; moreover, all are well over two times the case-hospitalization rate (0.92%) for the unvaccinated of ages 18–29 in Ontario over the period 12/14/20 – 12/12/21 [32, Table A1a]. [^3] The FDA case-hospitalization rates in Table 2 are all well over 5 times the Omicron rate 0.42% for the unvaccinated of all ages reported in Table S10 of [33] and based on data from the period 12/15/21 – 1/17/22. Based on information available at the time of the FDA’s mRNA-1273 assessment, one would expect the COVID-case hospitalization rate for the period 12/14/20 – 12/12/21 would be higher than that for Omicron; e.g., according to [34, p. 4], “The risk of hospital admission from emergency departments with Omicron was approximately one-third of that for Delta (Hazard Ratio 0.33, 95% CI: 0.30 to 0.37)”; see Supplement 1 for more examples. Furthermore, we expect the hospitalization rate for the age-group 18–25 to be lower than that for 18–29 [35]. To some extent, FDA’s exaggerated case-hospitalization rates for the 18–25 age range result from the agency’s assumption that hospital-admission rates were uniform for ages 18–45, contradicting the CDC’s own models [36]. Also, to some extent, the FDA’s exaggerated case-hospitalization rates result from the CDC’s inclusion of incidental COVID hospitalizations in its accounting. Our source for the 0.58% rate mitigated the incidental-hospitalization issue [28, p. 10], while our source [32] for the 0.92% rate did not.
$\bullet$ The FDA assumed in its most likely scenario (as well as Scenarios 5, and 6) that case and hospitalization rates would remain constant at peak 2021 Omicron levels for 5 months. One way to gauge the degree of implausibility of the FDA’s case-rate projection is to consider the corresponding projected number of infections. Throughout the pandemic, the number of COVID infections has always been significantly larger than the corresponding number of reported cases. For the period February 2020 through September 2021, the CDC estimated the number of infections to be four times the corresponding number of reported cases [26]. Moreover, the case-to-infection multiplier likely increased (e.g. owing to increased home testing) during the Omicron-dominant period in the U.S., which started mid-December 2021. If we assume that CDC’s 4:1 infection:case ratio holds over the evaluation period, we find the FDA’s evaluation-period case-count projection of 274,947 per 1 million unvaccinated 18–25-year-old males (Scenarios 1, 5, 6) corresponds to 4 · 274,947 = 1,099,788 infections, well over 1 million. Thus, the FDA’s case-total assumption is essentially equivalent to the assumption that among one million unvaccinated males, there would be, over 5 months, 1,000,000 infections. Having well over one million infections in a group of 1 million 18–25 year-olds over 5 months is implausible. Multiple infections over a five-month period would be unlikely [^4] The study [37] conducted over the period 11/1/21 – 4/30/22 provides evidence that multiple evaluation-period Omicron infections were rare: in the study there were 11,307 positive Omicron tests and none was a second positive Omicron test separated by 90 days from another such positive test. given the high protection from reinfection resulting from recovery from a recent infection, and any reinfections would be offset by individuals who did not become infected at all (e.g., owing to prior-infection protection, limited exposure to infectious persons, or, possibly, genetic factors [38]). [^5] The preprint [38], posted 6/10/21, was later published [39]. Even direct inoculation with SARS-CoV-2 does not necessarily produce infection: in the challenge study [40] of the 36 inoculated, 16 “remained uninfected.” Thus, for 3 of the FDA’s 5 Omicron-based scenarios, including its most likely scenario, the FDA essentially assumed that every unvaccinated male 18–25 would be infected by COVID over the 5-month evaluation period.
2.3. The FDA’s mRNA-1273 Risks Assessment
On the risk side, the FDA relied on its internal databases for the VAM/P case rates appearing in Table 1 above while using the CDC’s Vaccine Safety Datalink (VSD) system to estimate the percentage of VAM/P cases requiring hospitalization [9, Section 2.3.2].
The FDA underestimated VAM/P risks for mRNA-1273:
$\bullet$ The FDA omitted dose-one VAM/P events. For its mRNA-1273 analysis [9], the FDA computed the evaluation-period benefits of two doses of mRNA-1273 without accounting for the VAM/P risk associated with dose 1, potentially biasing the FDA’s analysis in favor of vaccination. [^6] The FDA explains its decision to exclude VAM/P events for dose 1 of mRNA-1273 as follows: “The risk post dose 1 was ignored since CDC has suggested people who have myocarditis after dose 1 take precaution to receive dose 2, and the majority of myocarditis cases were reported post dose 2” [9, Section 2.3.1]. In contrast, for BNT162b2-related VAM/P, the FDA added excess myo/pericarditis cases postdose 1 and postdose 2 to approximate the population-level risk of completing 1 million two-dose vaccination series [8, Section 2.3.1]. Hence, for the BNT162b2 analysis, the FDA did account for VAM/P risk associated with dose 1 but did not assess the benefits of dose 1 alone during the interval starting with dose 1 and ending at the point of full vaccination (7 days after dose 2), potentially biasing the analysis against vaccination. Given that adverse events of vaccination other than VAM/P possibly resulting in hospitalization were ignored by the FDA and that mRNA vaccination might have unknown long-term adverse health impacts, we believe that even in a benefit-risk analysis assessing benefits only postdose 2, the choice to include VAM/P risk associated with dose 1 is appropriate, being consistent with the first-do-no-harm principle, the FDA’s choice in its BNT162b2 assessment, and the FDA’s choice in its benefit-risk analysis for mRNA-1273 described in its BLA (Biologics License Application) Clinical Review Memorandum for mRNA-1273—see Section 2.4. Thus, we include VAM/P hospitalization risk for dose 1 of mRNA-1273 in our main analysis. In Supplement 2, we conduct a sensitivity analysis showing that accounting for dose 1 benefits during the period leading up to the point of full vaccination does not substantially change our benefit-risk conclusions.
$\bullet$ The FDA ignored evidence that VAM/P risk for mRNA-1273 was over twice that for BNT162b2. In a “Review Memorandum Addendum” [41], dated 11/19/21, relating to the emergency-use authorization of a booster dose of mRNA-1273, the FDA describes multiple sources of evidence indicating that VAM/P risk for mRNA-1273 is significantly greater than that for BNT162b2:
[E]vidence was communicated from other regulatory agencies, including agencies from Canada, France, Germany, Norway, and Japan that suggested that there was a higher rate of myocarditis/pericarditis observed in males under 40 years of age following the second dose of the Moderna COVID-19 Vaccine than with the Pfizer-BioNTech COVID-19 Vaccine. For example, Ontario, Canada reported a higher rate of myocarditis in individuals following the Moderna COVID-19 Vaccine than with the Pfizer-BioNTech COVID-19 Vaccine, particularly in males ages 18–24 years ([link provided to Ontario data, later analyzed by Buchan et al. [42]]). …[T]he most recent data presented by the Center for Disease Control and Prevention (CDC) analyzed from the Vaccine Safety Datalink (VSD) indicates a higher rate of myocarditis with the Moderna COVID-19 Vaccine than the Pfizer-BioNTech COVID-19 Vaccine ….
The “most recent data” from the VSD to which the FDA refers was presented at the CDC’s October 2021 Advisory Committee on Immunization Practices (ACIP) meeting [10]. Adjusted-rate ratios for 18–39 year-olds reported in the table on page 25 of [10] suggest that the 2nd-dose VAM/P rate for Moderna’s mRNA-1273 is 2.72 times that of Pfizer’s BNT162b2 and that the corresponding ratio for the 1st dose is 2.63. For 18–39-year-old males, the tabulated data on page 27 of [10] suggests a 2nd-dose rate ratio of 2.26, and, if we assume the same percentage decrease for the 1st dose, then that rate ratio would be 2.19 instead of 2.63. Data from other sources suggests higher rate ratios, especially for dose 2. For example, Buchan et al. found a 2nd dose rate ratio of 5.1 for males aged 18–24 [42, p. 5]. [^7] The specific data (299.5 per million second doses of Moderna vs. 59.2 per million of Pfizer) is the same in the publication [3] and its preprint [42] (available 12/5/2021). A study by Patone et al. [43] of myocarditis cases requiring hospitalization within a 28-day risk interval following a vaccine dose found among males under 40 a rate of 101 excess myocarditis hospitalizations per million 2nd doses of mRNA-1273 and approximately 12 excess myocarditis hospitalizations per 1 million 1st doses, with corresponding rates for BNT162b2 being 12 per million 2nd doses and 3 per million 1st doses [43, p. 4].888 Patone et al. [43] had a study population of over 42 million and a study period from 12/1/20 – 11/15/21. The published version [6] of the preprint [43] incorporates an additional month of data and suggests for males $<40$ a rate of 97 myocarditis hospitalizations per million 2nd doses of mRNA-1273 and a rate of 14 per million 1st doses with corresponding BNT162b2 rates being 11 and 4 [6, Table 4].
Despite all the evidence reviewed above, the FDA chose for its mRNA-1273 risk assessment [9] to rely on “preliminary findings” based on its internal databases suggesting “No clear difference in the rate of myocarditis/pericarditis [mRNA-1273 vs. BNT162b2]” [41, p. 2]. The “preliminary findings” to which the FDA refers have “large uncertainty” owing to “small numbers of observed events” [44, p. 16]. [^9] As noted in [44, p. 7] and in [9, Section 2.3.1], VAM/P-data-collection cutoff points, for the four databases on which the FDA relied for its mRNA-1273 risk assessment were: 8/21/21, 7/10/21, 7/31/21, and 6/30/21, each date being over four months before the time of the assessment. Based on these preliminary findings as well as ignoring dose-1 events, the FDA suggests in its mRNA-1273 assessment that for 18–25-year-old males VAM/P risk for mRNA-1273 is comparable to that for BNT162b2: 128 cases per million 2nd doses of mRNA-1273 [9, Table 4] vs. 131 cases per million full vaccinations of BNT162b2 (among 18–24-year-old males) [8, Table 3], with corresponding hospitalization projections of 110 per million 2nd doses of mRNA-1273 and 131 per million full BNT162b2 vaccinations. Throughout this paper, we assume that VAM/P rates for males 18–24 well approximate those for males 18–25.
Based on data from the preceding two paragraphs, we compute a VAM/P-rate estimate for mRNA-1273 using the FDA’s corresponding estimate for BNT162b2 in [8] and VSD VAM/P rate ratios [10]. FDA data available in October 2021 reports 5 myo/pericarditis events postdose 1 of BNT162b2 and 36 postdose 2 [44, p. 14]. Assuming the same ratio of dose 1 and dose 2 events is reflected in the FDA’s projected 131 VAM/P cases for Pfizer’s BNT162b2 [8, Table 3], we find the corresponding estimate for VAM/P cases resulting from fully mRNA-1273 vaccinating 1 million males 18–25 to be
[…]
where we have applied the VSD-based comparison of VAM/P rates, mRNA-1273 vs. BNT162b2, discussed above. The first summand on the left of Equation (2.3.1) represents a VAM/P incidence of 35 cases per million 1st doses of mRNA-1273 among males 18–25, while the second represents approximately 260 cases per million 2nd doses. Using the FDA’s VAM/P case-hospitalization rate of 86% [9], we find the overall case rate of 295 per million full vaccinations corresponds to a rate of approximately 254 VAM/P hospitalizations per million full vaccinations and 224 per million 2nd doses. Thus, the FDA’s estimate of 110 VAM/P hospitalizations per million 2nd doses of mRNA-1273 (used in Scenarios 1–4) is less than half the estimate of 224 per million 2nd doses derived from the FDA’s BNT162b2 estimate and VSD data.
$\bullet$ Internal FDA databases yielded VAM/P rates for mRNA-1273 substantially lower than those found in important studies.    The FDA’s internal databases not only yielded VAM/P-rate estimates for mRNA-1273 inconsistent with the agency’s BNT162b2 estimates and VSD data but also the estimates are substantially lower than those of major international studies by Buchan et al. [42] and Patone et al.  [43]. Furthermore, a U.S. study by Sharff et al. [45] found, with a wide confidence interval, a myo/pericarditis incidence rate of 537.1 per million second doses of an mRNA-vaccine for males ages 18–24 years. The conclusion of [45] is noteworthy: [^10] The preprint [45] and the later publication [4] share the same conclusion.
We identified additional valid cases of myopericarditis following an mRNA vaccination that would be missed by the VSD’s search algorithm, which depends on select hospital discharge diagnosis codes. The true incidence of myopericarditis is markedly higher than the incidence reported to US advisory committees. The VSD should validate its search algorithm to improve its sensitivity for myopericarditis.
We now compare the FDA’s mRNA-1273 projected VAM/P case and hospitalization rates to those of the other sources discussed above.
{adjustwidth}
-1.55cm
Table 3. Estimated VAM/P Events Per Million 2nd Doses of mRNA-1273 Among Males 18–24 or 18–25 years old.* Based on Data Available Before 1 January 2022
Description
FDA
Scenarios 1–4
(Scenarios 5 & 6)
[9, Table 2]
Buchan et al.**
[42]
Patone et al.
[43, 42]
Sharff et al.
[45]
Data From
FDA & VSD
[8, 44, 10]
Weighted
Average of
Highlighted
Data
Cases
per millon
128
(68 & 241)
  302   269   525   260 301
Hospitalizations***
per million
110
(58 & 207)
  242   231   452   224 250
Number of events
  on which estimate
  is based
21
[44, p. 14]
55
[42, Table 4]
  23
7
[45, Table 1]
  22
Total Events
Columns 3–6
107
* For discussions of the data in columns 3–5, see Section 6 of Supplement 1; in particular, data in column 4 are estimates based on findings from [43] for males under 40, adjusted using an age-based myocarditis-incidence ratio as well as pericarditis data from [42]. Data in column 6 is discussed above in the paragraph of Equation (2.3.1), with 22 being the number of Moderna myo/pericarditis cases contributing to the head-to-head comparison of [10, p. 25].
** All cases of myo/pericarditis contributing to Buchan et al.’s study were judged to meet the Brighton Collaboration’s case definitions (levels 1–3). Applying the Collaboration’s requirements reduced the reported 417 post-mRNA-vaccination study-period cases of myo/pericarditis to 297 [42, p. 5].
*** For its BNT162b2 assessment, the FDA assumed all VAM/P cases were hospitalized [8, Section 2.3.2.2]. For its mRNA-1273 assessment the FDA assumed VAM/P case-hospitalization rates decrease with age, starting with 86% for males 18–25 and ending with 77% for males 36 and over [9, Section 2.3.2.2]. We apply the 86% rate in columns 4–6 (which, for column 5, matches the case hospitalization rate for the 7 males 18–24 having myo/pericarditis after dose 2 of an mRNA COVID vaccine: 6 of the 7 were hospitalized). For column 3, Buchan et al. provide an overall case-hospitalization rate associated with the 2nd dose of an mRNA vaccine (all ages, males & females) of 74.3% [42, Table 1]; we apply a case hospitalization rate of 80% for males 18–24, with 80% being approximately the average of 74.3% and the FDA rate of 86%
$301\approx 302\cdot\frac{55}{107}.+269\cdot\frac{23}{107}+525\cdot\frac{7}{107}+260\cdot\frac{22}{107}$                              $\ddagger$ $250\approx 242\cdot\frac{55}{107}.+231\cdot\frac{23}{107}+452\cdot\frac{7}{107}+224\cdot\frac{22}{107}$
2.4. A Seventh Benefit Risk Scenario for mRNA-1273
The FDA’s BLA Clinical Review Memorandum [46] for mRNA-1273, completed 1/28/2022, describes a benefit-risk scenario that does not match any of the six scenarios appearing in the FDA’s published benefit-risk assessment of mRNA-1273 vaccination [9]. The BLA-Memo Scenario is described as follows in Section 4.7 of [46]:
The results of the modeling assessment predicted that in the population of adults at highest risk of vaccine associated myocarditis/pericarditis (males ages 18–25 years), vaccination of 1 million individuals with two primary series doses of mRNA-1273 vaccine would prevent 76,362 COVID-19 cases, 1,755 hospitalizations, 421 ICU visits and 4 deaths due to COVID-19, but would cause 148 myocarditis/pericarditis cases resulting in 128 hospitalizations, 47 ICU visits, and no deaths.
In Table 4 below, the BLA-Memo Scenario is compared to the FDA’s scenarios of [9, Table 2], in particular to Scenario 1, the “most likely scenario.” There are many noteworthy differences: e.g., dose-one VAM/P risk is included in the BLA-Memo Scenario but not in the scenarios of [9], and, in all scenarios of [9], the FDA predicts 0 ICU admissions owing to VAM/P while the BLA Memo predicts 1 million full vaccinations would cause 47.111111Subsequent analyses provide further evidence that the FDA underestimates VAM/P severity in [9]: A meta-analysis [47] of 23 studies including 854 patients ages 12–20 years found a 92.6% case-hospitalization rate for postvaccination myo/pericarditis with a mean length of stay (LOS) of 2.8 days and a 23.2% ICU-admission rate. Another pooled analysis [48] of 39 studies among 129 adults age 18 years or older (mean age 24) reported a median LOS of 3.0 days, 14% treated for heart failure, 8.1% admitted to the ICU, and 2 deaths (one after mRNA-1273 vaccination). A third meta-analysis [49] included 218 cases with a median LOS of 5.8 days with 3 deaths reported (one following mRNA-1273 vaccination), but provided no ICU-admission rate.
{adjustwidth}
-0.3cm
Table 4. FDA scenarios projecting benefit-risk outcomes per million 18–25-year-old males vaccinated with two primary series doses of Moderna RNA-1273.*
Benefits (Outcomes Prevented) VAM/P Risks
Scenario  Cases Hospitalizations
ICU
Visits
Deaths Cases Hospitalizations ICU Visits Deaths
BLA-Memo 76,362 1,755 421 4 148 128 47 0
1 82,484 4,766 1,144 51 128 110 0 0
2 26,705 2,088 501 48 128 110 0 0
3 3,903 635 152 7 128 110 0 0
4 219,958 5,957 1,430 63 128 110 0 0
5 82,484 4,766 1,144 51 68 58 0 0
6 82,484 4,766 1,144 51 247 207 0 0
* For Scenarios 1–6, the FDA estimated VAM/P risk only for dose 2. Outcomes for the BLA-Memo scenario are from Section 4.7 of [46]. Outcomes for Scenarios 1–6 are from Table 2 of the FDA publication [9] with Scenario 1 being the “most likely scenario.”
Consistent with the BLA Memo Scenario, the FDA’s 1/31/22 press release [12] announcing the full approval of the mRNA-1273 vaccine indicates some mRNA-1273 VAM/P cases do require ICU care. Both the BLA Memo [46, Section 4.7] and the FDA’s benefit-risk assessment of mRNA-1273 [9, Section 2.3.2.2] attribute their contradictory ICU risk estimates to VSD data. The BLA Memo lacks a VSD-data reference, while the FDA author team of [9] cite a source of VSD data [50] whose collection cutoff point is 21 August 2021, over four months before the team completed their mRNA-1273 benefit-risk analysis [9].
The dated VSD data from [50] upon which the FDA [9] relies is scant, reporting only 21 postvaccination myo/pericarditis cases for those 18–29 (males and females, BNT162b2 and mRNA-1273, doses 1 and 2), with 18 of the cases ($\approx 86\%$) requiring hospitalization [50, p. 16]. For comparison, note that Buchan et al.  report a total of 74 myo/pericarditis cases after the 2nd dose of mRNA-1273 for the age range 18–29 based on Ontario data from 6/1/21 to 9/4/21 [42, Table 4]. Moreover, all these Ontario cases occurred during a 7-day risk-window postvaccination while the 21 cases reported in [50, p.  16] occurred in a 21-day risk-window postvaccination. A comparison of Public Health Ontario’s myo/pericarditis data 6/1/21– 9/4/21 [42, Table 4] to VSD data through 10/9/21 [10, pp. 8, 16] shows that, among males and females ages 18–39, the number of myo/pericarditis cases per million second doses of mRNA-1273 detected by Public Health Ontario is 3.76 times the number detected by the CDC’s VSD system—see Appendix 3 of Supplement 2.
Because the BLA Memo’s quantification of the benefits and risks of mRNA-1273 vaccination does not closely relate to any of the scenarios of the FDA’s paper [9], it is unclear what role the analysis of [9] played in the FDA’s regulatory approval process for mRNA-1273. It is noteworthy that the FDA’s projection of no VAM/P-related ICU admissions for mRNA-1273 appears in a paper [9] submitted 5/19/22 well over three months after the appearance of FDA documents [12] and [46], with the former indicating ICU care is needed for some mRNA-1273 VAM/P cases and the latter projecting 47 VAM/P ICU admissions per million mRNA-1273 vaccinations of males 18–25.
The FDA’s published benefit-risk assessment [8] of Pfizer’s BNT162b2 omits ICU-admissions data for VAM/P. Funk et al. state in the last paragraph of the introduction of [8], ”Vaccine attributable myocarditis/pericarditis cases, hospitalizations, intensive care unit (ICU) admissions, and deaths were selected to pair with benefit endpoints for benefit-risk comparison.” However, in Section 2.3.1, Funk et al. indicate there are three risk endpoints, omitting ICU-admissions. There were, in fact, myo/pericarditis-related ICU admissions following BNT162b2 vaccination. The FDA acknowledges this in its 8/23/21 press release [11] announcing the full approval of BNT162b2 (Comirnaty). Moreover, VSD data collected through 8/21/21 and presented at the CDC’s ACIP meeting, August 2021, indicates that for the 12–17 age group there were 19 hospitalizations for post-BNT162b2 myo/pericarditis and 36.8% (7 of the 19) hospitalizations involved an ICU admission [50, p. 16].
2.5. Summary
In [9], the FDA substantially overestimated the hospitalizations-prevented benefit of mRNA-1273 vaccination for 18–25-year-old males by assuming COVID-hospitalization risk for this age group is equivalent to that for the 26–35 and 36–45 groups, by including incidental admissions, by ignoring prior-infection protection, and by assuming in its most likely scenario a COVID-19 case rate so high that the corresponding infection rate for the unvaccinated would be essentially 100% over 5 months. The FDA also substantially underestimated hospitalization risk for VAM/P, with available evidence indicating risk to be well over twice that of its most likely scenario—see Table 3 above. ICU-admission risk for VAM/P was underestimated as well (confirmed by FDA documents [46] and [12]).
3. Reanalysis: Methods
Recall that for its case and hospitalization endpoints, the FDA’s benefit-risk model consists of model inputs $C$ and $H$ (the number of COVID cases and hospitalizations in a test population of one million unvaccinated males) as well as VE and VEH (the fractions of these events prevented by full vaccination). These model inputs are combined to project the benefit of vaccinating the test population, and this projected benefit is compared to the model’s corresponding risk inputs, the number of VAM/P cases and hospitalizations projected for the 2nd mRNA-1273 dose. This comparison, or output of the model, depends on what values are used for each of the model inputs. The FDA performed the comparison under six different scenarios, each corresponding to a particular choice of input values.
Because the FDA relied on CDC data to supply projections of values for $C$ and $H$, the FDA’s model does not explicitly link cases to hospitalizations and similarly doesn’t explicitly link VE and VEH.
We extend and improve the FDA’s benefit-risk model for mRNA-1273 vaccination, focusing on the hospitalizations endpoint: those prevented by vaccination vs. those caused by VAM/P. Our model has the same structure as the FDA’s but has infections (not reported cases) as a key model input and links infections to hospitalizations via another key input, the infection hospitalization rate IHR. Additional inputs account for the fraction of the test population having a prior COVID infection as well as for the protection against infection and hospitalization provided by vaccination, by prior-infection, or by both (hybrid protection). Model inputs are combined to project the hospitalizations-prevented benefit of vaccinating a test population, and this projected benefit is compared to the model’s risk input, the projected number of VAM/P hospitalizations expected to occur in the course of achieving a fully vaccinated test population. In our modeling, ”infection” is any COVID infection, symptomatic or asymptomatic.
Test population. Our model is based on a hypothetical population of 1 million males 18–25, so that our test population consists of those at highest risk of VAM/P (among the groups considered in the FDA’s mRNA-1273 assessment). We assume that the test population is representative, as of 1 January 2022, of all males in this age group in terms of: distribution of ages within the 18–25 range, comorbidities (e.g., obesity, cardiovascular disease, diabetes), and the percentage having had a prior COVID-19 infection. We project the number of hospitalizations for Omicron infection that will occur in this test population over the evaluation period 1/1/22 – 5/31/22 (a) assuming all members remain unvaccinated throughout the period, and (b) assuming all members of the population are vaccinated throughout the period (all fully vaccinated, 14 days past dose 2, on 1/1/22). The difference—the projection of (a) minus that of (b)—is the hospitalizations-prevented benefit of vaccinating the test population. We call a test population consisting of infection naïve unvaccinated persons a baseline test population.
We begin by introducing the inputs to our model.
3.1. Model Inputs
Infection rate $I_{r}$. Rather than base our projections of COVID prevalence on reported cases, we base our modeling on $I_{r}$, the fraction of a baseline test population of unvaccinated infection-naïve males 18–25 projected to become infected (symptomatic and asymptomatic) over the FDA’s 5-month evaluation period.
Infection-hospitalization rate $H_{r}$. We translate infections to projected hospitalizations using an estimated infection-hospitalization rate $H_{r}$ for unvaccinated, infection-naïve males 18–25 years old. $H_{r}$ is the fraction of a population of unvaccinated individuals experiencing a primary COVID infection that will progress to severe disease requiring hospitalization.
Prior-infection fraction $F_{pi}$. This is the fraction of the test population having had a COVID-19 infection before the start of the 5-month evaluation period.
Vaccine effectiveness against infection $E_{v}$. This is the fraction by which the baseline infection rate $I_{r}$ is reduced due to the protection afforded by vaccination. That is, the infection rate for a fully vaccinated test population is $I_{r}(1-E_{v})$.
Prior-infection and hybrid effectiveness against infection, $E_{pi}$ and $E_{h}$. These are the fractions by which the baseline infection rate $I_{r}$ is reduced due, respectively, to the protection afforded by prior infection, or to both vaccination and prior infection (hybrid protection).
Hospitalization risk reduction $HRR_{v}$, $HRR_{pi}$, and $HRR_{h}$. These are the fractions by which the hospitalization rate $H_{r}$ is reduced for populations that are protected respectively by vaccination, prior infection, or both (hybrid protection).
VAM/P hospitalizations $H_{\rm VAM/P}$. The final input to our model is $H_{\rm VAM/P}$, which is the projected number of VAM/P hospitalizations expected to occur in the course of attaining a test population of 1 million fully mRNA-1273 vaccinated 18–25-year-old males. To vaccinate fully 1 million individuals with an mRNA COVID-vaccine, more than 2 million doses must be administered (assuming, e.g., that those who experience myo/pericarditis related to dose 1 do not take dose 2), and the adverse-events associated with each of the over 2 million doses reflect the cost to the population of obtaining the benefits of 1 million full vaccinations.
3.2. Combining Model Inputs to Produce Outputs
Recall that we are considering a test population of 1 million males 18–25 over the evaluation period 1/1/22 – 5/31/22. Let $U$ denote the expected number of Omicron hospitalizations occurring in this test population over the 5-month evaluation period assuming its members remain unvaccinated throughout the period. Let $V$ denote the corresponding number of expected hospitalizations assuming the population is vaccinated throughout the period. The difference $U-V$ is the hospitalizations-prevented benefit of vaccination for the test population, and the outputs $U$ and $V$ can be written in terms of the model inputs defined in the preceding section.
In our modeling, we consider only the possibility of hospitalizations associated with test-population members’ first infections during the evaluation period. We expect second infections over the course of a five-month period to be rare; moreover, infection-hospitalization risk will drop significantly after infection [16, 17, 18].121212 The meta-analyses [22, 23] suggest prior-infection provides greater protection against Omicron hospitalization than full vaccination among the COVID-naïve—see Section 7 of Supplement 1 for details. Our decision to analyze hospitalization risk associated with only first infections during the 5-month evaluation period allows us to assume, as did the FDA, that all effectiveness factors contributing to the analysis remain constant over the evaluation period.
Deriving $U$. To derive $U$, we will assume that the test population remains unvaccinated during the evaluation period. As a first step, suppose that the entire test population is infection naïve. Then, in this baseline population, we expect $(1\ {\rm m})\cdot I_{r}$ infections to develop over 5 months, resulting in $(1\ {\rm m})\cdot I_{r}\cdot H_{r}$ hospitalizations, where $(1\ {\rm m})$ is an abbreviation for (1 million).
Next suppose a fraction $F_{pi}$ of the unvaccinated test population has had a COVID-19 infection prior to the start of the evaluation period. We must account for the protection against infection and hospitalization afforded by prior infection. We accomplish this in two stages, first accounting for protection against infection, then for protection from progression to severe disease for infected individuals. The first stage depends on the model input $E_{pi}$, while the second stage uses $HRR_{pi}$.
So, with fraction $F_{pi}$ of the test population previously infected, $(1-F_{pi})\cdot(1{\rm m})$ are infection-naïve and will respond like the baseline population described above. That is, we expect $(1-F_{pi})\cdot(1{\rm m})\cdot I_{r}$ infections leading to $(1-F_{pi})\cdot(1{\rm m})\cdot I_{r}\cdot H_{r}$ hospitalizations from this group. Among the remaining $F_{pi}\cdot(1{\rm m})$ members of the test population with prior infection, we expect $F_{pi}\cdot(1{\rm m})\cdot I_{r}\cdot(1-E_{pi})$ infections resulting in $F_{pi}\cdot(1{\rm m})\cdot I_{r}\cdot(1-E_{pi})\cdot H_{r}\cdot(1-HRR_{pi})$ hospitalizations. Thus the total number of hospitalizations expected in the unvaccinated test population is
(3.2.1) $U=(1-F_{pi})(1{\rm m})I_{r}H_{r}+F_{pi}(1{\rm m})I_{r}(1-E_{pi})H_{r}(1-HRR_{pi}).$
Deriving $V$. Next we assume that every individual in the test population is fully vaccinated. Deriving the total number of Omicron hospitalizations $V$ for this population is analogous to deriving $U$, but more factors come into play because we must account for the protective effects of both vaccination and prior infection. Since all members of this test population are vaccinated, those with prior infection have hybrid protection. Once again we reason in two stages, first accounting for the protection against infection, then for the protection against hospitalization for infected individuals. The first stage depends on the model inputs $E_{v}$ and $E_{h}$, while the second stage uses $HRR_{v}$ and $HRR_{h}$. Once again, we assume a fraction $F_{pi}$ of the test population has had a prior COVID-19 infection.
Then in the fully vaccinated test population, the expected number of hospitalizations is $(1-F_{pi})\cdot(1{\rm m})\cdot I_{r}\cdot(1-E_{v})\cdot H_{r}\cdot(1-HRR_{v})$ for those without prior infection and $F_{pi}\cdot(1{\rm m})\cdot I_{r}\cdot(1-E_{h})\cdot H_{r}\cdot(1-HRR_{h})$ for those with hybrid protection. So the total number of hospitalizations in the population is
(3.2.2) $V=(1-F_{pi})(1{\rm m})I_{r}(1-E_{v})H_{r}(1-HRR_{v})+F_{pi}(1{\rm m})I_{r}(1-E_{h})H_{r}(1-HRR_{h}).$
Hospitalizations prevented by vaccination. The difference between total Omicron hospitalizations in an unvaccinated test population and total hospitalizations in a fully vaccinated test population is the hospitalizations-prevented benefit of vaccination for the test population:
[…]
We remark that disregarding the protection afforded by prior infection, as the FDA did, is equivalent to assuming in our model that all persons in the test population are infection-naïve, i.e. that $F_{pi}=0$. Notice that if $F_{pi}=0$ in Equation (3.2.3), then its right-hand side reduces to
[…]
Applying the model for benefits-risk assessment of vaccination. For the model to output a meaningful assessment of the benefits and risks of vaccination, we must make good estimates for the model inputs appearing in (3.2.3), as well as for the remaining model input: $H_{\rm VAM/P}$, which by itself describes the risk side of the assessment. Then the hospitalizations-prevented benefit $U-V$ can be compared to $H_{\rm VAM/P}$ to complete the analysis.
3.3. Estimating Values for Model Inputs
Like the FDA, we adopt a scenario-based structure for our benefit-risk assessment to account for uncertainties associated with pandemic dynamics as well as with the effectiveness of vaccine and prior-infection protection. Table 1 of Section 2 above summarizes the model-input assumptions the FDA used for each of the six scenarios of its mRNA-1273 assessment [9]. Observe that the first three scenarios differ only in their COVID-incidence assumptions, with Scenario 1 representing the FDA’s most likely scenario. Our first three scenarios, Scenarios A–C, are reanalyses, respectively, of the FDA’ s Scenarios 1–3 in that our COVID incidence assumptions are derived from the FDA’s corresponding assumptions. Based on information available at the time of the FDA’s mRNA-1273 analysis, we consider the COVID incidence assumption of our Scenario D more likely to be accurate than those of Scenarios A-C; also, for Scenario D our values for other input variables represent our best estimates. Thus, we consider Scenario D to be our most likely scenario. In Scenario E, our final scenario, we assume that prior infection and mRNA-1273 vaccination provide equivalent protection against Omicron, which, for reasons discussed in Section 2.2, we believe to be the minimum reasonable assumption of prior-infection’s protectiveness relative to vaccination’s protectiveness. Model inputs and outputs for our scenarios are described in Table 6 of the next section, while the methods we use to estimate values of model inputs are summarized below. Our estimates are based on data available to the FDA before 1/$2$
Estimating $I_{r}$. As we noted in Section 2.2, the FDA’s COVID-case rate assumption for its most likely scenario, Scenario 1, is so large that it corresponds to the assumption that essentially all unprotected (i.e., unvaccinated) 18–25-year-old males become infected over the 5-month evaluation period. Thus, in Scenario A, our reanalysis of Scenario 1, we too assume all unprotected persons in the test population become infected. However, in our modeling, the unprotected are the infection-naïve unvaccinated. Assuming all such persons become infected over 5 months is equivalent to setting $I_{r}=1$, which is our Scenario-A value for $I_{r}$. As discussed in Section 1 of Supplement 1, for our Scenarios B and C, values of $I_{r}$ are determined by the FDA’s COVID-incidence assumptions for its Scenarios 2 and 3, respectively, while for our Scenarios D and E, $I_{r}$ is derived from the assumption that over the evaluation period the increase in COVID-infection level among 18–25-year-old males is twice that of the general population during the second COVID wave in the U.S. (10/1/20–2/28/21).
Estimating $H_{r}$. Our estimate for $H_{r}$ is derived from an IHR model [51] developed via a seroprevalence-based meta-analysis anchored to the ancestral strain of SARS-CoV-2 and available as a preprint in October 2021 [52]. Applying the model, we obtain a “with COVID” IHR of 0.37% for the infection-naïve unvaccinated of ages 18–25 and reduce this rate assuming 40% of hospitalizations attributed to COVID are incidental, arriving at an IHR of $0.22\%\approx 0.6\cdot 0.37\%$ (for infection-naïve unvaccinated 18–25 year-olds). The 0.22% rate for the general population of 18–25 year-olds in the U.S. (males and females) is consistent with U.S. military COVID-19 hospitalization data from the period 1 February 2020 through 4 January 2021 [53]. See Supplement 1 for further details concerning our derivation of the IHR value $0.22\%$.
In its benefit-risk assessments [8] and [9] for, respectively, Pfizer’s and Moderna’s mRNA vaccines, the FDA relied on sex-specific CDC hospitalizations data. The case and hospitalization rates per 100,000 persons in Table 2 of [8] suggests that for the 18–24 age range, females are 2.53 times as likely as males to be hospitalized with COVID-19 and 1.21 times as likely to have a case of COVID-19.131313Moreover, these relationships are consistent with those of [54, Table 1] based on data from 3/1/20 to 3/1/22. In Section 2 of Supplement 1, we use hospitalization and case data from Table 2 of [8] to estimate the IHR for unvaccinated, infection-naïve males 18–25 to be $0.14\%$ (and for females 18–25 to be $0.29\%$).
Estimating $F_{pi}$. For our Scenarios A–C we use the value $F_{pi}=0.62$, which is the average of (a) the CDC’s estimate, mentioned earlier, that 54.9% of those 18–49 years old had been infected by COVID-19 before 10/1/21 [26] with (b) an estimate derived in Supplement 1 that 69% of unvaccinated male 18–25 year-olds had been infected by 1 January 2022. For Scenarios D and E, we use $F_{pi}=0.69$. These values of $F_{pi}$ may be underestimates because they rely on the CDC’s infected rate of 54.9% by 10/1/21 for the general population of 18–49 year-olds, and we would expect a somewhat higher infected rate among the unvaccinated.
Estimating $E_{v}$, $HRR_{v}$. Recall that the FDA used VE = 0.30 and VEH = 0.72 for all its Omicron-based scenarios [9, Table 1]. For reasons discussed in Supplement 1, for all our scenarios, we use $E_{v}=0.3$ and $HRR_{v}=0.67$. Hence, we are assuming a VEH of $1-(1-0.30)(1-0.67)\approx 0.77$, which exceeds the VEH used by FDA. We remark that the FDA’s choices of $VEH=0.72$ and $E_{v}=0.30$ are not consistent with the data and methods of the UK briefing [34] from which the FDA obtained its VEH estimate.
Estimating $E_{pi}$ and $E_{h}$. Our values $E_{pi}=0.47$ and $E_{h}=0.57$ are derived (see Supplement 1) from a December 2021 study [27], based on data from South Africa. [^14] As discussed in Supplement 1, our effectiveness estimates involving prior infection are consistent with those of meta-analyses [22] and [23].
[…]
{adjustwidth}
-1cm
Table 5. Summary of Model Inputs
Variables for $U$ and $V$ Range of Values Scenario A Values
$I_{r}$: Baseline-infection rate—fraction of an unvaccinated, infection-naïve test-population of 18–25-year-old males expected to contract an infection over the 5-month evaluation period. Infections need not accumulate at a constant rate $0\leq I_{r}\leq 1$ $1$
$H_{r}$: Baseline infection-hospitalization rate for infection-naïve unvaccinated males 18–25 $0\leq H_{r}\leq 0.01$ $0.0014$
$F_{pi}$: Fraction of a test population of 1 million unvaccinated males 18–25 having had a COVID-19 infection before the start of the 5-month evaluation period $0\leq F_{pi}\leq 1$ 0.62
$E_{pi}$: Effectiveness of prior-infection protection against COVID-19 infection $0\leq E_{pi}\leq 1$ 0.47
$E_{v}$: Effectiveness of vaccination against COVID-19 infection $0\leq E_{v}\leq 1$ $0.3$
$E_{h}$: Effectiveness of hybrid protection against COVID-19 infection $0\leq E_{h}\leq 1$ 0.57
$HRR_{pi}$: Hospitalization-risk reduction for those reinfected and unvaccinated $0\leq HRR_{pi}\leq 1$ $0.79$
$HRR_{v}$: Hospitalization-risk reduction for those fully vaccinated and experiencing their first COVID-19 infection $0\leq HRR_{v}\leq 1$ $0.67$
$HRR_{h}$: Hospitalization-risk reduction for those reinfected and fully vaccinated $0\leq HRR_{h}\leq 1$ $0.88$
$H_{\rm VAM/P}$: Projected number of VAM/P hospitalizations occurring in the course of attaining a test population of 1 million fully mRNA-1273 vaccinated 18–25-year-old males $H_{\rm VAM/P}\approx 268$ $268$
*Remarks. (i) All measures of effectiveness of protection against infection ($E_{pi}$, $E_{v}$, and $E_{h}$) are relative to infection-naïve, unvaccinated persons. (ii) All hospitalization-risk reductions ($HRR_{pi}$, $HRR_{v}$, and $HRR_{h}$) are relative to unvaccinated persons experiencing primary infections.
4. Reanalysis: Results
In Table 6 below, we compare the hospitalizations-prevented benefit of mRNA-1273 vaccination to corresponding VAM/P-hospitalizations risk through our Scenarios A–E, with Scenarios A–C being reanalyses of the corresponding FDA Scenarios 1–3.
{adjustwidth}
-1.2cm
Table 6. Scenarios projecting the hospitalizations-prevented benefit of mRNA-1273 vaccination of a hypothetical test population of 1 million males ages 18–25 compared to the estimated corresponding VAM/P hospitalization risk $H_{\rm VAM/P}$. For the FDA’s Scenarios 1–3, the values of $U$ and $V$ are determined by the FDA’s model described in Section 2.2, while, for our Scenarios A-E, the values of $U$ and $V$ determined .by our model described in Section 3. A green background for a scenario indicates benefit exceeds risk.
Scenario Description Assumptions* $U$ $V$ Hospitalizations Prevented Per Million Vaccinations $U-V$ $H_{\rm VAM/P}$ Benefit Risk Ratio
FDA
Scenario 1
Highest 2021 COVID in- cidence assumed. FDA’s “most likely scenario”
$U=6619$
$V=6619(1-0.72)\approx 1853$
6619 1853 4766 110 (2nd-dose only) 43.33
A Reanalysis of FDA’s Scenario 1 $I_{r}=1,H_{r}=0.0014,\\ F_{pi}=0.62,E_{pi}=0.47,\\ E_{v}=0.30,E_{h}=0.57,\\ HRR_{pi}=0.79,HRR_{v}=0.67,\\ HRR_{h}=0.88$ 629 168 461 268 1.72
FDA
Scenario 2
Average 2021 COVID incidence assumed
$U=2900$
$V=2900(1-0.72)\approx 812$
2900 812 2088 110 (2nd-dose only) 18.98
B Reanalysis of FDA’s Scenario 2 $I_{r}=0.356,H_{r}=0.0014,\\ F_{pi}=0.62,E_{pi}=0.47,\\ E_{v}=0.30,E_{h}=0.57,\\ HRR_{pi}=0.79,HRR_{v}=0.67,\\ HRR_{h}=0.88$ 224 60 164 268 0.61
FDA
Scenario 3
  Lowest incidence of pandemic assumed  (6/5/21) $U=882$ $V=882(1-0.72)\approx 247$ 882 247 635 110 (2nd-dose only) 5.77
C Reanalysis of FDA’s Scenario 3 $I_{r}=0.052,H_{r}=0.0014,\\ F_{pi}=0.62,E_{pi}=0.47,\\ E_{v}=0.30,E_{h}=0.57,\\ HRR_{pi}=0.79,HRR_{v}=0.67,\\ HRR_{h}=0.88$ 33 9 24 268 0.09
D COVID incidence   twice that for the 2nd COVID wave in the U.S. Our “most likely scenario” $I_{r}=0.456,H_{r}=0.0014,\\ F_{pi}=0.69,E_{pi}=0.47,\\ E_{v}=0.30,E_{h}=0.57,\\ HRR_{pi}=0.79,HRR_{v}=0.67,\\ HRR_{h}=0.88$ 247 68 179 268 0.67
E Equivalent protection for prior infected and COVID-naïve vaccinated; incidence as in Scenario D. $I_{r}=0.456,H_{r}=0.0014,\\ F_{pi}=0.69,E_{pi}=0.30,\\ E_{v}=0.30,E_{h}=0.57,\\ HRR_{pi}=0.67,HRR_{v}=0.67,\\ HRR_{h}=0.88$ 300 68 232 268 0.87
* For discussions of how “Assumptions” were derived, see Section 3.3 as well as Supplement 1.
To illustrate the application of our model, we carry out our Scenario-A computation. Using Scenario A assumptions for variable values and our definition (3.2.1) of the function $U$, we obtain
$U=\left(1-0.62\right)\cdot\left(1{\rm m}\right)\cdot 1\cdot 0.0014+0.62\cdot\left(1{\rm m}\right)\cdot 1\cdot\left(1-0.47\right)\cdot 0.0014\cdot\left(1-0.79\right)\approx 629,\$ and, from our definition (3.2.2) of the function $V$,
$V=\left(1-0.62\right)\cdot\left(1{\rm m}\right)\cdot 1\cdot\left(1-0.30\right)\cdot 0.0014\cdot\left(1-0.67\right)+0.62\cdot 1\cdot\left(1-0.57\right)\cdot 0.0014\cdot\left(1-0.88\right)\approx 168.$
[…]
Observe that our Scenario-D projection of 179 COVID-hospital admissions prevented by vaccination corresponds to $43\approx 0.24\cdot 179$ ICU admissions prevented [9, Table 3], which should be compared to the 47 ICU admissions for VAM/P projected by the BLA-Memo Scenario and 0 such admissions projected in the scenarios of [9].
In Supplement 2, we carry out sensitivity analyses exploring how the benefit-risk picture for mRNA-1273 vaccination changes with modified assumptions for key variables. We also show that in Scenarios B–E, when hospitalizations prevented by vaccination include those prevented by dose 1 during the period preceding the point of full vaccination, vaccine risks still outweigh benefits in terms of the hospitalizations endpoint.
4.1. Subpopulation Analyses Based on Prior-Infection Status
The FDA acknowledges, “The health condition of individuals may remarkably impact the B-R profile and its evaluation” [9, Section 4]; however, the authors of [9] do not explore this remarkable impact, explaining “[W]e did not investigate the benefits and risks for specific subpopulations, such as those with comorbidities, due to limited information for these populations.” Comorbidities such as obesity may triple COVID-hospitalization risk [55]. In this subsection, we consider subpopulations for which there was sufficient information before January 2022 to conduct conclusive benefit-risk assessments. We consider, in particular, subpopulations having had a COVID infection prior to the evaluation period 1/1/22 – 5/31/22.
Prior infection not only decreases the hospitalization-prevented benefit of vaccination but also increases the risk of a severe adverse event after vaccination: see, e.g., [56, 57], with [56] reporting prior infection is “associated with an increased risk of severe side effects leading to hospital care (1.56 [risk ratio] (95%CI: 1.14–2.12)).” Such “severe side effects” possibly linked to mRNA COVID vaccination and potentially requiring hospitalization include anaphylaxis [58], Multisystem Inflammatory Syndrome [59, 60], Guillain Barré Syndrome [61], posterior orthostatic tachycardia syndrome (POTS) [62], retinal vascular occlusion [63], and cerebral venous sinus thrombosis (CVST) [64]. Moreover, there is evidence that prior infection increases the risk of vaccine-associated myocarditis [43]—see Appendix 4 of our Supplement 2 for details. [^19] Survey results reported in [65] also suggest that prior infection may increase the risk of VAM/P; e.g., in Table 4, we find that 2.53% of those having a “laboratory-confirmed SARS-CoV-2 infection” before receiving dose 1 of mRNA-1273 reported “chest tightness/pain” during the 7-day period following the dose while only 0.54% of those not having a confirmed infection before dose 1 reported these symptoms; corresponding percentages for dose 2 (Table 5) are 1.78% and 0.98%.
Prior infection’s reducing the hospitalizations-prevented benefit of vaccination while increasing postvaccination adverse-event risk creates a compelling case for conducting a benefit-risk assessment for the subpopulation of unvaccinated 18–25-year-old males having prior-infection protection.
Consider the subpopulation of 18–25-year-old unvaccinated males having a COVID-infection prior to the evaluation period 1/1/22 – 5/31/22. Using our model, we can compute the expected number of COVID hospitalizations for this subpopulation in various scenarios simply by setting $F_{pi}=1$ and leaving all other variables unchanged. For example, with $F_{pi}=1$ in Scenario A, our model predicts $U\approx 155.82$ hospitalizations per million and $V\approx 72.24$, yielding a hospitalizations-prevented benefit of 85.8 per million vaccinated, far below VAM/P hospitalization risk.
The analysis of the preceding paragraph may be reinterpreted in terms of evaluation-period hospitalization risk for a typical unvaccinated 18–25-year-old male, COVID-infected prior to the evaluation period. For ease of exposition, we will call this male Bob. [^20] If we were also considering COVID-hospitalization risk for a typical 18–25-year-old female, we would call her Alice, choosing “Alice” and “Bob” because these names are customarily used for “person A” and “person B” in the field of cryptography. If Bob chooses not to vaccinate, our model estimates the probability that Bob will be hospitalized for COVID during the evaluation period to be 155.82 in 1 million (assuming $I_{r}=1$). Bob’s total hospitalization risk associated with the choice to vaccinate is essentially the sum of his probability of VAM/P hospitalization and the probability of his having a breakthrough-case COVID hospitalization during the evaluation period (assuming postvaccination VAM/P hospitalization is independent of postvaccination COVID-hospitalization).212121 Technically speaking, because there is a small chance—on the order of 0.02 in 1 million—that Bob might experience both a VAM/P hospitalization and a breakthrough-COVID hospitalization, Bob’s total hospitalization risk from vaccination is the sum of his VAM/P-related risk and his COVID-related risk minus the product of these two risks (with the product being the small chance that Bob experiences both a VAM/P hospitalization and a COVID hospitalization by our assumption of independence of these events).
While there is some evidence that VAM/P risk increases with prior infection, we assume only that VAM/P risk does not decrease. Suppose Bob decides to take dose 1 and at the same time decides to take dose 2 if VAM/P does not occur after dose 1; then, if $r_{1}$, $r_{2}$ are, respectively, the VAM/P-incidence rates for doses 1 and 2 for males 18–25, then the probability Bob will experience VAM/P is well approximated by $r_{1}+(1-r_{1})r_{2}.$ Because $r_{1}$ is quite small, the preceding probability is essentially $r_{1}+r_{2}$, which explains why, even on the individual level, the FDA should consider the sum of risks, postdose 1 and postdose 2 in assessing the VAM/P risk of full vaccination (as the agency did in its population-level analyses of BNT162b2 [8, Section 2.3.1]) and of mRNA-1273 in its BLA-Memo [46, Section 4.7]). Therefore 268/1,000,000 approximates the probability that Bob experiences hospitalization owing to VAM/P during the course of mRNA-1273 vaccination. The probability that he experiences COVID hospitalization owing to a breakthrough infection is 72.24 in 1 million, according to our model. Thus, his total hospitalization risk from vaccination is approximately 340.24 in 1 million. This is under the Scenario-A assumption that $I_{r}=1$, which we interpret to mean that evaluation-period infection is certain for an infection-naïve male 18–25, and Bob’s probability of infection is $I_{r}(1-E_{pi})=0.53$. The following Table shows how Bob’s hospitalization risk varies as his risk of infection varies.
{adjustwidth}
-1.2cm
Table 7. Hospitalization-risk ratios for a typical male 18–25 with prior-infection protection: risk of hospitalization after choosing mRNA-1273 vaccination (VAM/P or severe COVID-breakthrough infection*) divided by risk of COVID-hospitalization after choosing not to vaccinate.
Probability
of Infection $I_{r}$**
Probability of
COVID-19 Hospitalization
Without mRNA-1273 Vaccination
Probability of
COVID-19 Hospitalization
After mRNA-1273 Vaccination
Hospitalization Risk Ratio
vaccinated/unvaccinated, with
VAM/P Risk 268 in 1 million
$1$ 155.82 in 1 million 72.24 in 1 million $2.18\approx\frac{268+72.24}{155.82}$
$0.5$ 77.91 in 1 million 36.12 in 1 million $3.90\approx\frac{268+36.12}{77.91}$
$0.25$ 38.96 in 1 million 18.06 in 1 million $7.34\approx\frac{268+18.06}{38.955}$
*COVID hospitalization risk is considered over the 5-month evaluation period
**$I_{r}(1-E_{pi})$ is the probability the “typical male” (18–25 with prior-infection protection) will become COVID infected over the evaluation period if unvaccinated; e.g., 0.53 if $I_{r}=1$
In Section 2.3 of Supplement 2, based on seroprevalence and modeling information, we estimate that $I_{r}\approx 0.5$ (and is likely $<0.5$), that is, roughly 1/2 of unvaccinated, infection-naïve 18–25-year-old males in the U.S. became infected over the evaluation period (and Bob’s infection risk is $0.2651/2(1-E_{pi})=$ 0.265 because of his prior infection). Thus, according to Table 7, our modeling suggests that for our typical unvaccinated 18–25-year-old male Bob with prior-infection, the choice to vaccinate creates a hospitalization risk approximately 4 times that associated with the choice not to vaccinate. Moreover, if we account for the increased likelihood of postvaccination adverse-event hospitalizations for those with prior infection, then the risk ratios of Table 7 grow even larger. Also note that for every value of $I_{r}$ in Table 7, hospitalization risk owing to VAM/P alone exceeds COVID-hospitalization risk.
Applying the information in Table 7 is difficult because a “typical male” has an “average comorbidities level” and the meaning of this is unclear. However, if our typical unvaccinated male Bob with prior-infection protection also has no comorbidities, then his hospitalization-risk ratios should be larger than those of Table 7. (There is no evidence that being free of comorbidities decreases VAM/P risk; e.g., according to a 7/16/21 report of the chief science advisor of Canada “Patients presenting with this complication [vaccine-associated myo/pericarditis] typically have been healthy individuals.” [66]. [^22] See also [67, p. 383].)
Now assume that Bob is an unvaccinated 18–25-year-old male with prior infection without comorbidities. To establish conclusively that Bob’s hospitalization risk is greater with vaccination, we consider two bounding analyses, each overestimating the benefits of vaccination and underestimating the adverse-event risks associated with vaccination.
{adjustwidth}
0.3in0.3in Scenario I: To underestimate hospitalization risk associated with vaccination, we assume that vaccination is 100% effective in preventing breakthrough-infection hospitalization, and we obtain a lower bound on Bob’s VAM/P-hospitalization risk, by ignoring dose 1 VAM/P hospitalizations and using our estimate of 242 VAM/P hospitalizations per million second doses of mRNA-1273 among males 18–25 based on Ontario data (as described in Table 3). Thus, ignoring his first-dose VAM/P hospitalization risk, his hospitalization risk for postvaccination adverse events other than myo/pericarditis, and his likely increased risk for adverse-event hospitalization owing to his prior infection, we obtain an underestimate of Bob’s postvaccination hospitalization risk of 242 in 1 million (based on VAM/P alone).
To overestimate Bob’s COVID-hospitalization risk assuming he remains unvaccinated, we assume that evaluation-period infection is certain for an infection-naïve unvaccinated male ($I_{r}=1$) and that Bob’s being free of comorbidities does not decrease his COVID hospitalization risk which is 155.82 in 1 million according to Table 7.
We conclude that Bob is more likely—at least $1.55\approx\frac{242}{155.82}$ times as likely—to be hospitalized for VAM/P after choosing to vaccinate than to be hospitalized for COVID while remaining unvaccinated.
{adjustwidth}
0.3in0.3in Scenario II: If we assume no COVID-hospitalization risk-reduction from Bob’s being free of comorbidities, assume his VAM/P-hospitalization risk is only 110 in 1 million (the FDA’s assumption in its most likely scenario of [9]), ignore hospitalization risks for other postvaccination adverse events (likely increased owing to Bob’s prior infection), and assume every infection-naïve unvaccinated male becomes infected over the evaluation period, then we still conclude that Bob is more likely—at least $1.17\approx\frac{110+72.24}{155.82}$ times as likely—to be hospitalized for VAM/P or a breakthrough infection after choosing to vaccinate than to be hospitalized for COVID while remaining unvaccinated. Note that if we use the VAM/P hospitalization risk estimate of the FDA’s BLA-Memo Scenario the ratio 1.17 rises to $1.29\approx\frac{128+72.24}{155.82}$.
To further elucidate the important role that prior infection plays in a careful benefit-risk assessment of mRNA-1273 vaccination, we note that for an 18–25-year-old male without prior COVID infection, hospitalization-risk ratios become favorable for vaccination:
{adjustwidth}
-1.4cm0cm
Table 8. Hospitalization-risk ratios for a typical male 18—25 without prior-infection protection: risk of hospitalization after choosing mRNA-1273 vaccination (VAM/P or severe COVID-breakthrough infection*) divided by risk of COVID hospitalization after choosing not to vaccinate.
Probability
of Infection $I_{r}$
Probability of
COVID-19 Hospitalization
Without mRNA-1273 Vaccination
Probability of
COVID-19 Hospitalization
After mRNA-1273 Vaccination
Hospitalization Risk Ratio
vaccinated/unvaccinated, with
VAM/P Risk 268 in 1 million
$1$ 1400 in 1 million 323.4 in 1 million $0.42\approx\frac{268+323.4}{1400}$
$0.5$ 700 in 1 million 161.7 in 1 million $0.61\approx\frac{268+161.7}{700}$
$0.25$ 350 in 1 million 80.85 in 1 million $1.00\approx\frac{268+80.85}{350}$
*COVID hospitalization risk is considered over a 5-month evaluation period
4.2. Subpopulation Analyses Based on Prior-Infection Status and BMI
For the remainder of the paper, we no longer confine to footnotes discussions and citations of data unavailable at the time of the FDA’s benefit-risk analysis of mRNA-1273.
We illustrate how comorbidities may impact a benefit-risk assessment, focusing on obesity because evidence suggests that Body Mass Index (BMI) is the single most potent modifier of COVID-19 risk for those under 60 years old [68, 69]. According to the CDC [55], “Having obesity [BMI $\geq 30$] may triple the risk of hospitalization due to a COVID-19 infection.” This is consistent with a finding that hospitalization risk for the obese is 3.2 times that for the non-obese for an adolescent study population (ages 12–17) [70, Table IV]. Gao et al. [71] provide additional support for this CDC estimate based on a cohort study of 6,910,695 residents of England (data from the period 1/24/20 – 4/30/20). Gao et al. found that for those 20–39 years old COVID hospitalization risk is minimized when BMI $\approx 23$ and that for $23\leq\text{BMI}\leq 44$ risk is multiplied by a factor of approximately 1.09 (1.08, 1.10) for each unit increase in BMI. Gao et al. “found no difference in effect estimates by sex” [71, Figure 2]. Applying Gao’s model (see Appendix 5 of Supplement 2), we obtain the following:
Table 9. Estimated COVID hospitalization hazard ratio (HR) for 20–39 year-olds relative to those with BMI $\approx 23$ based on a model by Gao et al. [71]
BMI Range COVID Hospitalization HR
23–24 ref.
25–29 1.49
30–34 2.29
35–39 3.52
40–44 5.41
Assuming that for 18–25-year-old males obesity triples COVID-hospitalization risk and that 20.5% of this population is obese [72], we find not being obese drops baseline COVID hospitalization risk $H_{r}=0.0014$ to approximately $0.71\cdot H_{r}$ (see Appendix 5 of Supplement 2). We assume this same factor determines the drop in hospitalization risk for unvaccinated males 18–25 with prior infection (IHR $=0.71\cdot H_{r}\cdot(1-E_{pi})(1-HRR_{pi})$).
We assume that not being obese for an otherwise typical vaccinated 18–25-year-old male decreases IHR by a greater amount: IHR $=0.36\cdot H_{r}(1-E_{v})(1-HRR_{v})$ based on case-hospitalization rates computed from the raw data of Supplementary Table A2 of [73].
We provide below a COVID hospitalization benefit-risk analysis derived from Gao et al.’s BMI-based model [71] described above. We assume that the BMI range 23–24 for 18–25-year-old males provides minimum COVID-hospitalization risk, which we estimate by multiplying the hospitalization risks appearing in column 2 of Tables 7 and 8 by the factor 0.71 and those in column 3 by the factor of 0.36. For a BMI-based hazard ratio HR from Table 9, we compute the corresponding effect on the hospitalization risks in column 2 of Tables 7 and 8 via the factor $\text{HR}\cdot 0.71$ and for those in column 3 via the factor $\text{HR}\cdot 0.36$.
Table 10. Hospitalization-risk ratios (risk if vaccinated (VAM/P & breakthrough)risk if $\left(\frac{\text{risk if vaccinated (VAM/P \& breakthrough)}}{\text{risk if unvaccinated}}\right)$ based on BMI for males 18—25. Ratios displayed as pairs: “with prior-infection protection, without prior-infection projection.”
Probability
of Infection $I_{r}$*
BMI 23–24 BMI 25–29 BMI 30–34 BMI 35–39 BMI 40–44
$1$ 2.66, 0.39 1.86, 0.30 1.29 , 0.23 0.92, 0.19 0.68, 0.17
$0.5$ 5.08, 0.66 3.49, 0.48 ** 2.35, 0.35 1.61, 0.27 1.13, 0.22
$0.25$ 9.92 ,1.20 6.74 , 0.84 4.47 , 0.59 2.99, 0.42 2,03, 0.32
*With prior-infection protection, $I_{r}(1-E_{pi})$ is the probability of infection over the evaluation period.
**Sample computation: $\frac{268+36.12\cdot 1.49\cdot 0.36}{77.91\cdot 1.49\cdot 0.71},\frac{268+161.7\cdot 1.49\cdot 0.36}{700\cdot 1.49\cdot 0.71}\approx 3.49,0.48$
Consistent with Table 8, the preceding table suggests that for 18–25-year-old males without prior-infection protection, vaccination reduces hospitalization risk for all BMI ranges except for the ideal range (BMI 23–24) in a low-infection-risk scenario. Similarly, consistent with Table 7, we see that for 18–25-year-old males with prior-infection protection, vaccination increases hospitalization risk except for the those with very high BMI (35–44) in a high-infection-risk scenario. In a medium infection-risk scenario ($I_{r}=0.5$), our modeling suggests that for the non-obese (BMI $<30$) having prior-infection protection the choice to vaccinate creates a hospitalization risk 3.49 to 5.08 times that associated with the choice not to vaccinate.
5. Validating the Model
In this section, we discuss our model-validation work of Supplement 2, where we test our model as well as the FDA’s using COVID-hospitalizations data from Ontario over the evaluation period 1/1/22 – 5/31/22 and show that our model produces projections of the hospitalizations-prevented benefit of vaccination that are much more accurate than those of the FDA [9]. We test our model in two ways: (i) showing that it fits Ontario COVID-hospitalization data with inputs at plausible levels for Ontario over the evaluation period (see Section 8 of Supplement 2), and (ii) using Ontario COVID-hospitalization data to predict corresponding data for the U.S. and then comparing those predictions to ours (see Section 7 of Supplement 2). We validate our model using Ontario data for many reasons; e.g., unlike the CDC, Public Health Ontario chose to publish incidental COVID-hospitalization rates (starting 1/8/2022). We discuss other reasons in Section 3 of Supplement 2.
Because Public Health Ontario does not provide sex-specific COVID-hospitalizations data, to test our model using its data, we compute model outputs assuming no sex-related differences in COVID infection and hospitalization risk, so that $H_{r}=0.0022$, and, for Scenario D, $F_{pi}=0.76$ (see Supplement 1, Sections 2 and 3). If $H_{r}$ and $F_{pi}$ are increased to $0.0022$ and $0.76$ in Scenario D, leaving all other variables unchanged, then 231 is the hospitalizations-prevented benefit per million vaccinations.
Based on COVID hospitalization counts provided by Public Health Ontario as well as corresponding incidental hospitalization rates, we estimate that vaccinating one million 18–25-year-old Ontarians prevented 152 COVID hospitalizations in Ontario over the evaluation period—see Table S2.5. Because seroprevalence studies suggest that over the evaluation period, the COVID-infection rate in Ontario exceeded that in the U.S., we would anticipate a hospitalizations-prevented benefit of less than 152 per million in the U.S. provided we assume that 18–25 year-olds in the U.S. and in Ontario have equivalent comorbidity profiles. Making adjustments for better health of Ontarians, ignoring the greater infection-level rise in Ontario, and assuming all the vaccinated were 14 days postdose 2 on 1/1/22, we estimate that at most 214 evaluation-period hospitalizations were prevented by one million COVID vaccinations of 18–25 year-olds in the U.S. [Supplement 2, Table S2.7]. [^23] In addition to the greater infection-level rise in Ontario, other reasons why we view the rate of 214 likely to be an upper bound on the actual rate include the following: our ignoring evidence that incidental COVID hospitalization rates are higher among the young, ignoring evidence that hospitalization rates are higher among females, and applying a better-health adjustment such that “the better health of Ontarians [relative to Americans] is more likely overstated than understated”—see Section 4 of Supplement 2. Observe that our most likely scenario’s predicted benefit of 231 hospitalizations prevented per million vaccinations (assuming equivalent IHR for males and females) is only about 1.08 times larger than the estimated actual benefit of 214 [Supplement 2, Table S2.7], while FDA’s most likely scenario’s predicted benefit of 4,766 hospitalizations prevented per million vaccinations is over 22 times larger. Moreover, had the FDA assumed equivalent infection and hospitalization rates among males and females, then its corresponding benefit projection likely would have been over 6,000 hospitalizations prevented per million vaccinations.
Ontario data also supports the bounding analysis of Scenario I from Section 4.1 in which 155.82 in 1 million is used as an upper bound on the evaluation-period COVID-hospitalization risk for a typical unvaccinated male Bob having prior COVID infection and no comorbidities. Ontario data suggests that 151 in 1 million is an overestimate of Bob’s evaluation-period COVID hospitalization risk—see Section 2.3 of Supplement 2.
6. Discussion
Our reanalysis of the FDA’s benefit-risk assessment [9] of Moderna mRNA-1273 primary-series vaccination reveals substantial shortcomings in the modeling that supported the agency’s finding “the benefits of vaccination with the Moderna COVID-19 vaccine clearly outweigh the risks” even for males of ages 18–25 years. By accounting for prior-infection protection and using evidence-based model inputs, we find in our Scenarios B–E that risks outweighed benefits for 18–25-year-old males, assuming hospitalizations are the principal endpoint of analysis. In particular, for our scenarios predicting plausible infection-level increases over the evaluation period (Scenarios B, D, and E), our model suggests that vaccination of 18–25-year-old males generated between 16% and 63% more hospitalizations from vaccine-attributable myo/pericarditis (VAM/P) alone compared to COVID hospitalizations prevented. [^24] The range 16%–63% is consistent with our model-validation work discussed in Section 5: recall we estimate that at most 214 evaluation-period hospitalizations were prevented by one million mRNA-1273 vaccinations of 18–25-year-old males in the U.S. [Supplement 2, Table S2.7], and note that our corresponding VAM/P hospitalization projection of 268 is a little over 25% higher than 214. Moreover, for the subpopulation of 18–25-year-old males having prior-infection protection and no comorbidities, we have established via Scenarios I and II of Section 4.1 that the choice to mRNA-1273 vaccinate created a greater risk of hospitalization (owing to VAM/P or an evaluation-period breakthrough infection) than the risk of COVD hospitalization if unvaccinated.
The shortcomings of the FDA’s analysis included not accounting for prior-infection protection as well as incidental hospitalizations, and overestimating the COVID-hospitalization rate for 18–25-year-old males by assuming it to be equivalent to the rate for 18–45-year-old males. On the risk side, as discussed in Section 2.3, the FDA relied on a very small internal dataset (21 second-dose events for mRNA-1273 [44]) to determine mRNA-1273 VAM/P risk rather than relying on findings from regulatory agencies from 5 other countries, from the CDC’s VSD system, from major international studies [42, 43], and from a U.S. study [45]. These shortcomings contributed to the FDA’s obtaining in its most likely scenario a benefit-risk ratio of 43.33 for mRNA-1273 vaccination of 18–25-year-old males, a ratio that is over 60 times more favorable than that of our most likely scenario, 0.67.
6.1. The strength of prior-infection protection
The FDA decided to not account for prior-infection protection: “We did not factor in protection from potential exposure or infection with COVID-19 among unvaccinated individuals into our analysis since the degree of this protection and its time-dependency were not well understood at the time of the analysis” [9, Section 2.2.1.1]. However, the protection provided by mRNA-1273 vaccination against the Omicron variant and its time-dependency were also not well understood at the time of analysis (3rd week of January 2022). Nevertheless, the FDA assumed in every one of its five Omicron-based benefit-risk scenarios that mRNA-1273 vaccination would provide for at least 5 months’ protection against cases with 30% effectiveness and against hospitalization with 72% effectiveness. Even for its Delta-based scenario of [9] the FDA assumed prior-infection would provide no protection, despite (i) the CDC study [18] indicating that prior infection provides superior protection against Delta than does mRNA vaccination and (ii) a CDC Science Brief stating, “Available evidence shows that fully vaccinated individuals and those previously infected with SARS-CoV-2 each have a low risk of subsequent infection for at least 6 months” ([17], 10/29/21).
As noted in Section 2.2, the FDA’s assumption that prior infection would provide no protection against Omicron infection and hospitalization is contrary to immunological expectations and studies (e.g., [16, 17, 18, 27, 28]) available to the FDA at the time of its mRNA-1273 analysis all suggesting that prior-infection protection would likely equal or exceed vaccine protection among the COVID-naïve. As noted in Section 3.3, we have assumed that protection against hospitalization provided by vaccination is 77% while that provided by prior infection is 88.9%. A systematic review [25] involving 21 vaccine-effectiveness studies from the period 3 December 2021 – 7 April 2022 reports,
In contrast to vaccine effectiveness against delta severe disease, the majority of vaccine effectiveness estimates for omicron severe disease were below 75%; for example, thirteen (81%) of sixteen vaccine effectiveness estimates within three months of vaccination with the primary series were below 75% (Figure). Moreover, 13 (42%) vaccine-specific estimates fell below 50% at some point in time after vaccination.
In contrast, the meta-analysis [22] estimates prior-infection protection against severe Omicron to be over 80% through 6 months after infection, while the meta-analysis [23, Table S2] estimates protection to be over 88% through 45 weeks after infection, with the former analysis based on studies from the period 1/1/22 – 6/1/22 and the latter, on studies published up to 9/31/22. In early January of 2022, the minimum reasonable assumption the FDA might have made is that prior-infection protection equals vaccine protection (as in our Scenario E); instead, in all Omicron-based scenarios of [9], the agency ascribed to vaccination 72% effective protection against hospitalization while ascribing 0% effective protection to prior infection.
The FDA’s reliance on the UK Health Security Agency for its estimates of vaccine effectiveness against Omicron [34], instead of relying on the CDC, is surprising. The CDC might have had precisely the data the FDA needed for its mRNA-1273 assessment. For example, if the CDC extended its May–November 2021 study [18] through the end of December 2021, then the CDC might have had estimates of vaccine as well as prior-infection effectiveness against Omicron at the time of the FDA’s mRNA-1273 analysis. The FDA’s assessments of BNT162b2 [8] and mRNA-1273 [9] suggest that the FDA does not necessarily request from the CDC data needed to conduct rigorous analyses. For both assessments, the FDA failed to obtain data on incidental COVID hospitalizations. For both assessments, the FDA claims that age/sex specific VAM/P information in the CDC’s VSD system is not available; e.g., Funk et al. claim “[W]e have no access to VSD data for smaller age/sex subgroups” [8, Section 2.3.2.1]. Finally, for its mRNA-1273 assessment, the FDA failed to request that the CDC disaggregate COVID hospitalization data for the age range 18–45 into data for the subranges 18–25, 26–35, and 35–45. Also troubling is the withholding of and underreporting of ICU-admissions data for VAM/P in, respectively, [8] and [9] (see Section 2.4) as well as willingness to use international data [34] for the benefit analysis of [9] while eschewing international data relating to risks (e.g., VAM/P data from Public Health Ontario [3, 42] and postvaccination myocarditis data from England ([6, 43]).
6.2. Summary
We have shown that extending the FDA’s model to account for prior-infection protection as well as incidental COVID hospitalizations and making evidence-based model-input assumptions is sufficient to invert the calculus of seemingly overwhelming benefit of mRNA-1273 vaccination towards excessive risk of VAM/P in the general population of 18–25-year-old males. The failure of the FDA author team of [9] to perform a rigorous benefit-risk analysis led to COVID-vaccination recommendations that both failed reflect “The health condition [and prior-infection status] of individuals may remarkably impact the B-R profile and its evaluation” ([9, Section 4]) and that supported vaccination mandates enforced by many institutions placing certain subgroups at a greater risk of hospitalization for a serious adverse event owing to COVID vaccination than for hospitalization owing to COVID infection—while on an absolute basis this risk difference is small, institutional responsibility for adverse events owing to vaccination is not.
7. Recommendations
Our modeling provides a call to action to compel our policy-forming health agencies to employ multi-parameter models that can better stratify vaccination recommendations. To protect vaccine acceptance and promote public trust, [^25] A survey from the Robert Wood Johnson Foundation and Harvard School of Public Health released May of 2021 found only 37% of respondents had a great deal of trust in the FDA, while 38% somewhat trusted the agency, and 24% described their level of trust “Not very much/Not at all” [74]. the FDA must adopt a rigorous framework for benefit-risk assessment of new vaccines and revisit recommendations as new data becomes available. In particular, the framework must incorporate consideration of the protection that prior infection affords against development of severe disease.
A re-alignment of priorities of the public-health agencies of the United States is necessary. Two years after the first cases SARS-CoV-2 were identified in the U.S., the FDA decided to “not investigate the benefits and risks [of mRNA-1273 vaccination] for specific subpopulations, such as those with comorbidities, due to limited information for these populations” [9, Section 4]. Also, public-health agencies should have prioritized understanding the effectiveness of prior-infection protection against reinfection and hospitalization. That said, in Pfizer’s phase-3 trial for BNT162b2, 567 participants in the placebo group and 526 in the treatment arm had evidence of prior COVID-19 infection. In each of these cohorts, there was only 1 reinfection (or ¡0.2% for both) according to the primary endpoint definition [75, p. 27]. This alone should have prompted health agencies to aggressively pursue further data on protection from prior infection to stratify risk and recommendations. We note that by late fall of 2022, the percentage of the U.S. population having experienced at least one COVID infection was estimated to be 95% by an IHME study [76] and 94% by a Harvard study [77].
Given that nearly all persons in the United States now have significant protection from severe COVID conferred by prior-infection, there is an urgent need to reassess whether the following current blanket CDC recommendations [78] for COVID-19 vaccination advance public health: “ Everyone ages 6 months and older should get a 2024–2025 COVID-19 vaccine.” Compare the CDC’s policy to that of the UK: “All those aged 65 or over, all those in care homes for older people, and those aged 6 months or over who are in a clinical risk group will be eligible for vaccination this autumn” [79].
The urgency of the need for reassessment is reinforced by the fact that Pfizer’s and Moderna’s reformulated COVID-19 vaccines are available only via emergency-use authorization for individuals 6 months through 11 years of age (despite the expiration of the COVID-19 public health emergency 11 May 2023).
8. Limitations
We have assumed that young males are hospitalized for COVID at a higher rate than young females, as did the FDA in its benefit-risk assessments [8] and [9]. Data from [8] and [54] suggests this rate difference is substantial; however, the rates of [8] and [54] include incidental hospitalizations and incidental rates for males may differ from those for females. We have assumed, as did the FDA, that protection provided by vaccination and prior infection against hospitalization and infection is essentially equivalent for males and females, and this may not be the case. Also, we have based our assessment of the difference in hospitalization rates between young males and females on data in Table 2 of [8] representing mostly pre-Delta cases. Relative to Omicron, sex-based differences in hospitalization rates might have changed. However, the rate differences provided by Table 1 of [54], based on two years of data including that of the initial Omicron wave, are consistent with those of Table 2 of [8].
COVID-hospitalizations data from Public Health Ontario (see Supplement 2) and from the State of Connecticut (see the appendix of Supplement 1) suggests that that our Omicron IHR value $0.0014$ for 18–25-year-old males is a reasonable estimate for the Omicron IHR for all 18–25 year-olds (general population mix of males and females). Thus, even if we assume males and females in this age range are hospitalized for COVID at roughly equivalent rates, our IHR value of 0.14% for 18–25-year-old males appears to be a good estimate.
As noted in Section 3.3, based on data available before 1/22/22, we arrived at an Omicron IHR estimate for all 18–25 year-olds of 0.22%, not 0.14%. One possible explanation is that we may have underestimated the incidental COVID-hospitalization rate for 18–25 year-olds. Our IHR estimate for 18–25 year-olds of 0.22% results from assuming an incidental COVID-hospitalization rate of 40%, which is the rate suggested by [29] and [31] for hospitalizations with COVID among patients of all ages. However, there is strong evidence that incidental COVID-hospitalization rates are higher among the young than among the elderly; see, e.g., [30], [80], and Supplement 1. Figure 2 of [30] suggests an incidental rate of over 60% among 0–18 year-olds from April of 2020 through May of 2021. COVID-hospitalization data from the CDC and the state of Connecticut, discussed in the appendix of Supplement 1, suggests an incidental rate of about 72% among 15–24 year-olds in Connecticut during calendar year 2020. Assuming an incidental rate of about 62.2% (instead of 40%) transforms our IHR estimate of 0.22% into the estimate 0.14%.
In terms of applying our model to an emerging pathogen and vaccine campaign in the future, our functions $U$ and $V$ require reasonably accurate inputs to produce reasonably accurate projected hospitalizations as outputs. In particular, if public-health agencies do not provide accurate rates of hospitalizations (requiring treatment for the infection) based on age, vaccination-status, sex, and prior-infection status, then obtaining accurate infection-hospitalization rates and rate-reduction factors becomes difficult.
In applying our improvement of the FDA’s model for the general population of 18–25-year-old males, we constrained ourselves to using only information available at the time of the FDA’s benefit-risk analysis of mRNA-1273 .
Our assessment of risks associated with vaccination, like the FDA’s, ignores known rare but potentially severe vaccine-related adverse events other than myo/pericarditis that may require hospitalization, for example, anaphylaxis [58], Multisystem Inflammatory Syndrome [59, 60], Guillain Barré Syndrome [81], posterior orthostatic tachycardia syndrome (POTS) [82], retinal vascular occlusion [83], and cerebral venous sinus thrombosis (CVST) [64]. Our risk assessment as well as the FDA’s ignores the possibility that COVID-19 vaccination has long-term health effects; for instance, the FDA’s package-insert information [84, Section 13] for Moderna’s mRNA-1273 begins, “SPIKEVAX has not been evaluated for carcinogenic, mutagenic potential, or impairment of male fertility in animals.”
Finally, we have carried out a retrospective analysis illustrating how the FDA might have conducted its mRNA-1273 benefit-risk assessment in a more rigorous way. Data we rely on for assessing benefit and risks of vaccination pertains to the original, monovalent formulation of Moderna’s COVID vaccine, having a two-dose vaccination schedule.
Acknowledgments: We thank Retsef Levi for providing comments that improved both the structure and content of this paper.
Links to Supplements:

### v3

@@
-A Reanalysis of the FDA’s Benefit-Risk Assessment of Moderna’s mRNA-1273 COVID Vaccine: For 18–25-Year-Old Males, Risks Exceeded Benefits Relative to Hospitalizations
-Abstract: The Food and Drug Administration (FDA) conducted a benefit-risk assessment for Moderna’s COVID vaccine mRNA-1273 prior to its full approval (announced 1/31/2022).
-The FDA focused its assessment on males of ages 18–64 years because the agency’s risk analysis was limited to vaccine-attributable myocarditis/pericarditis (VAM/P) given the excess risk among males.
-The FDA analysis concluded that vaccine benefits clearly outweighed risks, even for 18–25-year-old males (those at highest VAM/P risk).
-In this paper, we reanalyze the FDA benefit-risk assessment using information available through the third week of January 2022 and focusing on 18–25-year-old males.
+A reanalysis of the FDA’s benefit-risk assessment of Moderna’s mRNA-1273 COVID vaccine stratified not only based on age and sex but also on prior-infection and comorbidity status
+Abstract.
+We demonstrate that the outcome of a benefit-risk assessment for vaccination may be dramatically impacted by accounting for the benefits derived from prior infection by the disease against which the vaccine is targeted.
+The United States Food and Drug Administration (FDA) conducted a benefit-risk assessment for Moderna’s COVID vaccine mRNA-1273 prior to its full approval, announced 1/31/2022.
+The FDA’s assessment focused on males of ages 18–64 years because the agency’s risk analysis was limited to vaccine-attributable myocarditis/pericarditis (VAM/P) given the excess risk among males.
+The FDA’s analysis concluded that vaccine benefits clearly outweighed risks, even for 18–25-year-old males (those at highest VAM/P risk).
+We reanalyze the FDA’s benefit-risk assessment using information available through the third week of January 2022 and focusing on 18–25-year-old males.
 We use the FDA’s framework but extend its model by accounting for protection derived from prior COVID infection, finer age-stratification in COVID-hospitalization rates, and incidental hospitalizations (those of patients who test positive for COVID but are being treated for something else).
@@
 With hospitalizations as the principal endpoint of the analysis (those prevented by vaccination vs.
-those caused by VAM/P), we find a net harm: vaccine risks outweighed benefits for 18–25-year-old males, except in scenarios projecting implausibly high Omicron-infection prevalence.
-Comparing the projections of FDA’s “most likely scenario” to ours, we find that the FDA’s benefit-risk ratio, 43.33, is over 60 times more favorable than ours, 0.67.
+those caused by VAM/P), our model finds vaccine risks outweighed benefits for 18–25-year-old males, except in scenarios projecting implausibly high Omicron-infection prevalence.
 Our assessment suggests that mRNA-1273 vaccination of 18–25-year-old males generated between 16% and 63% more hospitalizations from vaccine-attributable myocarditis/pericarditis alone compared to COVID hospitalizations prevented (over a five-month period of vaccine protection assumed by the FDA).
-The preceding assessment is based on assumptions drawn from data available before 1/22/22.
-Moreover, these assumptions are validated by subsequently available data.
+The preceding assessment derives from model inputs based on data available at the time of the FDA’s mRNA-1273 assessment.
+Moreover, these inputs as well as model outputs are validated by subsequently available data.
 1. Introduction
@@
 The Food and Drug Administration (FDA) issued emergency-use authorizations for mRNA COVID-19 vaccines from Pfizer (BNT162b2) and Moderna (mRNA-1273) in December 2020 [2].
-Once the mRNA COVID-19 vaccines were given to large populations, it became evident that the vaccines could cause myocarditis and pericarditis (hereafter myo/pericarditis), inflammation of the heart and of the membrane surrounding the heart, respectively, with studies indicating the highest risk among males younger than 40 years ([3, 4, 5, 6, 7]).
+Once the mRNA COVID-19 vaccines were given to large populations, it became evident that they could cause myocarditis or pericarditis (hereafter myo/pericarditis), inflammation of the heart or of the membrane surrounding the heart, respectively, with studies indicating the highest risk among males younger than 40 years [3, 4, 5, 6, 7, 8, 9].
 The FDA has published two benefit-risk analyses of primary series mRNA vaccination, the first for BNT162b2 [10] and the second for mRNA-1273 [11].
-The FDA presented its analyses through a number of scenarios involving different modeling assumptions to account for uncertainties associated with pandemic dynamics as well as with the benefits and risks of vaccination.
-To assess vaccine benefits, the FDA estimated vaccine-preventable COVID-19-cases, hospitalizations, intensive care unit (ICU) admissions, and deaths [8, 9].
+The FDA presented its analyses through scenarios involving different modeling assumptions to account for uncertainties in pandemic dynamics as well as in the benefits and risks of vaccination.
+To assess vaccine benefits, the FDA estimated vaccine-preventable COVID-19 cases, hospitalizations, intensive-care unit (ICU) admissions, and deaths [10, 11].
 To assess vaccine risks, the FDA estimated vaccine-attributable myo/pericarditis (VAM/P) cases, hospitalizations, ICU admissions, and deaths [10, 11].
-The FDA concluded that the benefits of the mRNA vaccines outweighed the risks of VAM/P even for males in its highest-risk age-groups: 16–17 years for BNT162b2 [8] and 18–25 years for mRNA-1273 [9].
-However, both of the FDA’s benefit-risk assessments omitted important inputs in the analysis of the hospitalizations-prevented benefit of vaccination as well as in the analysis of VAM/P risk.
-We focus mainly on the FDA’s benefit-risk analysis for Moderna’s mRNA-1273 to simplify our exposition—there are significant differences in modeling assumptions in the FDA’s Pfizer and Moderna assessments.
-Also, VAM/P risk is substantially greater for mRNA-1273 than for BNT162b2 [3, 5, 6, 10].
+The FDA concluded that the benefits of the mRNA vaccines outweighed VAM/P risk even for males in its highest-risk age-groups: 16–17 years for BNT162b2 [10] and 18–25 years for mRNA-1273 [11].
+However, these FDA benefit-risk assessments omitted important inputs in the analysis of both the hospitalizations-prevented benefit of vaccination and VAM/P risk.
+We focus mainly on the FDA’s analysis for Moderna’s mRNA-1273 to simplify our exposition—there are significant differences in modeling assumptions in the FDA’s Pfizer and Moderna assessments.
+Also, VAM/P risk is substantially greater for mRNA-1273 than for BNT162b2 [3, 4, 5, 6, 7, 8, 9].
 For the FDA’s assessment of mRNA-1273, benefits of vaccination were overestimated.
-First, the FDA did not account for the protection conferred by prior infection; and, at the time of the FDA’s Moderna analysis (mid-January 2022), it is likely that over 70% of 18–29 year-olds had been infected by COVID-19—see Supplement 1 for details.
+First, the FDA did not account for the protection conferred by prior infection; and, at the time of the FDA’s Moderna analysis (mid-January 2022), likely over 70% of 18–25 year-olds in the U.S.
+had been infected by COVID-19—see Supplement 1.
 Second, the FDA did not account for incidental COVID-19 hospitalizations.
@@
 Finally, the FDA assumed in its most likely scenario that case and hospitalization rates would remain constant at peak 2021 Omicron levels for 5 months.
-By contrast, the FDA underestimated VAM/P risks for mRNA-1273.
-The agency omitted dose-one mRNA-1273 VAM/P events [9, Section 2.3.1] (but included dose-one events for its BNT162b2 assessment [8, Section 2.3.1]).
-Moreover, the FDA’s analysis is inconsistent with evidence from the CDC that VAM/P risk for mRNA-1273 was over twice that for BNT162b2 [10, p.
-25].
-Finally, the FDA relied on internal FDA databases that yielded mRNA-1273 VAM/P rates substantially lower than those found in major international studies [3, 5, 6] as well as a U.S.
-study [4].
-In this study, our primary aim is to extend and improve the FDA’s benefit-risk model for mRNA-1273 vaccination by accounting for prior-infection protection, incidental COVID-19 hospitalizations, and age-specific infection-hospitalization rates, as well as for a more accurate assessment of VAM/P risk.
-We apply our model using data available to the FDA at the time of its benefit-risk analysis for mRNA-1273, which we take to be the 3rd week of January 2022 for reasons described Section 2.1 below.
+By contrast, the FDA underestimated mRNA-1273 VAM/P risks; see, e.g.,Table 3 of Section 2.3.
+Here, our primary aim is to demonstrate how to incorporate the protection derived from prior infection into a population-level quantitative benefit-risk model for a two-dose vaccination series.
+Our model has hospitalizations as its endpoint, comparing the hospitalizations-prevented benefit of vaccination to the hospitalization risk attributable to vaccination.
+We illustrate the utility of our model by applying it to extend and improve the FDA’s mRNA-1273 assessment relative to its hospitalizations endpoint, which can be viewed as the most important of the agency’s endpoints owing to the public-health concern that during waves of the COVID pandemic hospitals might be so overwhelmed that they would not be able to deliver high quality care to their patients.
+In addition to prior-infection protection, our modeling accounts for incidental hospitalizations, age-specific infection-hospitalization rates, and a more accurate assessment of VAM/P risk.
+We apply our model using data available to the FDA at the time of its benefit-risk analysis for mRNA-1273, assumed to be the 3rd week of January 2022 for reasons described in Section 2.1.
+We present evidence, using real-world data, that our model’s estimates of the hospitalizations-prevented benefit of mRNA-1273 vaccination are significantly more accurate than those the FDA obtained in [11].
 We also analyze benefits and risks of vaccination for specific subpopulations such as males 18–25 having prior-infection protection and no comorbidities.
-Our modeling not only addresses the shortcomings in the FDA’s assessment of the mRNA-1273 vaccine but also presents a framework for future benefit-risk assessments of vaccination.
+Our model can be adapted to produce for any vaccine designed for deployment on a population-wide scale a benefit-risk assessment stratified not only based on age and sex but also on prior-infection and comorbidity status.
 2. The FDA’s mRNA-1273 Benefit-Risk Assessment
-In [9], the FDA describes its benefit-risk modeling of Moderna’s mRNA-1273 (Spikevax) for males ages 18 to 64 years.
+In [11], the FDA describes its benefit-risk modeling of Moderna’s mRNA-1273 for males ages 18 to 64 years.
 Outcomes are provided via six scenarios reflecting uncertainties in three major model inputs: COVID-19 incidence, vaccine effectiveness against cases (VE) as well as hospitalization (VEH), and vaccine-attributable myo/pericarditis (VAM/P) rate.
-At the time of the FDA’s mRNA-1273 analysis, the Centers for Disease Control and Prevention (CDC) was recommending a 5-month interval between primary-series completion and a booster dose.
+At the time of the FDA’s mRNA-1273 analysis, the CDC was recommending a 5-month interval between primary-series completion and a booster dose.
 Consequently, the FDA evaluated benefits and risks over a 5-month period, which we take to be 1/1/22 – 5/31/22.
@@
 The FDA’s major modeling assumptions are summarized in Table 1 below.
-Table 1. The FDA’s mRNA-1273 model-input assumptions for 18–25-year-old males [9, Table 1].
-Scenario COVID-19 Incidence Vaccine Effectiveness VAM/P Rate Assumptions Per Million 2nd Vaccinations*
-$1$ Highest 2021 Incidence, 12/25–12/31 Omicron Dominant; VE = 30%, VEH=72% 128 cases, 110 hospitalizations
+Table 1. The FDA’s mRNA-1273 model-input assumptions for 18–25-year-old males [11, Table 1].
+Scenario COVID-19 Incidence Vaccine Effectiveness VAM/P Rate Per Million 2nd doses [11, Table 2]
+11 Highest 2021 Incidence, 12/25–12/31 Omicron Dominant; VE = 30%, VEH=72% 128 cases, 110 hospitalizations
 2 Average 2021 Incidence Same as Scenario 1 Same as Scenario 1
@@
 6 Same as Scenario 1 Same as Scenario 1 241 cases, 207 hospitalizations
-*From [9, Table 2].
-Using the modeling assumptions described in the first row of the Table 1, with incidence data drawn from [9, Table 3], the FDA concluded for Scenario 1, its “most likely scenario,” that vaccinating one million males ages 18–25 years with two doses would prevent, over 5 months, 82,484 COVID-19 cases, 4,766 COVID-19 hospitalizations, 1,144 ICU admissions, and 51 deaths due to COVID-19, while risking 128 VAM/P cases, 110 VAM/P hospitalizations, 0 ICU admissions, and 0 deaths due to VAM/P [9, Table 2].
-However, the FDA’s model overestimates benefits and underestimates risks.
+Using inputs from first row of Table 1 and incidence data from [11, Table 3], the FDA concluded for Scenario 1, its “most likely scenario,” that vaccinating one million males ages 18–25 years with two doses would prevent, over 5 months, 82,484 COVID-19 cases, 4,766 COVID-19 hospitalizations, 1,144 ICU admissions, and 51 deaths due to COVID-19, while risking 128 VAM/P cases, 110 VAM/P hospitalizations, 0 ICU admissions and 0 deaths from VAM/P [11, Table 2].
+.
 2.1. The Time of the FDA’s mRNA-1273 Analysis
-Although the FDA does not define in [9] the time of its benefit-risk analysis for mRNA-1273, a natural choice is the 3rd week of January 2022, which concluded 10 days prior to the FDA’s announcement of the full approval of mRNA-1273 on 1/31/2022 [12].
-This choice is natural because it is consistent with the FDA’s time-of-analysis definition for its BNT162b2 assessment: the 2nd week of August 2021 [8, Section 2.1], which concluded nine days prior to the FDA’s announcement of the full approval of BNT162b2 on 8/23/2021 [11].
-Consequently, our reanalysis of the FDA’s benefit-risk assessment [9] for the general population of 18–25-year-old males is based on assumptions derived from information available before 1/22/2022.
-Subsequent information validating our assumptions will often be provided but only via footnotes.
-Starting in Section 4.2, we no longer restrict ourselves to basing modeling assumptions only on information available before 1/22/2022; in this section, our intent is to demonstrate how our model might be modified to incorporate a comorbidity (obesity) as a risk modifier.
-Starting in Section 4.2, where we explore how vaccination’s benefit-risk picture changes with respect to Body Mass Index (BMI), we no longer restrict ourselves to basing modeling assumptions only on information available before 1/22/2022.
+Although the FDA does not define in [11] the time of its benefit-risk analysis for mRNA-1273, a natural choice is the 3rd week of January 2022, which concluded 10 days before the FDA’s announcement of the full approval of mRNA-1273 on 1/31/2022 [12].
+This choice is consistent with the FDA’s time-of-analysis definition for its BNT162b2 assessment: the 2nd week of August 2021 [10, Section 2.1], which concluded nine days before the FDA’s announcement of the full approval of BNT162b2 on 8/23/2021 [13].
+Consequently, our reanalysis of the FDA’s benefit-risk assessment [11] is based on model inputs derived from information available before 1/22/2022.
+Subsequent information validating our input choices will often be provided but only via footnotes.
+Starting in Section 4.2, we discontinue basing our modeling on information only available before 1/22/2022; in Section 4.2, our intent is to show how our model can be modified to incorporate a comorbidity (obesity) as a risk modifier.
 2.2. The FDA’s mRNA-1273 Benefits Assessment
@@
 For each scenario, CDC data is used to anticipate over the 5-month evaluation period a certain number of COVID cases $C$ and hospitalizations $H$ per million unvaccinated males, and the corresponding evaluation-period benefits per 1 million vaccinations are estimated to be
-[…]
+$C\cdot VE\ \text{cases prevented and}\ H\cdot VEH\ \text{hospitalizations prevented},$
 where $VE$ and $VEH$ are, respectively, the effectiveness of full mRNA-1273 vaccination against cases and against hospitalization; the FDA assumed that for Omicron $VE=0.30$ and $VEH=0.72$, as noted in Table  1 above.
-For Scenario 1, its most likely scenario, and two others, the FDA assumed that peak 2021 Omicron-case rates (late December 2021) would be sustained for 5 months, resulting in an exceptionally high projection of total COVID-19 cases among the unvaccinated; for example, $C=274,947$ for 18–25-year-old males.
-Similarly, for its most likely scenario (and two others), the FDA assumed that the CDC’s reported late-December 2021 Omicron hospitalization rate for unvaccinated 18–45-year-old males would be sustained for 5 months at approximately 43.54 cases per million daily [9, Table 3] and assumed that the same rate would apply to the younger 18–25 bracket, arriving at $H=6,619$ hospitalizations.
-Thus, for its most likely scenario (and two others), the FDA predicted that full mRNA-1273 vaccination of 1 million 18–25-year-old males would prevent $274,947\cdot 0.30\approx 82,484$ COVID cases and $6619\cdot 0.72\approx 4,766$ COVID hospitalizations [9, Table 2].
-To obtain its projections of ICU admissions prevented by vaccination, the FDA used CDC data to estimate age-based percentages of males hospitalized with COVID going to ICU [9, Table 3].
-According to Table 3 of [9], 24% of 18–25-year-old males admitted with COVID require ICU care; thus, the FDA predicted that, under its most likely scenario, fully vaccinating 1 million males in this age group with mRNA-1273 would prevent approximately 1,144 ICU admissions (24% of 4,766)[9, Table 2].
-The FDA’s projections of deaths prevented by vaccination utilized the same assumptions as hospitalizations prevented.
-Specifically, for its most likely scenario, the FDA assumed the CDC’s reported late-December 2021 Omicron death rate for unvaccinated males 18–25 would be sustained for 5 months resulting in a total of about 71 deaths per million and assumed vaccine effectiveness against death to be the same as VEH, estimating $71\cdot 0.72\approx 51$ COVID deaths prevented per million vaccinations.
+For Scenario 1, its most likely scenario, and two others, the FDA assumed that peak 2021 Omicron-case rates (late December 2021) would be sustained for 5 months, resulting in an exceptionally high projection of COVID-19 cases among the unvaccinated; e.g., $C=274,947$ for 18–25-year-old males.
+Similarly, for these scenarios, the FDA assumed that the CDC’s reported late-December 2021 Omicron hospitalization rate for unvaccinated 18–45-year-old males would be sustained for 5 months at approximately 43.54 cases per million daily [11, Table 3] and assumed that the same rate would apply to the younger 18–25 bracket, arriving at $H=6,619$ hospitalizations.
+Thus, for these scenarios, the FDA predicted that full mRNA-1273 vaccination of 1 million 18–25-year-old males would prevent $274,947\cdot 0.30\approx 82,484$ COVID cases and $6619\cdot 0.72\approx 4,766$ COVID hospitalizations over the 5-month evaluation period [11, Table 2].
 The following factors contributed to the FDA’s overestimation of vaccine benefits:
-$\bullet$ The FDA did not account for the protection conferred by prior infection.
-The FDA’s not accounting for prior-infection protection in its mRNA-1273 assessment is difficult to justify.
-A study of SARS-CoV-2–specific immune memory found “Substantial immune memory is generated after COVID-19,” with about 95% of subjects retaining memory at  6 months post infection [15].
+$\bullet$
+The FDA did not account for the protection conferred by prior infection.
+The FDA’s disregarding prior-infection protection in its mRNA-1273 assessment is difficult to justify.
+Dan et al.
+found “About 95% of subjects retained immune memory at $\sim$6 months after [COVID] infection” [14].
 A meta-analysis based on studies published December 2020 through August 2021 demonstrated that protection from prior COVID infection “is, at least, equivalent to the protection afforded by complete vaccination of COVID-naïve populations” [15].
-The CDC acknowledged in a science brief released 10/29/21 that “The immunity provided by vaccine and prior infection are both high” [17], while another CDC study found that during the Delta wave unvaccinated persons with prior COVID infection had significantly greater protection against infection and hospitalization than fully vaccinated, COVID-naïve persons [18].
-Moreover, by the end of December 2021, the FDA acknowledged, “The omicron variant has significantly more mutations than previous SARS-CoV-2 variants, particularly in its S-gene, the gene that encodes the virus’s spike protein”[19], suggesting that prior infection, which entails exposure to all SARS-CoV-2 proteins (such as the nucleoprotein), would likely yield greater protection against Omicron than would mRNA-1273 vaccination, which entails exposure only to the spike protein of the ancestral strain of SARS-CoV-2.
-The summary of findings of a study published 7/20/21 concludes, “CD4+ T cell responses equally target several SARS-CoV-2 proteins, whereas the CD8+ T cell responses preferentially target the nucleoprotein, highlighting the potential importance of including the nucleoprotein in future vaccines.
-Taken together, these results suggest that broad and effective immunity may persist long-term in recovered COVID-19 patients[20].
-[^1] Consistent with the immunological considerations discussed in this paragraph, a study from Qatar found prior infection provides 52.2% effective protection against symptomatic Omicron infection and 91% effective protection against “severe, critical, or fatal” Omicron infection; whereas, the study’s corresponding effectiveness estimates for 2 doses of mRNA-1273 are 2.2% and 66.3%, respectively [21, Figure 2C, 2D] (study period 12/23/21 – 2/21/22).
-In Section 7 of Supplement 1, we explain how the meta-analyses and systematic reviews [22, 23, 24, 25] provide convincing evidence that prior infection provided greater protection against Omicron during the evaluation period than did two-dose vaccination among the COVID-naïve.
-Additional reasons why the FDA’s not accounting for prior-infection protection is difficult to justify include (i) according to the CDC, 54.9% of those 18–49 years old had been infected by COVID-19 by the end of September 2021 [26]; (ii) a South African study [27] released 12/14/21 suggested overall effectiveness of prior-infection protection against Omicron to be about 47% (see Supplement 1), while data from England [28] released 12/22/21 suggests that prior infection among the unvaccinated provides significant protection from Omicron hospitalization (see Supplement 1); and (iii) the FDA adopted a scenario-based structure for its benefit-risk assessment to account for “the impact of uncertainty associated with pandemic dynamics” [9, Abstract], yet the agency did not consider prior-infection protection in any of its six scenarios.
+The CDC acknowledged in a science brief released 10/29/21 that “The immunity provided by vaccine and prior infection are both high” [16], while another CDC study found that during the Delta wave unvaccinated persons with prior COVID infection had significantly greater protection against infection and hospitalization than did fully vaccinated, COVID-naïve persons [17].
+Moreover, in December 2021, the FDA acknowledged, “The omicron variant has significantly more mutations than previous SARS-CoV-2 variants, particularly in its S-gene, the gene that encodes the virus’s spike protein”[18], suggesting that prior infection, entailing exposure to all SARS-CoV-2 proteins (such as the nucleoprotein), would likely yield greater protection against Omicron than would mRNA-1273 vaccination, which entails exposure only to the spike protein of the ancestral strain of SARS-CoV-2.
+Among the findings of a study of SARS-CoV-2 specific B and T cell memory is “CD8+ T cell responses preferentially target the nucleoprotein, highlighting the potential importance of including the nucleoprotein in future vaccines” [19].
+[^1] Consistent with the evidence presented in this paragraph, a Qatari study found prior infection provided 52.2% effective protection against symptomatic Omicron infection and 91% effective protection against “severe, critical, or fatal” Omicron, with corresponding effectiveness estimates for 2 doses of mRNA-1273 being 2.2% and 66.3%, respectively [20, Figure 2C, 2D] (study period 12/23/21 – 2/21/22).
+Findings from the meta-analyses and systematic reviews [21, 22, 23, 24] provide convincing evidence that prior infection provided greater protection against Omicron during the evaluation period than did two-dose vaccination among the COVID-naïve [Supplement 1, Section 7].
+Further evidence of the need to account for prior-infection protection includes (i) a CDC estimate that 54.9% of 18–49 year-olds had been infected by COVID-19 by 9/30/2021 [25], (ii) a South African study [26] released 12/14/21 suggesting overall effectiveness of prior-infection protection against Omicron to be about 47% [Supplement 1, Section 4], and (iii) data from England [27] released 12/22/21 suggested that prior infection among the unvaccinated provides significant protection from Omicron hospitalization [Supplement 1, Section 5].
 $\bullet$ The FDA did not account for incidental COVID-19 hospitalizations.
-At the time of its mRNA-1273 assessment, it was well known that incidental-hospitalization rates were substantial.
-The day after CDC director R.
-Walensky [29] stated during a 1/9/22 interview “[U]p to 40 percent of the patients who are coming in with COVID-19 are coming in …with something else,” [^2] We searched unsuccessfully for published CDC data on incidental COVID-hospitalization rates.
+At the time of its mRNA-1273 assessment, incidental-hospitalization rates were well known to be substantial.
+On 1/9/22, the director of the CDC stated, “[U]p to 40 percent of the patients who are coming in with COVID-19 are coming in …with something else” [28].
+[^2] We searched unsuccessfully for published CDC incidental-COVID-hospitalizations data.
 Figure 2 of a large U.S.
-study suggests an incidental rate for 0–18 year-olds typically well above 50% during the study period 3/1/2020–1/31/22 and for those 19–64 an incidental rate above 40% for 1/1/2022–1/31/2022 [30].
-Canada’s province of Ontario started publishing incidental rates, with rates for January 8–15 exceeding 45% [31].
-There is reason to believe incidental rates are especially high for the young—see Supplement 1.
+study suggests an incidental rate for 0–18 year-olds typically well above 50% during its study period 3/1/2020–1/31/22 and for those 19–64 an incidental rate above 40% for 1/1/2022–1/31/2022 [29].
+On 1/10/22, Ontario began publishing incidental rates, with rates for January 8–15 exceeding 45% [30].
+There is evidence that incidental rates are especially high for the young.
+For example, a major study [31] found, “As of May 2021, the cumulative rate for hospitalization associated with SARS-CoV-2 infection was 35.9 per 10,000 children” and “When limiting the analysis to include only patients with COVID-19 who required therapeutic interventions, the hospitalization rate decreased 5.5-fold to 6.5 per 10,000 children,” which suggests an incidental rate of over 80% “for children” (persons 0-17-years-old).
+See Supplement 1 for additional incidental-hospitalization data.
 $\bullet$ The FDA overestimated the proportion of projected COVID cases requiring hospitalization.
-One way to gauge the degree of implausibility of the FDA’s evaluation-period hospitalization projections is to consider the corresponding case-hospitalization rates:
-Table 2. Case-hospitalization-rate projections for unvaccinated 18–25-year-old males over the evaluation period 1/1/22 – 5/31/22 for the Omicron-based scenarios of [9, Table 2].*
+All of the FDA’s projected case-hospitalization rates, which are described in Table 2 below, are over four times the Omicron case-hospitalization rate ($0.58\%\approx 56/9,585$) for the unvaccinated of all ages based on Omicron cases in England tested between 1 December 2021 and 14 December 202 [27, Table 3]; moreover, all are well over twice the case-hospitalization rate (0.92%) for the unvaccinated of ages 18–29 in Ontario 12/14/20 – 12/12/21 [32, Table A1a].
+[^3] The FDA case-hospitalization rates in Table 2 are all well over 5 times the Omicron rate 0.42% for the unvaccinated of all ages reported in Table S10 of [33] (data from 12/15/21 – 1/17/22).
+Based on information available at the time of the FDA’s mRNA-1273 assessment, we expect the COVID-case hospitalization rate for the period 12/14/20 – 12/12/21 to be higher than that for Omicron; e.g., “The risk of hospital admission from emergency departments with Omicron was approximately one-third of that for Delta (Hazard Ratio 0.33, 95% CI: 0.30 to 0.37)” [34, p.
+4]; see Supplement 1 for more examples.
+Furthermore, we expect the hospitalization rate for those 18–25 to be lower than that for 18–29 [35].
+Factors contributing to the FDA’s overestimating case-hospitalization rates for the 18–25 age range include (i) assuming hospital-admission rates were uniform for ages 18–45, contradicting CDC models [36], and (ii) including incidental COVID hospitalizations.
+Our source for the 0.58% rate mitigated the incidental-hospitalization issue [27, p.
+10], while our source [32] for the 0.92% rate did not.
+Table 2. The FDA’s case-hospitalization-rate projections for unvaccinated 18–25-year-old males over the evaluation period 1/1/22 – 5/31/22 for the Omicron-based scenarios of [11, Table 2].*
 Scenario Projected Cases Projected Hospitalizations Projected Case-Hospitalization Rate
-$1$   274,947   6,619 $\frac{6619}{274,947}\approx 2.41\%$
+11   274,947   6,619 $\frac{6619}{274,947}\approx 2.41\%$
 2 89,017 2,900 $\frac{2900}{89,017}\approx 3.26\%$
 3 13,010 882 $\frac{882}{13,010}\approx 6.78\%$
-$5$ 274,947 6,619 $\frac{6619}{274,947}\approx 2.41\%$
-$6$ 274,947 6,619 $\frac{6619}{274,947}\approx 2.41\%$
+55 274,947 6,619 $\frac{6619}{274,947}\approx 2.41\%$
+66 274,947 6,619 $\frac{6619}{274,947}\approx 2.41\%$
 *Case and hospitalization projections are obtained from Table 2 of [11] by dividing, respectively, cases-prevented estimates by 0.30 and hospitalizations-prevented estimates by 0.72.
-The FDA’s projected case-hospitalization rates are all over four times the Omicron case-hospitalization rate ($0.58\%\approx 56/9,585$) for the unvaccinated of all ages based on Omicron cases in England tested between 1 December 2021 and 14 December 2021, inclusive [28, Table 3]; moreover, all are well over two times the case-hospitalization rate (0.92%) for the unvaccinated of ages 18–29 in Ontario over the period 12/14/20 – 12/12/21 [32, Table A1a].
-[^3] The FDA case-hospitalization rates in Table 2 are all well over 5 times the Omicron rate 0.42% for the unvaccinated of all ages reported in Table S10 of [33] and based on data from the period 12/15/21 – 1/17/22.
-Based on information available at the time of the FDA’s mRNA-1273 assessment, one would expect the COVID-case hospitalization rate for the period 12/14/20 – 12/12/21 would be higher than that for Omicron; e.g., according to [34, p.
-4], “The risk of hospital admission from emergency departments with Omicron was approximately one-third of that for Delta (Hazard Ratio 0.33, 95% CI: 0.30 to 0.37)”; see Supplement 1 for more examples.
-Furthermore, we expect the hospitalization rate for the age-group 18–25 to be lower than that for 18–29 [35].
-To some extent, FDA’s exaggerated case-hospitalization rates for the 18–25 age range result from the agency’s assumption that hospital-admission rates were uniform for ages 18–45, contradicting the CDC’s own models [36].
-Also, to some extent, the FDA’s exaggerated case-hospitalization rates result from the CDC’s inclusion of incidental COVID hospitalizations in its accounting.
-Our source for the 0.58% rate mitigated the incidental-hospitalization issue [28, p.
-10], while our source [32] for the 0.92% rate did not.
 $\bullet$ The FDA assumed in its most likely scenario (as well as Scenarios 5, and 6) that case and hospitalization rates would remain constant at peak 2021 Omicron levels for 5 months.
-One way to gauge the degree of implausibility of the FDA’s case-rate projection is to consider the corresponding projected number of infections.
 Throughout the pandemic, the number of COVID infections has always been significantly larger than the corresponding number of reported cases.
@@
 owing to increased home testing) during the Omicron-dominant period in the U.S., which started mid-December 2021.
-If we assume that CDC’s 4:1 infection:case ratio holds over the evaluation period, we find the FDA’s evaluation-period case-count projection of 274,947 per 1 million unvaccinated 18–25-year-old males (Scenarios 1, 5, 6) corresponds to 4 · 274,947 = 1,099,788 infections, well over 1 million.
-Thus, the FDA’s case-total assumption is essentially equivalent to the assumption that among one million unvaccinated males, there would be, over 5 months, 1,000,000 infections.
+If we conservatively assume that the CDC’s 4:1 infection:case ratio holds over the evaluation period, we find the FDA’s evaluation-period case-count projection of 274,947 per 1 million unvaccinated 18–25-year-old males (Scenarios 1, 5, 6) corresponds to 4 · 274,947 = 1,099,788 infections, well over 1 million.
+Rounding down, we deem the FDA’s case-total assumption to be equivalent to the assumption that among one million unvaccinated males, there would be, over 5 months, 1,000,000 infections.
 Having well over one million infections in a group of 1 million 18–25 year-olds over 5 months is implausible.
-Multiple infections over a five-month period would be unlikely [^4] The study [37] conducted over the period 11/1/21 – 4/30/22 provides evidence that multiple evaluation-period Omicron infections were rare: in the study there were 11,307 positive Omicron tests and none was a second positive Omicron test separated by 90 days from another such positive test.
-given the high protection from reinfection resulting from recovery from a recent infection, and any reinfections would be offset by individuals who did not become infected at all (e.g., owing to prior-infection protection, limited exposure to infectious persons, or, possibly, genetic factors [38]).
+Multiple infections over a five-month period would be unlikely given the high protection from reinfection resulting from recovery from a recent infection, and any reinfections would be offset by individuals who did not become infected at all (e.g., owing to prior-infection protection, limited exposure to infectious persons, or, possibly, genetic factors [37]).
 [^4] The preprint [37], posted 6/10/21, was later published [38].
-Even direct inoculation with SARS-CoV-2 does not necessarily produce infection: in the challenge study [40] of the 36 inoculated, 16 “remained uninfected.” Thus, for 3 of the FDA’s 5 Omicron-based scenarios, including its most likely scenario, the FDA essentially assumed that every unvaccinated male 18–25 would be infected by COVID over the 5-month evaluation period.
+Even direct inoculation with SARS-CoV-2 does not necessarily produce infection: in the challenge study [39] of the 36 inoculated, 16 “remained uninfected.” Thus, for 3 of the FDA’s 5 Omicron-based scenarios, including its most likely scenario, the FDA essentially assumed that every unvaccinated male 18–25 would be infected by COVID over the evaluation period.
 2.3. The FDA’s mRNA-1273 Risks Assessment
-On the risk side, the FDA relied on its internal databases for the VAM/P case rates appearing in Table 1 above while using the CDC’s Vaccine Safety Datalink (VSD) system to estimate the percentage of VAM/P cases requiring hospitalization [9, Section 2.3.2].
-The FDA underestimated VAM/P risks for mRNA-1273:
+The FDA relied on its Biologics Effectiveness and Safety (BEST) System for the VAM/P case rates appearing in Table 1 above and used the CDC’s Vaccine Safety Datalink (VSD) system to estimate the percentage of VAM/P cases requiring hospitalization [11, Section 2.3.2].
+VAM/P-case rates vary considerably depending on the case definition employed and the extent of stratification of risk, based on age, sex, vaccine manufacturer and dose, with most cases occurring postdose 2 [9].
+For its mRNA-1273 assessment, the FDA four BEST-System data partners reported, with wide confidence intervals, the following myo/pericarditis incidence rates for 18–25-year-old males during the 7-day interval following dose 2: approximately 35, 70, 146, and 256 cases per million second doses [40, p, 12].
+The FDA’s BEST System uses health-claims data not validated by a complete review of patients’ medical charts [11, Section 4], and has “inherent limitations, such as small sample sizes and imperfect sensitivity of ICD-10 codes to identify these rare outcomes” [10, Section 4].
+We rely on four sources of data to estimate the VAM/P rate for dose 2 among 18–25-year-old males: (i) Buchan et al.
+[41] who apply the Brighton Collaboration’s case definition, (ii) Patone et al.
+[42] who consider only hospitalized cases of myocarditis, (iii) Sharff et al.
+[43] who apply the CDC’s case definition [44], and (iv) the FDA (with VSD-based adjustments).
+The FDA underestimated mRNA-1273 VAM/P risk:
 $\bullet$ The FDA omitted dose-one VAM/P events.
 For its mRNA-1273 analysis [11], the FDA computed the evaluation-period benefits of two doses of mRNA-1273 without accounting for the VAM/P risk associated with dose 1, potentially biasing the FDA’s analysis in favor of vaccination.
-[^6] The FDA explains its decision to exclude VAM/P events for dose 1 of mRNA-1273 as follows: “The risk post dose 1 was ignored since CDC has suggested people who have myocarditis after dose 1 take precaution to receive dose 2, and the majority of myocarditis cases were reported post dose 2” [9, Section 2.3.1].
-In contrast, for BNT162b2-related VAM/P, the FDA added excess myo/pericarditis cases postdose 1 and postdose 2 to approximate the population-level risk of completing 1 million two-dose vaccination series [8, Section 2.3.1].
-Hence, for the BNT162b2 analysis, the FDA did account for VAM/P risk associated with dose 1 but did not assess the benefits of dose 1 alone during the interval starting with dose 1 and ending at the point of full vaccination (7 days after dose 2), potentially biasing the analysis against vaccination.
-Given that adverse events of vaccination other than VAM/P possibly resulting in hospitalization were ignored by the FDA and that mRNA vaccination might have unknown long-term adverse health impacts, we believe that even in a benefit-risk analysis assessing benefits only postdose 2, the choice to include VAM/P risk associated with dose 1 is appropriate, being consistent with the first-do-no-harm principle, the FDA’s choice in its BNT162b2 assessment, and the FDA’s choice in its benefit-risk analysis for mRNA-1273 described in its BLA (Biologics License Application) Clinical Review Memorandum for mRNA-1273—see Section 2.4.
-Thus, we include VAM/P hospitalization risk for dose 1 of mRNA-1273 in our main analysis.
-In Supplement 2, we conduct a sensitivity analysis showing that accounting for dose 1 benefits during the period leading up to the point of full vaccination does not substantially change our benefit-risk conclusions.
-$\bullet$ The FDA ignored evidence that VAM/P risk for mRNA-1273 was over twice that for BNT162b2.
-In a “Review Memorandum Addendum” [41], dated 11/19/21, relating to the emergency-use authorization of a booster dose of mRNA-1273, the FDA describes multiple sources of evidence indicating that VAM/P risk for mRNA-1273 is significantly greater than that for BNT162b2:
-[E]vidence was communicated from other regulatory agencies, including agencies from Canada, France, Germany, Norway, and Japan that suggested that there was a higher rate of myocarditis/pericarditis observed in males under 40 years of age following the second dose of the Moderna COVID-19 Vaccine than with the Pfizer-BioNTech COVID-19 Vaccine.
-For example, Ontario, Canada reported a higher rate of myocarditis in individuals following the Moderna COVID-19 Vaccine than with the Pfizer-BioNTech COVID-19 Vaccine, particularly in males ages 18–24 years ([link provided to Ontario data, later analyzed by Buchan et al.
-[42]]).
-…[T]he most recent data presented by the Center for Disease Control and Prevention (CDC) analyzed from the Vaccine Safety Datalink (VSD) indicates a higher rate of myocarditis with the Moderna COVID-19 Vaccine than the Pfizer-BioNTech COVID-19 Vaccine ….
-The “most recent data” from the VSD to which the FDA refers was presented at the CDC’s October 2021 Advisory Committee on Immunization Practices (ACIP) meeting [10].
-Adjusted-rate ratios for 18–39 year-olds reported in the table on page 25 of [10] suggest that the 2nd-dose VAM/P rate for Moderna’s mRNA-1273 is 2.72 times that of Pfizer’s BNT162b2 and that the corresponding ratio for the 1st dose is 2.63.
-For 18–39-year-old males, the tabulated data on page 27 of [10] suggests a 2nd-dose rate ratio of 2.26, and, if we assume the same percentage decrease for the 1st dose, then that rate ratio would be 2.19 instead of 2.63.
-Data from other sources suggests higher rate ratios, especially for dose 2.
-For example, Buchan et al.
-found a 2nd dose rate ratio of 5.1 for males aged 18–24 [42, p.
-5].
-[^7] The specific data (299.5 per million second doses of Moderna vs.
-59.2 per million of Pfizer) is the same in the publication [3] and its preprint [42] (available 12/5/2021).
+In contrast, for BNT162b2-related VAM/P, the FDA added excess myo/pericarditis cases postdose 1 and postdose 2 to approximate the population-level risk of completing 1 million two-dose vaccination series [10, Section 2.3.1].
+Hence, for BNT162b2, the FDA did account for 1st-dose VAM/P risk but did not assess the benefits of dose 1 alone during the interval starting with dose 1 and ending at the point of full vaccination, potentially biasing the analysis against vaccination.
+Because adverse events of vaccination other than VAM/P were disregarded by the FDA and mRNA vaccination might have unknown long-term adverse health impacts, we believe that even in a benefit-risk analysis assessing benefits only postdose 2, the choice to include 1st-dose VAM/P risk is appropriate, being consistent with the first-do-no-harm principle, the FDA’s choice in its BNT162b2 assessment, and the FDA’s choice in the benefit-risk analysis of its BLA (Biologics License Application) Clinical Review Memorandum for mRNA-1273—see Section 2.4.
+Thus, we include 1st-dose VAM/P risk for mRNA-1273 in our main analysis.
+In Supplement 2, we conduct a sensitivity analysis showing that accounting for dose 1 benefits does not substantially change our benefit-risk conclusions.
+$\bullet$ The FDA disregarded evidence that VAM/P risk for mRNA-1273 was over twice that for BNT162b2.
+In a “Review Memorandum Addendum” [45], dated 11/19/21, the FDA describes multiple sources of evidence that VAM/P risk for mRNA-1273 is significantly greater than that for BNT162b2, including reports from regulatory agencies of five countries, data from Ontario later analyzed by Buchan et al.
+[41], and data from the CDC: “[T]he most recent data presented by the Center for Disease Control and Prevention (CDC) analyzed from the Vaccine Safety Datalink (VSD) indicates a higher rate of myocarditis with the Moderna COVID-19 Vaccine than the Pfizer-BioNTech COVID-19 Vaccine ….”
+The “most recent data” from the VSD to which the FDA refers was presented at the CDC’s October 2021 ACIP meeting [46].
+Adjusted-rate ratios for 18–39-year-old males from the table on page 27 of [46] suggest the 2nd-dose VAM/P rate for mRNA-1273 is 2.26 times that of BNT162b2.
+Data from other sources suggests higher rate ratios e.g,, a 2nd-dose rate ratio of 5.1 was reported by Buchan et al.
+for males 18–24 [41, p.
+5] and of 5.2 was reported by Husby et al.
+for males 12–39 [47, paragraph after Figure 3].
 A study by Patone et al.
-[43] of myocarditis cases requiring hospitalization within a 28-day risk interval following a vaccine dose found among males under 40 a rate of 101 excess myocarditis hospitalizations per million 2nd doses of mRNA-1273 and approximately 12 excess myocarditis hospitalizations per 1 million 1st doses, with corresponding rates for BNT162b2 being 12 per million 2nd doses and 3 per million 1st doses [43, p.
-4].888 Patone et al.
-[43] had a study population of over 42 million and a study period from 12/1/20 – 11/15/21.
-The published version [6] of the preprint [43] incorporates an additional month of data and suggests for males $<40$ a rate of 97 myocarditis hospitalizations per million 2nd doses of mRNA-1273 and a rate of 14 per million 1st doses with corresponding BNT162b2 rates being 11 and 4 [6, Table 4].
-Despite all the evidence reviewed above, the FDA chose for its mRNA-1273 risk assessment [9] to rely on “preliminary findings” based on its internal databases suggesting “No clear difference in the rate of myocarditis/pericarditis [mRNA-1273 vs.
-BNT162b2]” [41, p.
-2].
-The “preliminary findings” to which the FDA refers have “large uncertainty” owing to “small numbers of observed events” [44, p.
-16].
-[^9] As noted in [44, p.
-7] and in [9, Section 2.3.1], VAM/P-data-collection cutoff points, for the four databases on which the FDA relied for its mRNA-1273 risk assessment were: 8/21/21, 7/10/21, 7/31/21, and 6/30/21, each date being over four months before the time of the assessment.
-Based on these preliminary findings as well as ignoring dose-1 events, the FDA suggests in its mRNA-1273 assessment that for 18–25-year-old males VAM/P risk for mRNA-1273 is comparable to that for BNT162b2: 128 cases per million 2nd doses of mRNA-1273 [9, Table 4] vs.
-131 cases per million full vaccinations of BNT162b2 (among 18–24-year-old males) [8, Table 3], with corresponding hospitalization projections of 110 per million 2nd doses of mRNA-1273 and 131 per million full BNT162b2 vaccinations.
+of myocarditis cases requiring hospitalization within a 28-day risk interval after a vaccine dose found among males under 40 a rate of 101 excess myocarditis hospitalizations per million 2nd doses of mRNA-1273 and 12 per million 1st doses, with corresponding BNT162b2 rates being 12 per million 2nd doses and 3 per million 1st doses [42, p.
+4].
+Based on “preliminary findings” [45, p.
+2], having “large uncertainty” owing to “small numbers of observed events” [40, p.
+16], as well as omitting dose-1 cases, the FDA suggests that for 18–25-year-old males VAM/P risk for mRNA-1273 is comparable to that for BNT162b2: 128 cases per million 2nd doses of mRNA-1273 [11, Table 4] vs.
+131 cases per million full vaccinations of BNT162b2 (among 18–24-year-old males) [10, Table 3], with corresponding hospitalization projections of 110 per million 2nd doses of mRNA-1273 and 131 per million full BNT162b2 vaccinations.
 Throughout this paper, we assume that VAM/P rates for males 18–24 well approximate those for males 18–25.
-Based on data from the preceding two paragraphs, we compute a VAM/P-rate estimate for mRNA-1273 using the FDA’s corresponding estimate for BNT162b2 in [8] and VSD VAM/P rate ratios [10].
-FDA data available in October 2021 reports 5 myo/pericarditis events postdose 1 of BNT162b2 and 36 postdose 2 [44, p.
-14].
-Assuming the same ratio of dose 1 and dose 2 events is reflected in the FDA’s projected 131 VAM/P cases for Pfizer’s BNT162b2 [8, Table 3], we find the corresponding estimate for VAM/P cases resulting from fully mRNA-1273 vaccinating 1 million males 18–25 to be
-[…]
-where we have applied the VSD-based comparison of VAM/P rates, mRNA-1273 vs.
-BNT162b2, discussed above.
-The first summand on the left of Equation (2.3.1) represents a VAM/P incidence of 35 cases per million 1st doses of mRNA-1273 among males 18–25, while the second represents approximately 260 cases per million 2nd doses.
-Using the FDA’s VAM/P case-hospitalization rate of 86% [9], we find the overall case rate of 295 per million full vaccinations corresponds to a rate of approximately 254 VAM/P hospitalizations per million full vaccinations and 224 per million 2nd doses.
-Thus, the FDA’s estimate of 110 VAM/P hospitalizations per million 2nd doses of mRNA-1273 (used in Scenarios 1–4) is less than half the estimate of 224 per million 2nd doses derived from the FDA’s BNT162b2 estimate and VSD data.
-$\bullet$ Internal FDA databases yielded VAM/P rates for mRNA-1273 substantially lower than those found in important studies.
-The FDA’s internal databases not only yielded VAM/P-rate estimates for mRNA-1273 inconsistent with the agency’s BNT162b2 estimates and VSD data but also the estimates are substantially lower than those of major international studies by Buchan et al.
+Using the rate ratios and BNT162b2 VAM/P-risk assessment discussed above, we compute in Section 6.1 of Supplement 1 a VAM/P hospitalization rate of 224 per million 2nd doses of mRNA-1273 for males 18–25 years old.
+The FDA’s BEST System also provided evidence that VAM/P risk for mRNA-1273 might be over twice that for BNT162b2.
+An FDA briefing document for a meeting held 10/14/2021 indicates that one of the agency’s BEST datapartners reported a postvaccination myo/pericarditis incidence rate of 283.7 (95% CI 145.2-573.5) cases per million full mRNA-1273 vaccinations (7 day risk window after each dose) [48, Section 2.3.3].
+Applying a background rate of 4 cases per million per week for 18–25-year-old males [Supplement 1, Section 6.3] to each dose, we find the VAM/P case rate corresponding to 283.7 is 275.7 per million full vaccinations.
+Assuming a case-hospitalization rate of 100%, as in FDA’s BNT162b assessment, we obtain 275.7 VAM/P hospitalizations per million full mRNA-1273 vaccinations vs.
+131 per million full BNT162b vaccinations.
+$\bullet$ The FDA’s meta-analysis of its BEST-System data [11, Section 2.3.1] yielded VAM/P-rate estimates substantially lower than those of major international studies by Buchan et al.
 [41] and Patone et al.
-[43].
-Furthermore, a U.S.
+[42], as well as a U.S.
 study by Sharff et al.
-[45] found, with a wide confidence interval, a myo/pericarditis incidence rate of 537.1 per million second doses of an mRNA-vaccine for males ages 18–24 years.
-The conclusion of [45] is noteworthy: [^10] The preprint [45] and the later publication [4] share the same conclusion.
-We identified additional valid cases of myopericarditis following an mRNA vaccination that would be missed by the VSD’s search algorithm, which depends on select hospital discharge diagnosis codes.
-The true incidence of myopericarditis is markedly higher than the incidence reported to US advisory committees.
-The VSD should validate its search algorithm to improve its sensitivity for myopericarditis.
-We now compare the FDA’s mRNA-1273 projected VAM/P case and hospitalization rates to those of the other sources discussed above.
-{adjustwidth}
--1.55cm
-Table 3. Estimated VAM/P Events Per Million 2nd Doses of mRNA-1273 Among Males 18–24 or 18–25 years old.* Based on Data Available Before 1 January 2022
+[43] [^5] The conclusion of [43] includes “We identified additional valid cases of myopericarditis following an mRNA vaccination that would be missed by the VSD’s search algorithm, which depends on select hospital discharge diagnosis codes.
+The true incidence of myopericarditis is markedly higher than the incidence reported to US advisory committees.” .
+In Table 3 below, we compare the FDA’s projected mRNA-1273 VAM/P case and hospitalization rates to those of the other sources discussed above.
+(We have excluded VAM/P-rate sources, such as [49], depending on strictly passive surveillance—see Section 6.8 of Supplement 1.)
+Table 3. Estimated VAM/P Events Per Million 2nd Doses of mRNA-1273 Among Males 18–24 or 18–25 years old.a Based on Data Available Before 1 January 2022
 Description
@@
 (Scenarios 5 & 6)
-[9, Table 2]
-Buchan et al.**
+[11, Table 2]
+Buchan et al.b
 [41]
@@
 (68 & 241)
-  302   269   525   260 301
-Hospitalizations***
+  302c   269   525   260 301
+Hospitalizationsd
 per million
@@
 21
-[44, p. 14]
+[40, p. 14]
 55
-[42, Table 4]
+[41, Table 4]
   23
 7
-[45, Table 1]
+[43, Table 1]
   22
@@
 107
-* For discussions of the data in columns 3–5, see Section 6 of Supplement 1; in particular, data in column 4 are estimates based on findings from [43] for males under 40, adjusted using an age-based myocarditis-incidence ratio as well as pericarditis data from [42].
-Data in column 6 is discussed above in the paragraph of Equation (2.3.1), with 22 being the number of Moderna myo/pericarditis cases contributing to the head-to-head comparison of [10, p.
-25].
-** All cases of myo/pericarditis contributing to Buchan et al.’s study were judged to meet the Brighton Collaboration’s case definitions (levels 1–3).
-Applying the Collaboration’s requirements reduced the reported 417 post-mRNA-vaccination study-period cases of myo/pericarditis to 297 [42, p.
+a For discussions of the data in columns 3–6, see Section 6 of Supplement 1; in particular, data in column 4 are estimates based on findings from [42] for males under 40, adjusted using an age-based myocarditis-incidence ratio as well as pericarditis data from [41].
+To obtain the column-5 VAM/P case rate of 525 per million we applied the background myo/pericarditis rate suggested by [41].
+b Buchan et al.
+applied the Brighton Collaboration’s case definition for myo/pericarditisx, which reduced the reported 417 post-mRNA-vaccination study-period cases of myo/pericarditis to 297 [41, p.
 5].
-*** For its BNT162b2 assessment, the FDA assumed all VAM/P cases were hospitalized [8, Section 2.3.2.2].
-For its mRNA-1273 assessment the FDA assumed VAM/P case-hospitalization rates decrease with age, starting with 86% for males 18–25 and ending with 77% for males 36 and over [9, Section 2.3.2.2].
-We apply the 86% rate in columns 4–6 (which, for column 5, matches the case hospitalization rate for the 7 males 18–24 having myo/pericarditis after dose 2 of an mRNA COVID vaccine: 6 of the 7 were hospitalized).
+c This estimated VAM/P-case rate is irrespective of the corresponding 1st-dose type.
+For homologous dosing (1st dose also mRNA-1273), Table 3 of [41] indicates the rate rises to at least 320 per million when the 2nd dose is taken less than 8 weeks after the 1st (with 4–6 weeks being the recommended interval [50]).
+d For its BNT162b2 assessment, the FDA assumed all VAM/P cases were hospitalized [10, Section 2.3.2.2].
+For its mRNA-1273 assessment the FDA assumed VAM/P case-hospitalization rates decrease with age, starting with 86% for males 18–25 and ending with 77% for males 36 and over [11, Section 2.3.2.2].
+We apply the 86% rate in columns 4–6 (which, for column 5, matches the case hospitalization rate: 6 of the 7 cases were hospitalized).
 For column 3, Buchan et al.
-provide an overall case-hospitalization rate associated with the 2nd dose of an mRNA vaccine (all ages, males & females) of 74.3% [42, Table 1]; we apply a case hospitalization rate of 80% for males 18–24, with 80% being approximately the average of 74.3% and the FDA rate of 86%
+report a case-hospitalization rate postdose 2 of an mRNA vaccine (all ages, males & females) of 79.2% ($\approx 164/207\times 100$%) [41, Table 1]; we expect the rate to be higher for the age range 18–24, but, conservatively, choose a hospitalization rate of 80%.
 $301\approx 302\cdot\frac{55}{107}.+269\cdot\frac{23}{107}+525\cdot\frac{7}{107}+260\cdot\frac{22}{107}$           $\ddagger$ $250\approx 242\cdot\frac{55}{107}.+231\cdot\frac{23}{107}+452\cdot\frac{7}{107}+224\cdot\frac{22}{107}$
 2.4. A Seventh Benefit Risk Scenario for mRNA-1273
-The FDA’s BLA Clinical Review Memorandum [46] for mRNA-1273, completed 1/28/2022, describes a benefit-risk scenario that does not match any of the six scenarios appearing in the FDA’s published benefit-risk assessment of mRNA-1273 vaccination [9].
-The BLA-Memo Scenario is described as follows in Section 4.7 of [46]:
-The results of the modeling assessment predicted that in the population of adults at highest risk of vaccine associated myocarditis/pericarditis (males ages 18–25 years), vaccination of 1 million individuals with two primary series doses of mRNA-1273 vaccine would prevent 76,362 COVID-19 cases, 1,755 hospitalizations, 421 ICU visits and 4 deaths due to COVID-19, but would cause 148 myocarditis/pericarditis cases resulting in 128 hospitalizations, 47 ICU visits, and no deaths.
-In Table 4 below, the BLA-Memo Scenario is compared to the FDA’s scenarios of [9, Table 2], in particular to Scenario 1, the “most likely scenario.” There are many noteworthy differences: e.g., dose-one VAM/P risk is included in the BLA-Memo Scenario but not in the scenarios of [9], and, in all scenarios of [9], the FDA predicts 0 ICU admissions owing to VAM/P while the BLA Memo predicts 1 million full vaccinations would cause 47.111111Subsequent analyses provide further evidence that the FDA underestimates VAM/P severity in [9]: A meta-analysis [47] of 23 studies including 854 patients ages 12–20 years found a 92.6% case-hospitalization rate for postvaccination myo/pericarditis with a mean length of stay (LOS) of 2.8 days and a 23.2% ICU-admission rate.
-Another pooled analysis [48] of 39 studies among 129 adults age 18 years or older (mean age 24) reported a median LOS of 3.0 days, 14% treated for heart failure, 8.1% admitted to the ICU, and 2 deaths (one after mRNA-1273 vaccination).
-A third meta-analysis [49] included 218 cases with a median LOS of 5.8 days with 3 deaths reported (one following mRNA-1273 vaccination), but provided no ICU-admission rate.
-{adjustwidth}
--0.3cm
-Table 4. FDA scenarios projecting benefit-risk outcomes per million 18–25-year-old males vaccinated with two primary series doses of Moderna RNA-1273.*
+The FDA’s BLA Clinical Review Memorandum [51] for mRNA-1273, completed 1/28/2022, describes a benefit-risk scenario not matching any scenario of the FDA’s published benefit-risk assessment of mRNA-1273 [11].
+The projections of the BLA-Memo Scenario [51, Section 4.7] are described in the initial row of data of Table 4 below and compared to the FDA’s scenarios of [11, Table 2] in particular to Scenario 1, the most likely scenario.
+Differences include the BLA-Memo Scenario’s accounting for 1st-dose VAM/P risk and projecting 1 million full vaccinations would cause 47 ICU admissions (vs.
+0 admissions).
+[^6] Subsequent analyses provide further evidence that the FDA may underestimate VAM/P severity in [11]; e.g., a meta-analysis [52] of 23 studies including 854 patients ages 12–20 years found a 92.6% case-hospitalization rate for postvaccination myo/pericarditis and a 23.2% ICU-admission rate.
+Table 4. FDA scenarios projecting benefit-risk outcomes per million 18–25-year-old males vaccinated with two primary series doses of mRNA-1273.*
 Benefits (Outcomes Prevented) VAM/P Risks
@@
 6 82,484 4,766 1,144 51 247 207 0 0
-* For Scenarios 1–6, the FDA estimated VAM/P risk only for dose 2.
-Outcomes for the BLA-Memo scenario are from Section 4.7 of [46].
-Outcomes for Scenarios 1–6 are from Table 2 of the FDA publication [9] with Scenario 1 being the “most likely scenario.”
+* For Scenarios 1–6, the FDA estimated VAM/P risk only for dose 2. Projections for the FDA’a Scenarios 1–6 are from [11, Table 2]. BLA-Memo-scenario projections are from [51, Section 4.7].”
 Consistent with the BLA Memo Scenario, the FDA’s 1/31/22 press release [12] announcing the full approval of the mRNA-1273 vaccine indicates some mRNA-1273 VAM/P cases do require ICU care.
-Both the BLA Memo [46, Section 4.7] and the FDA’s benefit-risk assessment of mRNA-1273 [9, Section 2.3.2.2] attribute their contradictory ICU risk estimates to VSD data.
-The BLA Memo lacks a VSD-data reference, while the FDA author team of [9] cite a source of VSD data [50] whose collection cutoff point is 21 August 2021, over four months before the team completed their mRNA-1273 benefit-risk analysis [9].
-The dated VSD data from [50] upon which the FDA [9] relies is scant, reporting only 21 postvaccination myo/pericarditis cases for those 18–29 (males and females, BNT162b2 and mRNA-1273, doses 1 and 2), with 18 of the cases ($\approx 86\%$) requiring hospitalization [50, p.
+Both the BLA Memo [51, Section 4.7] and the FDA’s mRNA-1273 benefit-risk assessment [11, Section 2.3.2.2] attribute their contradictory ICU risk estimates to VSD data.
+The BLA Memo lacks a specific reference, while the FDA in [11] cites a source of VSD data [53] with collection cutoff point 8/21/2021, over four months before completion of the FDA’s mRNA-1273 assessment.
+The VSD data from [53] cited by the FDA [11] is scant, reporting only 21 postvaccination myo/pericarditis cases for those 18–29 [53, p.
 16].
-For comparison, note that Buchan et al.
-report a total of 74 myo/pericarditis cases after the 2nd dose of mRNA-1273 for the age range 18–29 based on Ontario data from 6/1/21 to 9/4/21 [42, Table 4].
-Moreover, all these Ontario cases occurred during a 7-day risk-window postvaccination while the 21 cases reported in [50, p.
-16] occurred in a 21-day risk-window postvaccination.
-A comparison of Public Health Ontario’s myo/pericarditis data 6/1/21– 9/4/21 [42, Table 4] to VSD data through 10/9/21 [10, pp.
-8, 16] shows that, among males and females ages 18–39, the number of myo/pericarditis cases per million second doses of mRNA-1273 detected by Public Health Ontario is 3.76 times the number detected by the CDC’s VSD system—see Appendix 3 of Supplement 2.
-Because the BLA Memo’s quantification of the benefits and risks of mRNA-1273 vaccination does not closely relate to any of the scenarios of the FDA’s paper [9], it is unclear what role the analysis of [9] played in the FDA’s regulatory approval process for mRNA-1273.
-It is noteworthy that the FDA’s projection of no VAM/P-related ICU admissions for mRNA-1273 appears in a paper [9] submitted 5/19/22 well over three months after the appearance of FDA documents [12] and [46], with the former indicating ICU care is needed for some mRNA-1273 VAM/P cases and the latter projecting 47 VAM/P ICU admissions per million mRNA-1273 vaccinations of males 18–25.
-The FDA’s published benefit-risk assessment [8] of Pfizer’s BNT162b2 omits ICU-admissions data for VAM/P.
-Funk et al.
-state in the last paragraph of the introduction of [8], ”Vaccine attributable myocarditis/pericarditis cases, hospitalizations, intensive care unit (ICU) admissions, and deaths were selected to pair with benefit endpoints for benefit-risk comparison.” However, in Section 2.3.1, Funk et al.
-indicate there are three risk endpoints, omitting ICU-admissions.
-There were, in fact, myo/pericarditis-related ICU admissions following BNT162b2 vaccination.
-The FDA acknowledges this in its 8/23/21 press release [11] announcing the full approval of BNT162b2 (Comirnaty).
-Moreover, VSD data collected through 8/21/21 and presented at the CDC’s ACIP meeting, August 2021, indicates that for the 12–17 age group there were 19 hospitalizations for post-BNT162b2 myo/pericarditis and 36.8% (7 of the 19) hospitalizations involved an ICU admission [50, p.
-16].
-2.5. Summary
-In [9], the FDA substantially overestimated the hospitalizations-prevented benefit of mRNA-1273 vaccination for 18–25-year-old males by assuming COVID-hospitalization risk for this age group is equivalent to that for the 26–35 and 36–45 groups, by including incidental admissions, by ignoring prior-infection protection, and by assuming in its most likely scenario a COVID-19 case rate so high that the corresponding infection rate for the unvaccinated would be essentially 100% over 5 months.
-The FDA also substantially underestimated hospitalization risk for VAM/P, with available evidence indicating risk to be well over twice that of its most likely scenario—see Table 3 above.
-ICU-admission risk for VAM/P was underestimated as well (confirmed by FDA documents [46] and [12]).
+A comparison of Buchan et al.’s myo/pericarditis data from Ontario [41, Table 4, collected 6/1/21– 9/4/21] to VSD data through 10/9/21 [46, pp.
+8, 16] shows that, among males and females ages 18–39, the number of myo/pericarditis cases per million second doses of mRNA-1273 detected by Public Health Ontario is 3.76 times the number detected by the CDC’s VSD system [Supplement 2, Appendix 3].
 3. Reanalysis: Methods
-Recall that for its case and hospitalization endpoints, the FDA’s benefit-risk model consists of model inputs $C$ and $H$ (the number of COVID cases and hospitalizations in a test population of one million unvaccinated males) as well as VE and VEH (the fractions of these events prevented by full vaccination).
-These model inputs are combined to project the benefit of vaccinating the test population, and this projected benefit is compared to the model’s corresponding risk inputs, the number of VAM/P cases and hospitalizations projected for the 2nd mRNA-1273 dose.
-This comparison, or output of the model, depends on what values are used for each of the model inputs.
-The FDA performed the comparison under six different scenarios, each corresponding to a particular choice of input values.
-Because the FDA relied on CDC data to supply projections of values for $C$ and $H$, the FDA’s model does not explicitly link cases to hospitalizations and similarly doesn’t explicitly link VE and VEH.
-We extend and improve the FDA’s benefit-risk model for mRNA-1273 vaccination, focusing on the hospitalizations endpoint: those prevented by vaccination vs.
-those caused by VAM/P.
-Our model has the same structure as the FDA’s but has infections (not reported cases) as a key model input and links infections to hospitalizations via another key input, the infection hospitalization rate IHR.
-Additional inputs account for the fraction of the test population having a prior COVID infection as well as for the protection against infection and hospitalization provided by vaccination, by prior-infection, or by both (hybrid protection).
+Recall that for its case and hospitalization endpoints, the FDA’s benefit-risk model has inputs $C$ and $H$ (the numbers of evaluation-period COVID cases and hospitalizations, respectively, in a test population of one million unvaccinated males) as well as VE and VEH (the fractions of these events prevented by full vaccination).
+These model inputs are combined to project the cases-prevented and hospitalizations-prevented benefits of vaccinating the test population, and these benefits are compared to the model’s corresponding risk inputs, the numbers of VAM/P cases and hospitalizations projected for the 2nd mRNA-1273 dose.
+Our model, which has hospitalizations as its endpoint, is algebraic (like the FDA’s) and thus is widely accessible.
+Our model extends the FDA’s by not only incorporating prior-infection protection but also by being essentially equivalent to the FDA’s model if prior infection is assumed to provide no protection.
+Because the FDA relied on CDC data to supply projected values for $C$ and $H$, the FDA’s model does not explicitly link cases to hospitalizations and similarly doesn’t explicitly link VE and VEH.
+In contrast, our model does provide these links.
+For example, it links infections, a key model input replacing reported cases, to hospitalizations via another key input—the infection hospitalization rate IHR.
+Additional input variables are described in Section 3.1 below and summarized in Table 5.
 Model inputs are combined to project the hospitalizations-prevented benefit of vaccinating a test population, and this projected benefit is compared to the model’s risk input, the projected number of VAM/P hospitalizations expected to occur in the course of achieving a fully vaccinated test population.
@@
 Test population.
-Our model is based on a hypothetical population of 1 million males 18–25, so that our test population consists of those at highest risk of VAM/P (among the groups considered in the FDA’s mRNA-1273 assessment).
-We assume that the test population is representative, as of 1 January 2022, of all males in this age group in terms of: distribution of ages within the 18–25 range, comorbidities (e.g., obesity, cardiovascular disease, diabetes), and the percentage having had a prior COVID-19 infection.
+Our model is based on a hypothetical test population of 1 million males 18–25.
+We assume that the test population is representative, as of 1/1/2022, of all males in this age group in terms of: distribution of ages within the 18–25 range, comorbidities (e.g., obesity, cardiovascular disease, diabetes), and the percentage having had a prior COVID-19 infection.
 We project the number of hospitalizations for Omicron infection that will occur in this test population over the evaluation period 1/1/22 – 5/31/22 (a) assuming all members remain unvaccinated throughout the period, and (b) assuming all members of the population are vaccinated throughout the period (all fully vaccinated, 14 days past dose 2, on 1/1/22).
@@
 We call a test population consisting of infection naïve unvaccinated persons a baseline test population.
-We begin by introducing the inputs to our model.
 3.1. Model Inputs
 Infection rate $I_{r}$.
-Rather than base our projections of COVID prevalence on reported cases, we base our modeling on $I_{r}$, the fraction of a baseline test population of unvaccinated infection-naïve males 18–25 projected to become infected (symptomatic and asymptomatic) over the FDA’s 5-month evaluation period.
+We base our modeling on $I_{r}$, the fraction of a baseline test population of unvaccinated infection-naïve males 18–25 projected to become infected (symptomatic and asymptomatic) over the FDA’s 5-month evaluation period.
 Infection-hospitalization rate $H_{r}$.
@@
 $H_{r}$ is the fraction of a population of unvaccinated individuals experiencing a primary COVID infection that will progress to severe disease requiring hospitalization.
-Prior-infection fraction $F_{pi}$. This is the fraction of the test population having had a COVID-19 infection before the start of the 5-month evaluation period.
+We emphasize that $H_{r}$ is a sex-specific rate.
+The FDA provides data in Table 2 of [10] indicating that for the 18–24 age-range $H_{r}$ for males is significantly less that that for females.
+[^7] The data of Table 1 of [54] is consistent with that of [10, Table 2].
+Prior-infection fraction $F_{pi}$. This is the fraction of the test population having had a COVID-19 infection prior to the evaluation period.
 Vaccine effectiveness against infection $E_{v}$.
-This is the fraction by which the baseline infection rate $I_{r}$ is reduced due to the protection afforded by vaccination.
-That is, the infection rate for a fully vaccinated test population is $I_{r}(1-E_{v})$.
+This is the fraction by which the baseline infection rate $I_{r}$ is reduced due to the protection afforded by vaccination; i.e., the infection rate for a fully vaccinated test population is $I_{r}(1-E_{v})$.
 Prior-infection and hybrid effectiveness against infection, $E_{pi}$ and $E_{h}$.
@@
 VAM/P hospitalizations $H_{\rm VAM/P}$.
-The final input to our model is $H_{\rm VAM/P}$, which is the projected number of VAM/P hospitalizations expected to occur in the course of attaining a test population of 1 million fully mRNA-1273 vaccinated 18–25-year-old males.
+Our final model input $H_{\rm VAM/P}$ is the projected number of VAM/P hospitalizations expected to occur in attaining a test population of 1 million fully mRNA-1273 vaccinated 18–25-year-old males.
 To vaccinate fully 1 million individuals with an mRNA COVID-vaccine, more than 2 million doses must be administered (assuming, e.g., that those who experience myo/pericarditis related to dose 1 do not take dose 2), and the adverse-events associated with each of the over 2 million doses reflect the cost to the population of obtaining the benefits of 1 million full vaccinations.
+In Section 2.3, based on data available before 1/1/22, we obtained an estimated VAM/P hospitalization rate of 250 per million 2nd doses of mRNA-1273 given to 18–25-year-old males.
+In Section 6.7 of Supplement 1, we discuss VAM/P-hospitalization risk among 18–25-year-old males for dose 1 of mRNA-1273, justifying a choice of 18 VAM/P hospitalizations per million 1st doses, which is the 1st-dose rate of the FDA’s BLA-Memo Scenario—compare the first two rows of Table 4 above.
+Thus, for all our scenarios, we set $H_{\rm VAM/P}=268$.
+[^8] In Supplement 1 (Appendix 2, Section 2.4), we show that our VAM/P hospitalization-rate estimate of 268 per million full mRNA-1273 vaccinations among 18–25-year-old males is consistent with VAM/P data of three major studies [5, 7, 8] appearing in 2022 but available only after 1/22/22, showing the data suggests a VAM/P hospitalization rate between 240 and 277 per million full, homologous mRNA-1273 vaccinations among males 18–25.
 3.2. Combining Model Inputs to Produce Outputs
@@
 Let $V$ denote the corresponding number of expected hospitalizations assuming the population is vaccinated throughout the period.
-The difference $U-V$ is the hospitalizations-prevented benefit of vaccination for the test population, and the outputs $U$ and $V$ can be written in terms of the model inputs defined in the preceding section.
+The difference UVU-V is the hospitalizations-prevented benefit of vaccination for the test population, and the outputs $U$ and $V$ can be written in terms of the model inputs defined in the preceding section (thus, our model is deterministic).
 In our modeling, we consider only the possibility of hospitalizations associated with test-population members’ first infections during the evaluation period.
-We expect second infections over the course of a five-month period to be rare; moreover, infection-hospitalization risk will drop significantly after infection [16, 17, 18].121212 The meta-analyses [22, 23] suggest prior-infection provides greater protection against Omicron hospitalization than full vaccination among the COVID-naïve—see Section 7 of Supplement 1 for details.
+We expect second infections over a 5-month period to be rare; moreover, infection-hospitalization risk will drop significantly after infection [15, 16, 17].
 Our decision to analyze hospitalization risk associated with only first infections during the 5-month evaluation period allows us to assume, as did the FDA, that all effectiveness factors contributing to the analysis remain constant over the evaluation period.
+Each effectiveness value represents average effectiveness over the evaluation period; in this way, our modeling accounts for waning of vaccine and prior-infection protection.
 Deriving $U$.
@@
 As a first step, suppose that the entire test population is infection naïve.
-Then, in this baseline population, we expect $(1\ {\rm m})\cdot I_{r}$ infections to develop over 5 months, resulting in $(1\ {\rm m})\cdot I_{r}\cdot H_{r}$ hospitalizations, where $(1\ {\rm m})$ is an abbreviation for (1 million).
+Then, in this baseline population, we expect $(1\ {\rm m})\cdot I_{r}$ infections to develop over 5 months, resulting in $(1\ {\rm m})\cdot I_{r}\cdot H_{r}$ hospitalizations, where $1\ {\rm m}=$ 1 million.
 Next suppose a fraction $F_{pi}$ of the unvaccinated test population has had a COVID-19 infection prior to the start of the evaluation period.
@@
 Next we assume that every individual in the test population is fully vaccinated.
-Deriving the total number of Omicron hospitalizations $V$ for this population is analogous to deriving $U$, but more factors come into play because we must account for the protective effects of both vaccination and prior infection.
 Since all members of this test population are vaccinated, those with prior infection have hybrid protection.
 Once again we reason in two stages, first accounting for the protection against infection, then for the protection against hospitalization for infected individuals.
-The first stage depends on the model inputs $E_{v}$ and $E_{h}$, while the second stage uses $HRR_{v}$ and $HRR_{h}$.
 Once again, we assume a fraction $F_{pi}$ of the test population has had a prior COVID-19 infection.
-Then in the fully vaccinated test population, the expected number of hospitalizations is $(1-F_{pi})\cdot(1{\rm m})\cdot I_{r}\cdot(1-E_{v})\cdot H_{r}\cdot(1-HRR_{v})$ for those without prior infection and $F_{pi}\cdot(1{\rm m})\cdot I_{r}\cdot(1-E_{h})\cdot H_{r}\cdot(1-HRR_{h})$ for those with hybrid protection.
+Then, in the fully vaccinated test population, the expected number of hospitalizations is $(1-F_{pi})\cdot(1{\rm m})\cdot I_{r}\cdot(1-E_{v})\cdot H_{r}\cdot(1-HRR_{v})$ for those without prior infection and $F_{pi}\cdot(1{\rm m})\cdot I_{r}\cdot(1-E_{h})\cdot H_{r}\cdot(1-HRR_{h})$ for those with hybrid protection.
 So the total number of hospitalizations in the population is
 (3.2.2) $V=(1-F_{pi})(1{\rm m})I_{r}(1-E_{v})H_{r}(1-HRR_{v})+F_{pi}(1{\rm m})I_{r}(1-E_{h})H_{r}(1-HRR_{h}).$
-Hospitalizations prevented by vaccination.
-The difference between total Omicron hospitalizations in an unvaccinated test population and total hospitalizations in a fully vaccinated test population is the hospitalizations-prevented benefit of vaccination for the test population:
+Hospitalizations prevented by vaccination. As noted above, the difference UVU-V is the hospitalizations-prevented benefit of vaccination for the test population:
 […]
-We remark that disregarding the protection afforded by prior infection, as the FDA did, is equivalent to assuming in our model that all persons in the test population are infection-naïve, i.e.
+We remark that disregarding the protection afforded by prior infection, as the FDA did, is equivalent in our model to assuming that all persons in the test population are infection-naïve, i.e.
 that $F_{pi}=0$.
-Notice that if $F_{pi}=0$ in Equation (3.2.3), then its right-hand side reduces to
+With $F_{pi}=0$ in Equation (3.2.3), its right-hand side is
+(3.2.4) $\left[\text{Expected Baseline Hospitalizations}\right]\left(1-(1-E_{v})(1-HRR_{v})\right),$
 […]
-[…]
-Applying the model for benefits-risk assessment of vaccination.
-For the model to output a meaningful assessment of the benefits and risks of vaccination, we must make good estimates for the model inputs appearing in (3.2.3), as well as for the remaining model input: $H_{\rm VAM/P}$, which by itself describes the risk side of the assessment.
-Then the hospitalizations-prevented benefit $U-V$ can be compared to $H_{\rm VAM/P}$ to complete the analysis.
 3.3. Estimating Values for Model Inputs
-Like the FDA, we adopt a scenario-based structure for our benefit-risk assessment to account for uncertainties associated with pandemic dynamics as well as with the effectiveness of vaccine and prior-infection protection.
-Table 1 of Section 2 above summarizes the model-input assumptions the FDA used for each of the six scenarios of its mRNA-1273 assessment [9].
+Like the FDA, we account for uncertainties in model inputs by adopting a scenario-based assessment.
+The FDA’s model-input assumptions for its six scenarios of [11] are summarized in Table 1 of Section 2.
 Observe that the first three scenarios differ only in their COVID-incidence assumptions, with Scenario 1 representing the FDA’s most likely scenario.
-Our first three scenarios, Scenarios A–C, are reanalyses, respectively, of the FDA’ s Scenarios 1–3 in that our COVID incidence assumptions are derived from the FDA’s corresponding assumptions.
-Based on information available at the time of the FDA’s mRNA-1273 analysis, we consider the COVID incidence assumption of our Scenario D more likely to be accurate than those of Scenarios A-C; also, for Scenario D our values for other input variables represent our best estimates.
-Thus, we consider Scenario D to be our most likely scenario.
+Our first three scenarios, Scenarios A–C, are reanalyses, respectively, of the FDA’s Scenarios 1–3 because our COVID incidence assumptions are derived from the FDA’s corresponding assumptions.
+Based on information available at the time of the FDA’s mRNA-1273 analysis, we consider the COVID incidence assumption of our Scenario D more likely to be accurate than those of Scenarios A-C; also, for Scenario D our values for other inputs represent our best estimates.
+Thus, Scenario D is our most likely scenario.
 In Scenario E, our final scenario, we assume that prior infection and mRNA-1273 vaccination provide equivalent protection against Omicron, which, for reasons discussed in Section 2.2, we believe to be the minimum reasonable assumption of prior-infection’s protectiveness relative to vaccination’s protectiveness.
-Model inputs and outputs for our scenarios are described in Table 6 of the next section, while the methods we use to estimate values of model inputs are summarized below.
-Our estimates are based on data available to the FDA before 1/$2$
-Estimating $I_{r}$.
-As we noted in Section 2.2, the FDA’s COVID-case rate assumption for its most likely scenario, Scenario 1, is so large that it corresponds to the assumption that essentially all unprotected (i.e., unvaccinated) 18–25-year-old males become infected over the 5-month evaluation period.
-Thus, in Scenario A, our reanalysis of Scenario 1, we too assume all unprotected persons in the test population become infected.
-However, in our modeling, the unprotected are the infection-naïve unvaccinated.
-Assuming all such persons become infected over 5 months is equivalent to setting $I_{r}=1$, which is our Scenario-A value for $I_{r}$.
-As discussed in Section 1 of Supplement 1, for our Scenarios B and C, values of $I_{r}$ are determined by the FDA’s COVID-incidence assumptions for its Scenarios 2 and 3, respectively, while for our Scenarios D and E, $I_{r}$ is derived from the assumption that over the evaluation period the increase in COVID-infection level among 18–25-year-old males is twice that of the general population during the second COVID wave in the U.S.
-(10/1/20–2/28/21).
-Estimating $H_{r}$.
-Our estimate for $H_{r}$ is derived from an IHR model [51] developed via a seroprevalence-based meta-analysis anchored to the ancestral strain of SARS-CoV-2 and available as a preprint in October 2021 [52].
-Applying the model, we obtain a “with COVID” IHR of 0.37% for the infection-naïve unvaccinated of ages 18–25 and reduce this rate assuming 40% of hospitalizations attributed to COVID are incidental, arriving at an IHR of $0.22\%\approx 0.6\cdot 0.37\%$ (for infection-naïve unvaccinated 18–25 year-olds).
-The 0.22% rate for the general population of 18–25 year-olds in the U.S.
-(males and females) is consistent with U.S.
-military COVID-19 hospitalization data from the period 1 February 2020 through 4 January 2021 [53].
-See Supplement 1 for further details concerning our derivation of the IHR value $0.22\%$.
-In its benefit-risk assessments [8] and [9] for, respectively, Pfizer’s and Moderna’s mRNA vaccines, the FDA relied on sex-specific CDC hospitalizations data.
-The case and hospitalization rates per 100,000 persons in Table 2 of [8] suggests that for the 18–24 age range, females are 2.53 times as likely as males to be hospitalized with COVID-19 and 1.21 times as likely to have a case of COVID-19.131313Moreover, these relationships are consistent with those of [54, Table 1] based on data from 3/1/20 to 3/1/22.
-In Section 2 of Supplement 1, we use hospitalization and case data from Table 2 of [8] to estimate the IHR for unvaccinated, infection-naïve males 18–25 to be $0.14\%$ (and for females 18–25 to be $0.29\%$).
-Estimating $F_{pi}$.
-For our Scenarios A–C we use the value $F_{pi}=0.62$, which is the average of (a) the CDC’s estimate, mentioned earlier, that 54.9% of those 18–49 years old had been infected by COVID-19 before 10/1/21 [26] with (b) an estimate derived in Supplement 1 that 69% of unvaccinated male 18–25 year-olds had been infected by 1 January 2022.
-For Scenarios D and E, we use $F_{pi}=0.69$.
-These values of $F_{pi}$ may be underestimates because they rely on the CDC’s infected rate of 54.9% by 10/1/21 for the general population of 18–49 year-olds, and we would expect a somewhat higher infected rate among the unvaccinated.
-Estimating $E_{v}$, $HRR_{v}$.
-Recall that the FDA used VE = 0.30 and VEH = 0.72 for all its Omicron-based scenarios [9, Table 1].
-For reasons discussed in Supplement 1, for all our scenarios, we use $E_{v}=0.3$ and $HRR_{v}=0.67$.
-Hence, we are assuming a VEH of $1-(1-0.30)(1-0.67)\approx 0.77$, which exceeds the VEH used by FDA.
-We remark that the FDA’s choices of $VEH=0.72$ and $E_{v}=0.30$ are not consistent with the data and methods of the UK briefing [34] from which the FDA obtained its VEH estimate.
-Estimating $E_{pi}$ and $E_{h}$.
-Our values $E_{pi}=0.47$ and $E_{h}=0.57$ are derived (see Supplement 1) from a December 2021 study [27], based on data from South Africa.
-[^14] As discussed in Supplement 1, our effectiveness estimates involving prior infection are consistent with those of meta-analyses [22] and [23].
-[…]
-[…]
-{adjustwidth}
--1cm
-Table 5. Summary of Model Inputs
+Table 5 below summarizes model inputs with footnotes providing some comments on the estimation of their values—see Supplement 1 for details.
+Table 5. Summary of Input Variables*
 Variables for $U$ and $V$ Range of Values Scenario A Values
 $I_{r}$: Baseline-infection rate—fraction of an unvaccinated, infection-naïve test-population of 18–25-year-old males expected to contract an infection over the 5-month evaluation period.
-Infections need not accumulate at a constant rate $0\leq I_{r}\leq 1$ $1$
-$H_{r}$: Baseline infection-hospitalization rate for infection-naïve unvaccinated males 18–25 $0\leq H_{r}\leq 0.01$ $0.0014$
+Infections need not accumulate at a constant rate $0\leq I_{r}\leq 1$ 11
+$H_{r}$: Baseline infection-hospitalization rate for infection-naïve unvaccinated males 18–25 $\leq H_{r}\leq 0.01 0.0$0140.0014
 $F_{pi}$: Fraction of a test population of 1 million unvaccinated males 18–25 having had a COVID-19 infection before the start of the 5-month evaluation period $0\leq F_{pi}\leq 1$ 0.62
@@
 *Remarks.
-(i) All measures of effectiveness of protection against infection ($E_{pi}$, $E_{v}$, and $E_{h}$) are relative to infection-naïve, unvaccinated persons.
-(ii) All hospitalization-risk reductions ($HRR_{pi}$, $HRR_{v}$, and $HRR_{h}$) are relative to unvaccinated persons experiencing primary infections.
+(i) Estimates are based on data available before 1/22/22.
+(ii) $E_{pi}$, $E_{v}$, and $E_{h}$ are relative to infection-naïve, unvaccinated persons.
+(iii) $HRR_{pi}$, $HRR_{v}$, and $HRR_{h}$ are relative to unvaccinated persons experiencing primary infections.
+(iv) Our Scenario-A value $I_{r}=1$ reflects the assumption that all infection-naïve unvaccinated males 18–25 become infected over the evaluation period, consistent with FDA Scenario-1 assumptions (see Section 2.2).
+(v) We estimate $H_{r}$ using an IHR model [55] developed via a seroprevalence-based meta-analysis anchored to the ancestral strain of SARS-CoV-2 and available as a preprint in October 2021 [56].
+We assume a 40% incidental-hospitalization rate and make sex-based adjustments based on [10, Table 2].
+(See [Supplement 1, Section 2]) (vi) $F_{pi}=0.62$ is the average of (a) the CDC’s estimate that 54.9% of those 18–49 years old had been infected by COVID-19 before 10/1/21 [25] and (b) an estimate derived in Supplement 1 that 69% of unvaccinated male 18–25 year-olds had been infected by 1/1/2022.
+(vii) For reasons discussed in Supplement 1, for all scenarios, we use $E_{v}=0.3$ and $HRR_{v}=0.67$; thus, we are assuming a VEH of $1-(1-0.30)(1-0.67)\approx 0.77$, which exceeds the FDA’s VEH (viii) Estimates for $E_{pi},E_{h},HRR_{pi}$, and $HRR_{h}$, are based on data from [26] and [27].
+$HRR_{pi}=0.79$ and $HRR_{h}=0.88$ combine with $E_{pi}=0.47$ and $E_{h}=0.57$ to yield overall estimates for protection against hospitalization—88.9% for prior infection and 94% for hybrid protection—consistent with those of meta-analyses [21] and [22].
 4. Reanalysis: Results
-In Table 6 below, we compare the hospitalizations-prevented benefit of mRNA-1273 vaccination to corresponding VAM/P-hospitalizations risk through our Scenarios A–E, with Scenarios A–C being reanalyses of the corresponding FDA Scenarios 1–3.
-{adjustwidth}
--1.2cm
+In Table 6 below, we compare the hospitalizations-prevented benefit of mRNA-1273 vaccination to corresponding VAM/P-hospitalizations risk, with Scenarios A–C being reanalyses of the corresponding FDA Scenarios 1–3.
 Table 6.
-Scenarios projecting the hospitalizations-prevented benefit of mRNA-1273 vaccination of a hypothetical test population of 1 million males ages 18–25 compared to the estimated corresponding VAM/P hospitalization risk $H_{\rm VAM/P}$.
-For the FDA’s Scenarios 1–3, the values of $U$ and $V$ are determined by the FDA’s model described in Section 2.2, while, for our Scenarios A-E, the values of $U$ and $V$ determined .by our model described in Section 3.
-A green background for a scenario indicates benefit exceeds risk.
-Scenario Description Assumptions* $U$ $V$ Hospitalizations Prevented Per Million Vaccinations $U-V$ $H_{\rm VAM/P}$ Benefit Risk Ratio
-FDA
-Scenario 1
-Highest 2021 COVID in- cidence assumed. FDA’s “most likely scenario”
+Scenarios comparing the hospitalizations-prevented benefit of mRNA-1273 vaccination of a hypothetical test population of 1 million males 18–25 to the corresponding VAM/P hospitalization risk $H_{\rm VAM/P}$.
+For the FDA’s Scenarios 1–3, the values of $U$ and $V$ are determined by the FDA’s model described in Section 2.2, while, for Scenarios A-E, the values of $U$ and $V$ are determined .by our model described in Section 3.
+A green background indicates benefit exceeds risk.
+Scenario Description Assumptions* $U$ $V$ Hospitalizations Prevented Per Million Vaccinations UVU-V $H_{\rm VAM/P}$ Benefit Risk Ratio
+FDA Scenario 1 Highest 2021 COVID incidence assumed. FDA’s “most likely scenario”
 $U=6619$
 $V$ = 6619(1-0.72) $\approx$ 1853
-6619 1853 4766 110 (2nd-dose only) 43.33
-A Reanalysis of FDA’s Scenario 1 $I_{r}=1,H_{r}=0.0014,\\ F_{pi}=0.62,E_{pi}=0.47,\\ E_{v}=0.30,E_{h}=0.57,\\ HRR_{pi}=0.79,HRR_{v}=0.67,\\ HRR_{h}=0.88$ 629 168 461 268 1.72
-FDA
-Scenario 2
-Average 2021 COVID incidence assumed
-$U=2900$
-$V=2900(1-0.72)\approx 812$
-2900 812 2088 110 (2nd-dose only) 18.98
-B Reanalysis of FDA’s Scenario 2 $I_{r}=0.356,H_{r}=0.0014,\\ F_{pi}=0.62,E_{pi}=0.47,\\ E_{v}=0.30,E_{h}=0.57,\\ HRR_{pi}=0.79,HRR_{v}=0.67,\\ HRR_{h}=0.88$ 224 60 164 268 0.61
-FDA
-Scenario 3
-  Lowest incidence of pandemic assumed  (6/5/21) $U=882$ $V=882(1-0.72)\approx 247$ 882 247 635 110 (2nd-dose only) 5.77
-C Reanalysis of FDA’s Scenario 3 $I_{r}=0.052,H_{r}=0.0014,\\ F_{pi}=0.62,E_{pi}=0.47,\\ E_{v}=0.30,E_{h}=0.57,\\ HRR_{pi}=0.79,HRR_{v}=0.67,\\ HRR_{h}=0.88$ 33 9 24 268 0.09
-D COVID incidence   twice that for the 2nd COVID wave in the U.S.
-Our “most likely scenario” $I_{r}=0.456,H_{r}=0.0014,\\ F_{pi}=0.69,E_{pi}=0.47,\\ E_{v}=0.30,E_{h}=0.57,\\ HRR_{pi}=0.79,HRR_{v}=0.67,\\ HRR_{h}=0.88$ 247 68 179 268 0.67
+6619 1853 4766 110 (Dose 2 only) 43.33
+[…]
+FDA Scenario 2 Average 2021 COVID incidence assumed $U$ = 2900 VV= 2900(1-0.72) $\approx$ 812 2900 812 2088 110 (Dose 2 only) 18.98
+[…]
+FDA Scenario 3 Lowest 2021 COVID incidence assumed $U$ = 882 VV= 882(1-0.72) $\approx$ 247 882 247 635 110 (Dose 2 only) 5.77
+[…]
+[…]
 E Equivalent protection for prior infected and COVID-naïve vaccinated; incidence as in Scenario D.
-$I_{r}=0.456,H_{r}=0.0014,\\ F_{pi}=0.69,E_{pi}=0.30,\\ E_{v}=0.30,E_{h}=0.57,\\ HRR_{pi}=0.67,HRR_{v}=0.67,\\ HRR_{h}=0.88$ 300 68 232 268 0.87
-* For discussions of how “Assumptions” were derived, see Section 3.3 as well as Supplement 1.
-To illustrate the application of our model, we carry out our Scenario-A computation. Using Scenario A assumptions for variable values and our definition (3.2.1) of the function $U$, we obtain
-$U=\left(1-0.62\right)\cdot\left(1{\rm m}\right)\cdot 1\cdot 0.0014+0.62\cdot\left(1{\rm m}\right)\cdot 1\cdot\left(1-0.47\right)\cdot 0.0014\cdot\left(1-0.79\right)\approx 629,\$ and, from our definition (3.2.2) of the function $V$,
-$V=\left(1-0.62\right)\cdot\left(1{\rm m}\right)\cdot 1\cdot\left(1-0.30\right)\cdot 0.0014\cdot\left(1-0.67\right)+0.62\cdot 1\cdot\left(1-0.57\right)\cdot 0.0014\cdot\left(1-0.88\right)\approx 168.$
-[…]
-Observe that our Scenario-D projection of 179 COVID-hospital admissions prevented by vaccination corresponds to $43\approx 0.24\cdot 179$ ICU admissions prevented [9, Table 3], which should be compared to the 47 ICU admissions for VAM/P projected by the BLA-Memo Scenario and 0 such admissions projected in the scenarios of [9].
-In Supplement 2, we carry out sensitivity analyses exploring how the benefit-risk picture for mRNA-1273 vaccination changes with modified assumptions for key variables.
-We also show that in Scenarios B–E, when hospitalizations prevented by vaccination include those prevented by dose 1 during the period preceding the point of full vaccination, vaccine risks still outweigh benefits in terms of the hospitalizations endpoint.
+$I_{r}=0.456$, $H_{r}=0.0014$, $F_{pi}=0.69$, $E_{pi}=0.30$, $E_{v}=0.30$, $E_{h}=0.57$, $HRR_{pi}$$HRR_{v}=0.67$, $HRR_{h}=0.88$ 300 68 232 268 0.87
+* For discussions of “Assumptions,” see Section 3.3 as well as Supplement 1.
+We illustrate our model’s application by carrying out Scenario-A computations, Using Scenario A assumptions from Table 6 and our definitions (3.2.1) and (3.2.2) of $U$ and $V$, we obtain
+$U=\left(1-0.62\right)\cdot\left(1{\rm m}\right)\cdot 1\cdot 0.0014+0.62\cdot\left(1{\rm m}\right)\cdot 1\cdot\left(1-0.47\right)\cdot 0.0014\cdot\left(1-0.79\right)\approx 629,{\rm and}\\$ $V=\left(1-0.62\right)\cdot\left(1{\rm m}\right)\cdot 1\cdot\left(1-0.30\right)\cdot 0.0014\cdot\left(1-0.67\right)+0.62\cdot 1\cdot\left(1-0.57\right)\cdot 0.0014\cdot\left(1-0.88\right)\approx 168.$
+Thus, for Scenario A, the hospitalizations-prevented benefit of vaccination is projected to be UV= 629168=461U-V=\ 629-168=461 per million full mRNA-1273 vaccinations, roughly one-tenth of the FDA’s Scenario-1 estimate 4,766 [11].
+Although Scenario A has a benefit-risk ratio above 1, its assumption that $I_{r}=1$ makes this scenario highly unlikely.
+[^9] Based on seroprevalence and modeling information, we estimate $I_{r}\leq 0.5$ [Supplement 2, Section 2.3].
+For the more likely scenarios B, D, and E, our modeling suggests that mRNA-1273 vaccination of 18–25-year-old males generated between 16% and 63% more hospitalizations from VAM/P alone compared to COVID hospitalizations prevented.
+Comparing the projections of FDA’s “most likely scenario” to ours, we find that the FDA’s benefit-risk ratio, 43.33, is over 60 times more favorable than ours, 0.67.
+Though Scenario D is our most likely scenario, Scenario E is equally important, being a “best-case scenario” for vaccination incorporating the minimum reasonable assumption of prior-infection protection vs.
+vaccine protection.
+In Section 2 of Supplement 2, we carry out sensitivity analyses exploring how modifying model inputs impacts the benefit-risk picture.
+For example, if we fix our model-input values, except for $I_{r}$, at their Scenario-D levels, then benefit-risk equipoise is reached when $I_{r}\approx 0.685$.
+We also illustrate how slider-plots can provide insights concerning equipoise questions involving variation of more than one model input.
 4.1. Subpopulation Analyses Based on Prior-Infection Status
-The FDA acknowledges, “The health condition of individuals may remarkably impact the B-R profile and its evaluation” [9, Section 4]; however, the authors of [9] do not explore this remarkable impact, explaining “[W]e did not investigate the benefits and risks for specific subpopulations, such as those with comorbidities, due to limited information for these populations.” Comorbidities such as obesity may triple COVID-hospitalization risk [55].
-In this subsection, we consider subpopulations for which there was sufficient information before January 2022 to conduct conclusive benefit-risk assessments.
-We consider, in particular, subpopulations having had a COVID infection prior to the evaluation period 1/1/22 – 5/31/22.
-Prior infection not only decreases the hospitalization-prevented benefit of vaccination but also increases the risk of a severe adverse event after vaccination: see, e.g., [56, 57], with [56] reporting prior infection is “associated with an increased risk of severe side effects leading to hospital care (1.56 [risk ratio] (95%CI: 1.14–2.12)).” Such “severe side effects” possibly linked to mRNA COVID vaccination and potentially requiring hospitalization include anaphylaxis [58], Multisystem Inflammatory Syndrome [59, 60], Guillain Barré Syndrome [61], posterior orthostatic tachycardia syndrome (POTS) [62], retinal vascular occlusion [63], and cerebral venous sinus thrombosis (CVST) [64].
-Moreover, there is evidence that prior infection increases the risk of vaccine-associated myocarditis [43]—see Appendix 4 of our Supplement 2 for details.
-[^19] Survey results reported in [65] also suggest that prior infection may increase the risk of VAM/P; e.g., in Table 4, we find that 2.53% of those having a “laboratory-confirmed SARS-CoV-2 infection” before receiving dose 1 of mRNA-1273 reported “chest tightness/pain” during the 7-day period following the dose while only 0.54% of those not having a confirmed infection before dose 1 reported these symptoms; corresponding percentages for dose 2 (Table 5) are 1.78% and 0.98%.
-Prior infection’s reducing the hospitalizations-prevented benefit of vaccination while increasing postvaccination adverse-event risk creates a compelling case for conducting a benefit-risk assessment for the subpopulation of unvaccinated 18–25-year-old males having prior-infection protection.
-Consider the subpopulation of 18–25-year-old unvaccinated males having a COVID-infection prior to the evaluation period 1/1/22 – 5/31/22.
-Using our model, we can compute the expected number of COVID hospitalizations for this subpopulation in various scenarios simply by setting $F_{pi}=1$ and leaving all other variables unchanged.
-For example, with $F_{pi}=1$ in Scenario A, our model predicts $U\approx 155.82$ hospitalizations per million and $V\approx 72.24$, yielding a hospitalizations-prevented benefit of 85.8 per million vaccinated, far below VAM/P hospitalization risk.
-The analysis of the preceding paragraph may be reinterpreted in terms of evaluation-period hospitalization risk for a typical unvaccinated 18–25-year-old male, COVID-infected prior to the evaluation period.
-For ease of exposition, we will call this male Bob.
-[^20] If we were also considering COVID-hospitalization risk for a typical 18–25-year-old female, we would call her Alice, choosing “Alice” and “Bob” because these names are customarily used for “person A” and “person B” in the field of cryptography.
-If Bob chooses not to vaccinate, our model estimates the probability that Bob will be hospitalized for COVID during the evaluation period to be 155.82 in 1 million (assuming $I_{r}=1$).
-Bob’s total hospitalization risk associated with the choice to vaccinate is essentially the sum of his probability of VAM/P hospitalization and the probability of his having a breakthrough-case COVID hospitalization during the evaluation period (assuming postvaccination VAM/P hospitalization is independent of postvaccination COVID-hospitalization).212121 Technically speaking, because there is a small chance—on the order of 0.02 in 1 million—that Bob might experience both a VAM/P hospitalization and a breakthrough-COVID hospitalization, Bob’s total hospitalization risk from vaccination is the sum of his VAM/P-related risk and his COVID-related risk minus the product of these two risks (with the product being the small chance that Bob experiences both a VAM/P hospitalization and a COVID hospitalization by our assumption of independence of these events).
-While there is some evidence that VAM/P risk increases with prior infection, we assume only that VAM/P risk does not decrease.
-Suppose Bob decides to take dose 1 and at the same time decides to take dose 2 if VAM/P does not occur after dose 1; then, if $r_{1}$, $r_{2}$ are, respectively, the VAM/P-incidence rates for doses 1 and 2 for males 18–25, then the probability Bob will experience VAM/P is well approximated by $r_{1}+(1-r_{1})r_{2}.$ Because $r_{1}$ is quite small, the preceding probability is essentially $r_{1}+r_{2}$, which explains why, even on the individual level, the FDA should consider the sum of risks, postdose 1 and postdose 2 in assessing the VAM/P risk of full vaccination (as the agency did in its population-level analyses of BNT162b2 [8, Section 2.3.1]) and of mRNA-1273 in its BLA-Memo [46, Section 4.7]).
-Therefore 268/1,000,000 approximates the probability that Bob experiences hospitalization owing to VAM/P during the course of mRNA-1273 vaccination.
-The probability that he experiences COVID hospitalization owing to a breakthrough infection is 72.24 in 1 million, according to our model.
-Thus, his total hospitalization risk from vaccination is approximately 340.24 in 1 million.
-This is under the Scenario-A assumption that $I_{r}=1$, which we interpret to mean that evaluation-period infection is certain for an infection-naïve male 18–25, and Bob’s probability of infection is $I_{r}(1-E_{pi})=0.53$.
-The following Table shows how Bob’s hospitalization risk varies as his risk of infection varies.
-{adjustwidth}
--1.2cm
+There is a compelling case for conducting a benefit-risk assessment for the subpopulation of unvaccinated 18–25-year-old males having prior-infection protection because prior infection not only decreases the hospitalizations-prevented benefit of vaccination but also increases the risk of severe adverse events101010e.g., those described in [57, 58, 59, 60, 61, 62, 63] after vaccination [64, 65], with [64] reporting prior infection is “associated with an increased risk of severe side effects leading to hospital care (1.56 [risk ratio] (95%CI: 1.14–2.12)).” Moreover, prior infection may increase the risk of VAM/P [42]—see Appendix 4 of Supplement 2 for details.
+[^11] Survey results reported in [66] also suggest that prior infection may increase the risk of VAM/P; e.g., in Table 4, we see that 2.53% of those having a confirmed SARS-CoV-2 infection before dose 1 of mRNA-1273 reported “chest tightness/pain” during the 7-day period after the dose while only 0.54% of those not having a confirmed infection before dose 1 reported these symptoms; corresponding percentages for dose 2 (Table 5) are 1.78% and 0.98%.
+For the subpopulation of 18–25-year-old unvaccinated males having a COVID-infection prior to the evaluation period 1/1/22 – 5/31/22, we can use our model to compute the expected number of COVID hospitalizations in various scenarios simply by setting $F_{pi}=1$ while leaving all other inputs unchanged.
+For example, with $F_{pi}=1$ in Scenario A, our model predicts $U\approx 155.82$ hospitalizations per million and $V\approx 72.24$, yielding a hospitalizations-prevented benefit of 83.58 per million vaccinated, far below VAM/P hospitalization risk ($\approx$ 268 per million vaccinated).
+As we discuss in Appendix 3 of Supplement 1, the preceding paragraph’s analysis may be reinterpreted in terms of evaluation-period hospitalization risk for a typical unvaccinated 18–25-year-old male, COVID-infected prior to the evaluation period.
+For such a male, the total hospitalization risk of the choice to vaccinate is essentially the sum of his probability of VAM/P hospitalization and his probability of having a breakthrough-case COVID hospitalization during the evaluation period.
+Our analysis is summarized in Tables 7 and 8 below in terms of hospitalization-risk ratios, with ratios <1<1 indicating the choice to vaccinate decreases risk and >1>1 indicating it increases risk.
 Table 7.
@@
 VAM/P Risk 268 in 1 million
-$1$ 155.82 in 1 million 72.24 in 1 million $2.18\approx\frac{268+72.24}{155.82}$
+11 155.82 in 1 million 72.24 in 1 million $2.18\approx\frac{268+72.24}{155.82}$
 $0.5$ 77.91 in 1 million 36.12 in 1 million $3.90\approx\frac{268+36.12}{77.91}$
@@
 *COVID hospitalization risk is considered over the 5-month evaluation period
-**$I_{r}(1-E_{pi})$ is the probability the “typical male” (18–25 with prior-infection protection) will become COVID infected over the evaluation period if unvaccinated; e.g., 0.53 if $I_{r}=1$
-In Section 2.3 of Supplement 2, based on seroprevalence and modeling information, we estimate that $I_{r}\approx 0.5$ (and is likely $<0.5$), that is, roughly 1/2 of unvaccinated, infection-naïve 18–25-year-old males in the U.S.
-became infected over the evaluation period (and Bob’s infection risk is $0.2651/2(1-E_{pi})=$ 0.265 because of his prior infection).
-Thus, according to Table 7, our modeling suggests that for our typical unvaccinated 18–25-year-old male Bob with prior-infection, the choice to vaccinate creates a hospitalization risk approximately 4 times that associated with the choice not to vaccinate.
-Moreover, if we account for the increased likelihood of postvaccination adverse-event hospitalizations for those with prior infection, then the risk ratios of Table 7 grow even larger.
-Also note that for every value of $I_{r}$ in Table 7, hospitalization risk owing to VAM/P alone exceeds COVID-hospitalization risk.
-Applying the information in Table 7 is difficult because a “typical male” has an “average comorbidities level” and the meaning of this is unclear.
-However, if our typical unvaccinated male Bob with prior-infection protection also has no comorbidities, then his hospitalization-risk ratios should be larger than those of Table 7.
-(There is no evidence that being free of comorbidities decreases VAM/P risk; e.g., according to a 7/16/21 report of the chief science advisor of Canada “Patients presenting with this complication [vaccine-associated myo/pericarditis] typically have been healthy individuals.” [66].
-[^22] See also [67, p.
-383].)
-Now assume that Bob is an unvaccinated 18–25-year-old male with prior infection without comorbidities.
-To establish conclusively that Bob’s hospitalization risk is greater with vaccination, we consider two bounding analyses, each overestimating the benefits of vaccination and underestimating the adverse-event risks associated with vaccination.
-{adjustwidth}
-0.3in0.3in Scenario I: To underestimate hospitalization risk associated with vaccination, we assume that vaccination is 100% effective in preventing breakthrough-infection hospitalization, and we obtain a lower bound on Bob’s VAM/P-hospitalization risk, by ignoring dose 1 VAM/P hospitalizations and using our estimate of 242 VAM/P hospitalizations per million second doses of mRNA-1273 among males 18–25 based on Ontario data (as described in Table 3).
-Thus, ignoring his first-dose VAM/P hospitalization risk, his hospitalization risk for postvaccination adverse events other than myo/pericarditis, and his likely increased risk for adverse-event hospitalization owing to his prior infection, we obtain an underestimate of Bob’s postvaccination hospitalization risk of 242 in 1 million (based on VAM/P alone).
-To overestimate Bob’s COVID-hospitalization risk assuming he remains unvaccinated, we assume that evaluation-period infection is certain for an infection-naïve unvaccinated male ($I_{r}=1$) and that Bob’s being free of comorbidities does not decrease his COVID hospitalization risk which is 155.82 in 1 million according to Table 7.
-We conclude that Bob is more likely—at least $1.55\approx\frac{242}{155.82}$ times as likely—to be hospitalized for VAM/P after choosing to vaccinate than to be hospitalized for COVID while remaining unvaccinated.
-{adjustwidth}
-0.3in0.3in Scenario II: If we assume no COVID-hospitalization risk-reduction from Bob’s being free of comorbidities, assume his VAM/P-hospitalization risk is only 110 in 1 million (the FDA’s assumption in its most likely scenario of [9]), ignore hospitalization risks for other postvaccination adverse events (likely increased owing to Bob’s prior infection), and assume every infection-naïve unvaccinated male becomes infected over the evaluation period, then we still conclude that Bob is more likely—at least $1.17\approx\frac{110+72.24}{155.82}$ times as likely—to be hospitalized for VAM/P or a breakthrough infection after choosing to vaccinate than to be hospitalized for COVID while remaining unvaccinated.
-Note that if we use the VAM/P hospitalization risk estimate of the FDA’s BLA-Memo Scenario the ratio 1.17 rises to $1.29\approx\frac{128+72.24}{155.82}$.
-To further elucidate the important role that prior infection plays in a careful benefit-risk assessment of mRNA-1273 vaccination, we note that for an 18–25-year-old male without prior COVID infection, hospitalization-risk ratios become favorable for vaccination:
-{adjustwidth}
--1.4cm0cm
+$I_{r}(1-E_{pi})$ is the probability the “typical male” (18–25 with prior-infection protection) will become COVID infected over the evaluation period if unvaccinated; e.g., 0.53 if $I_{r}=1$
+According to Table 7, our modeling suggests that for a typical unvaccinated 18–25-year-old male with prior-infection, the choice to vaccinate created a hospitalization risk over twice that associated with the choice not to vaccinate even if $I_{r}=1$ (evaluation-period infection is certain for an infection-naïve male 18–25).
+Because a “typical male” has an “average comorbidities level,” the information in Table 7 is difficult to apply.
+However, if our typical unvaccinated male of Table 7 also has no comorbidities, then his hospitalization-risk ratios should exceed those of Table 7, [^12] See, e.g., the BMI 23–24 column of Table 9 because risk-ratio denominators should decrease given no comorbidities and there is no evidence that being free of comorbidities reduces VAM/P risk [67].
+[^13] See also [68, p.
+383].
+For an 18–25-year-old male without prior COVID infection, hospitalization-risk ratios become favorable for vaccination:
 Table 8.
-Hospitalization-risk ratios for a typical male 18—25 without prior-infection protection: risk of hospitalization after choosing mRNA-1273 vaccination (VAM/P or severe COVID-breakthrough infection*) divided by risk of COVID hospitalization after choosing not to vaccinate.
+Hospitalization-risk ratios for a typical male 18—25 without prior-infection protection: risk of hospitalization after choosing mRNA-1273 vaccination (VAM/P or severe COVID-breakthrough infection) divided by risk of COVID hospitalization after choosing not to vaccinate.
 Probability
@@
 VAM/P Risk 268 in 1 million
-$1$ 1400 in 1 million 323.4 in 1 million $0.42\approx\frac{268+323.4}{1400}$
+11 1400 in 1 million 323.4 in 1 million $0.42\approx\frac{268+323.4}{1400}$
 $0.5$ 700 in 1 million 161.7 in 1 million $0.61\approx\frac{268+161.7}{700}$
 $0.25$ 350 in 1 million 80.85 in 1 million $1.00\approx\frac{268+80.85}{350}$
-*COVID hospitalization risk is considered over a 5-month evaluation period
 4.2. Subpopulation Analyses Based on Prior-Infection Status and BMI
-For the remainder of the paper, we no longer confine to footnotes discussions and citations of data unavailable at the time of the FDA’s benefit-risk analysis of mRNA-1273.
-We illustrate how comorbidities may impact a benefit-risk assessment, focusing on obesity because evidence suggests that Body Mass Index (BMI) is the single most potent modifier of COVID-19 risk for those under 60 years old [68, 69].
-According to the CDC [55], “Having obesity [BMI $\geq 30$] may triple the risk of hospitalization due to a COVID-19 infection.” This is consistent with a finding that hospitalization risk for the obese is 3.2 times that for the non-obese for an adolescent study population (ages 12–17) [70, Table IV].
-Gao et al.
-[71] provide additional support for this CDC estimate based on a cohort study of 6,910,695 residents of England (data from the period 1/24/20 – 4/30/20).
-Gao et al.
-found that for those 20–39 years old COVID hospitalization risk is minimized when BMI $\approx 23$ and that for $23\leq\text{BMI}\leq 44$ risk is multiplied by a factor of approximately 1.09 (1.08, 1.10) for each unit increase in BMI.
-Gao et al.
-“found no difference in effect estimates by sex” [71, Figure 2].
-Applying Gao’s model (see Appendix 5 of Supplement 2), we obtain the following:
-Table 9. Estimated COVID hospitalization hazard ratio (HR) for 20–39 year-olds relative to those with BMI $\approx 23$ based on a model by Gao et al. [71]
-BMI Range COVID Hospitalization HR
-23–24 ref.
-25–29 1.49
-30–34 2.29
-35–39 3.52
-40–44 5.41
-Assuming that for 18–25-year-old males obesity triples COVID-hospitalization risk and that 20.5% of this population is obese [72], we find not being obese drops baseline COVID hospitalization risk $H_{r}=0.0014$ to approximately $0.71\cdot H_{r}$ (see Appendix 5 of Supplement 2).
-We assume this same factor determines the drop in hospitalization risk for unvaccinated males 18–25 with prior infection (IHR $=0.71\cdot H_{r}\cdot(1-E_{pi})(1-HRR_{pi})$).
-We assume that not being obese for an otherwise typical vaccinated 18–25-year-old male decreases IHR by a greater amount: IHR $=0.36\cdot H_{r}(1-E_{v})(1-HRR_{v})$ based on case-hospitalization rates computed from the raw data of Supplementary Table A2 of [73].
-We provide below a COVID hospitalization benefit-risk analysis derived from Gao et al.’s BMI-based model [71] described above.
-We assume that the BMI range 23–24 for 18–25-year-old males provides minimum COVID-hospitalization risk, which we estimate by multiplying the hospitalization risks appearing in column 2 of Tables 7 and 8 by the factor 0.71 and those in column 3 by the factor of 0.36.
-For a BMI-based hazard ratio HR from Table 9, we compute the corresponding effect on the hospitalization risks in column 2 of Tables 7 and 8 via the factor $\text{HR}\cdot 0.71$ and for those in column 3 via the factor $\text{HR}\cdot 0.36$.
+At this point, we discontinue confining to footnotes discussions and citations of data unavailable at the time of the FDA’s benefit-risk analysis of mRNA-1273.
+The FDA acknowledges, “The health condition of individuals may remarkably impact the B-R profile and its evaluation” [11, Section 4].
+Comorbidities such as obesity may triple COVID-hospitalization risk [69].
+Our model can produce a benefit-risk analysis based on comorbidity status by appropriately adjusting its infection-hospitalization-rate input $H_{r}$.
+We illustrate the process in Appendix 3 of Supplement 1, focusing on obesity because evidence suggests that Body Mass Index (BMI) is the single most potent modifier of COVID-19 risk for those under 60 years old [70, 71].
+We use an assessment of COVID-hospitalization risk based on BMI that finds risk is minimized when BMI $\approx 23$ and that for $23\leq\text{BMI}\leq 44$ risk is multiplied by a factor of approximately 1.09 (1.08, 1.10) for each unit increase in BMI [72].
+The results of our modeling appear in Table 9 below.
 Table 9.
-Hospitalization-risk ratios (risk if vaccinated (VAM/P & breakthrough)risk if $\left(\frac{\text{risk if vaccinated (VAM/P \& breakthrough)}}{\text{risk if unvaccinated}}\right)$ based on BMI for males 18—25.
+Hospitalization-risk ratios $\left(\frac{\text{risk if vaccinated (VAM/P \& breakthrough)}}{\text{risk if unvaccinated}}\right)$ for males 18–25 based on BMI .
 Ratios displayed as pairs: “with prior-infection protection, without prior-infection projection.”
@@
 BMI 23–24 BMI 25–29 BMI 30–34 BMI 35–39 BMI 40–44
-$1$ 2.66, 0.39 1.86, 0.30 1.29 , 0.23 0.92, 0.19 0.68, 0.17
-$0.5$ 5.08, 0.66 3.49, 0.48 ** 2.35, 0.35 1.61, 0.27 1.13, 0.22
-$0.25$ 9.92 ,1.20 6.74 , 0.84 4.47 , 0.59 2.99, 0.42 2,03, 0.32
+11 2.66, 0.39 1.86, 0.30 1.29 , 0.23 0.92, 0.19 0.68, 0.17
+$0.5$ 5.08, 0.66 3.49, 0.48 2.35, 0.35 1.61, 0.27 1.13, 0.22
+$0.25$ 9.92 ,1.20 6.74 , 0.84 4.47 , 0.59 2.99, 0.42 2.03, 0.32
 *With prior-infection protection, $I_{r}(1-E_{pi})$ is the probability of infection over the evaluation period.
-**Sample computation: $\frac{268+36.12\cdot 1.49\cdot 0.36}{77.91\cdot 1.49\cdot 0.71},\frac{268+161.7\cdot 1.49\cdot 0.36}{700\cdot 1.49\cdot 0.71}\approx 3.49,0.48$
-Consistent with Table 8, the preceding table suggests that for 18–25-year-old males without prior-infection protection, vaccination reduces hospitalization risk for all BMI ranges except for the ideal range (BMI 23–24) in a low-infection-risk scenario.
-Similarly, consistent with Table 7, we see that for 18–25-year-old males with prior-infection protection, vaccination increases hospitalization risk except for the those with very high BMI (35–44) in a high-infection-risk scenario.
-In a medium infection-risk scenario ($I_{r}=0.5$), our modeling suggests that for the non-obese (BMI $<30$) having prior-infection protection the choice to vaccinate creates a hospitalization risk 3.49 to 5.08 times that associated with the choice not to vaccinate.
-5. Validating the Model
-In this section, we discuss our model-validation work of Supplement 2, where we test our model as well as the FDA’s using COVID-hospitalizations data from Ontario over the evaluation period 1/1/22 – 5/31/22 and show that our model produces projections of the hospitalizations-prevented benefit of vaccination that are much more accurate than those of the FDA [9].
-We test our model in two ways: (i) showing that it fits Ontario COVID-hospitalization data with inputs at plausible levels for Ontario over the evaluation period (see Section 8 of Supplement 2), and (ii) using Ontario COVID-hospitalization data to predict corresponding data for the U.S.
-and then comparing those predictions to ours (see Section 7 of Supplement 2).
-We validate our model using Ontario data for many reasons; e.g., unlike the CDC, Public Health Ontario chose to publish incidental COVID-hospitalization rates (starting 1/8/2022).
-We discuss other reasons in Section 3 of Supplement 2.
-Because Public Health Ontario does not provide sex-specific COVID-hospitalizations data, to test our model using its data, we compute model outputs assuming no sex-related differences in COVID infection and hospitalization risk, so that $H_{r}=0.0022$, and, for Scenario D, $F_{pi}=0.76$ (see Supplement 1, Sections 2 and 3).
-If $H_{r}$ and $F_{pi}$ are increased to $0.0022$ and $0.76$ in Scenario D, leaving all other variables unchanged, then 231 is the hospitalizations-prevented benefit per million vaccinations.
-Based on COVID hospitalization counts provided by Public Health Ontario as well as corresponding incidental hospitalization rates, we estimate that vaccinating one million 18–25-year-old Ontarians prevented 152 COVID hospitalizations in Ontario over the evaluation period—see Table S2.5.
-Because seroprevalence studies suggest that over the evaluation period, the COVID-infection rate in Ontario exceeded that in the U.S., we would anticipate a hospitalizations-prevented benefit of less than 152 per million in the U.S.
-provided we assume that 18–25 year-olds in the U.S.
-and in Ontario have equivalent comorbidity profiles.
-Making adjustments for better health of Ontarians, ignoring the greater infection-level rise in Ontario, and assuming all the vaccinated were 14 days postdose 2 on 1/1/22, we estimate that at most 214 evaluation-period hospitalizations were prevented by one million COVID vaccinations of 18–25 year-olds in the U.S.
-[Supplement 2, Table S2.7].
-[^23] In addition to the greater infection-level rise in Ontario, other reasons why we view the rate of 214 likely to be an upper bound on the actual rate include the following: our ignoring evidence that incidental COVID hospitalization rates are higher among the young, ignoring evidence that hospitalization rates are higher among females, and applying a better-health adjustment such that “the better health of Ontarians [relative to Americans] is more likely overstated than understated”—see Section 4 of Supplement 2.
-Observe that our most likely scenario’s predicted benefit of 231 hospitalizations prevented per million vaccinations (assuming equivalent IHR for males and females) is only about 1.08 times larger than the estimated actual benefit of 214 [Supplement 2, Table S2.7], while FDA’s most likely scenario’s predicted benefit of 4,766 hospitalizations prevented per million vaccinations is over 22 times larger.
-Moreover, had the FDA assumed equivalent infection and hospitalization rates among males and females, then its corresponding benefit projection likely would have been over 6,000 hospitalizations prevented per million vaccinations.
-Ontario data also supports the bounding analysis of Scenario I from Section 4.1 in which 155.82 in 1 million is used as an upper bound on the evaluation-period COVID-hospitalization risk for a typical unvaccinated male Bob having prior COVID infection and no comorbidities.
-Ontario data suggests that 151 in 1 million is an overestimate of Bob’s evaluation-period COVID hospitalization risk—see Section 2.3 of Supplement 2.
+Consistent with Tables 7 and 8, Table 9 indicates that for those having BMI in the intermediate range 25–34, prior-infection status determines whether vaccination is beneficial.
+However, for the those with high BMI (35–44) in a high-infection-risk scenario Table 9 indicates that vaccination becomes beneficial even for those with prior-infection protection.
+4.3. Validating the Model
+In Supplement 2, we use Ontario COVID-hospitalizations data from the evaluation period 1/1/22– 5/31/22 to show that our model’s projections of Table 6 for the hospitalizations-prevented benefit of vaccination for 18–25-year-old males are much more accurate than those of the FDA [11].
+[^14] We use Ontario data for many reasons; e.g., unlike the CDC, Public Health Ontario chose to publish incidental COVID-hospitalization rates (starting 1/8/2022) [30]; see [Supplement 2, Section 3] for other reasons.
+We test our model in two ways: (i) showing that it fits Ontario COVID-hospitalization data with inputs at plausible levels for Ontario over the evaluation period [Supplement 2 Section 8], and (ii) using Ontario COVID-hospitalization data to predict corresponding data for the U.S.
+[Supplement 2, Section 7].
+For (i), we fit our model to Ontario data for the general-population mix of 18–25-year-old male-and-female Ontarians by using Mathematica to solve a sequence of least-squares problems depending on the value of $H_{r}$.
+For $H_{r}=0.0011$ and $0.0010$, we obtain model inputs in plausible ranges for Ontario that yield a perfect modeling fit and note that these values of $H_{r}$ are consistent with COVID hospitalization data from the state of Connecticut.
+A brief discussion of (ii) is provided in the next paragraph.
+Based on COVID hospitalization counts from Public Health Ontario as well as corresponding incidental hospitalization rates, we estimate that vaccinating one million 18–25-year-old Ontarians prevented 152 COVID hospitalizations over the evaluation period—see Table S2.5.
+Making various adjustments e.g., accounting for the better health of Ontarians vs.
+Americans, we obtain an upper-bound estimate that 214 evaluation-period hospitalizations were prevented by one million full COVID vaccinations of 18–25–year-old males in the U.S.—see [Supplement 2, Table S2.7] and the discussion that follows it.
+Observe that 214 is an upper bound for our projections of Scenarios B–E (see Table 6) but not an upper bound for any of the FDA’s projections.
+Ontario data also supports the definitiveness of the results of Table 7 above given the male has no comorbidities, suggesting that 151 in 1 million is an overestimate of his evaluation-period COVID hospitalization risk [Supplement 2, Section 2.3].
 5. Discussion
-Our reanalysis of the FDA’s benefit-risk assessment [9] of Moderna mRNA-1273 primary-series vaccination reveals substantial shortcomings in the modeling that supported the agency’s finding “the benefits of vaccination with the Moderna COVID-19 vaccine clearly outweigh the risks” even for males of ages 18–25 years.
-By accounting for prior-infection protection and using evidence-based model inputs, we find in our Scenarios B–E that risks outweighed benefits for 18–25-year-old males, assuming hospitalizations are the principal endpoint of analysis.
+By accounting for prior-infection protection and using evidence-based model inputs, we find in our Scenarios B–E of Table 6 that risks of mRNA-1273 vaccination outweighed benefits for the general population of 18–25-year-old males, assuming hospitalizations are the principal endpoint of analysis.
 In particular, for our scenarios predicting plausible infection-level increases over the evaluation period (Scenarios B, D, and E), our model suggests that vaccination of 18–25-year-old males generated between 16% and 63% more hospitalizations from vaccine-attributable myo/pericarditis (VAM/P) alone compared to COVID hospitalizations prevented.
-[^24] The range 16%–63% is consistent with our model-validation work discussed in Section 5: recall we estimate that at most 214 evaluation-period hospitalizations were prevented by one million mRNA-1273 vaccinations of 18–25-year-old males in the U.S.
+[^15] The range 16%–63% is consistent with our model-validation work discussed in Section 4.3: recall we estimate that at most 214 evaluation-period hospitalizations were prevented by one million mRNA-1273 vaccinations of 18–25-year-old males in the U.S.
 [Supplement 2, Table S2.7], and note that our corresponding VAM/P hospitalization projection of 268 is a little over 25% higher than 214.
-Moreover, for the subpopulation of 18–25-year-old males having prior-infection protection and no comorbidities, we have established via Scenarios I and II of Section 4.1 that the choice to mRNA-1273 vaccinate created a greater risk of hospitalization (owing to VAM/P or an evaluation-period breakthrough infection) than the risk of COVD hospitalization if unvaccinated.
-The shortcomings of the FDA’s analysis included not accounting for prior-infection protection as well as incidental hospitalizations, and overestimating the COVID-hospitalization rate for 18–25-year-old males by assuming it to be equivalent to the rate for 18–45-year-old males.
-On the risk side, as discussed in Section 2.3, the FDA relied on a very small internal dataset (21 second-dose events for mRNA-1273 [44]) to determine mRNA-1273 VAM/P risk rather than relying on findings from regulatory agencies from 5 other countries, from the CDC’s VSD system, from major international studies [42, 43], and from a U.S.
-study [45].
-These shortcomings contributed to the FDA’s obtaining in its most likely scenario a benefit-risk ratio of 43.33 for mRNA-1273 vaccination of 18–25-year-old males, a ratio that is over 60 times more favorable than that of our most likely scenario, 0.67.
-6.1. The strength of prior-infection protection
-The FDA decided to not account for prior-infection protection: “We did not factor in protection from potential exposure or infection with COVID-19 among unvaccinated individuals into our analysis since the degree of this protection and its time-dependency were not well understood at the time of the analysis” [9, Section 2.2.1.1].
-However, the protection provided by mRNA-1273 vaccination against the Omicron variant and its time-dependency were also not well understood at the time of analysis (3rd week of January 2022).
-Nevertheless, the FDA assumed in every one of its five Omicron-based benefit-risk scenarios that mRNA-1273 vaccination would provide for at least 5 months’ protection against cases with 30% effectiveness and against hospitalization with 72% effectiveness.
-Even for its Delta-based scenario of [9] the FDA assumed prior-infection would provide no protection, despite (i) the CDC study [18] indicating that prior infection provides superior protection against Delta than does mRNA vaccination and (ii) a CDC Science Brief stating, “Available evidence shows that fully vaccinated individuals and those previously infected with SARS-CoV-2 each have a low risk of subsequent infection for at least 6 months” ([17], 10/29/21).
-As noted in Section 2.2, the FDA’s assumption that prior infection would provide no protection against Omicron infection and hospitalization is contrary to immunological expectations and studies (e.g., [16, 17, 18, 27, 28]) available to the FDA at the time of its mRNA-1273 analysis all suggesting that prior-infection protection would likely equal or exceed vaccine protection among the COVID-naïve.
-As noted in Section 3.3, we have assumed that protection against hospitalization provided by vaccination is 77% while that provided by prior infection is 88.9%.
-A systematic review [25] involving 21 vaccine-effectiveness studies from the period 3 December 2021 – 7 April 2022 reports,
+Recalling that the FDA found, “[T]he benefits of vaccination with the Moderna COVID-19 vaccine clearly outweigh the risks, even among the male adolescent population” [11, Section 4], we have demonstrated that the conclusions of a benefit-risk assessment for vaccination may be dramatically impacted by accounting for the benefits conferred by prior infection.
+We have also demonstrated in Tables 79 how the benefit-risk picture for vaccination may vary based on an individual’s prior-infection and comorbidity status.
+Consistent with our model’s finding that for 18–25-year-old males in the U.S.
+the risks of mRNA-1273 vaccination outweighed the benefits relative to hospitalizations, we note that France suspended mRNA-1273 use on 10/15/2021, making it available again on 11/8/21 but at half-dose and only for those over 30 years [73, p.
+4].
+We emphasize that the cardiac risk of COVID infection is incorporated into our modeling because our estimates of COVID hospitalizations include those for COVID-related cardiac issues.
+We address here assertions such as “Importantly, the risk of myocarditis after any dose of COVID-19 vaccine, in any population, is substantially lower than the risk of myocarditis after SARS-CoV-2 infection,” which appeared, without supporting references, in an April 2024 editorial of the European Heart Journal [74].
+Contrary to the editorial’s assertion, Patone et al.
+find that “In men younger than 40 years old, the number of excess myocarditis events [leading to hospitalization] per million people was higher after a second dose of mRNA-1273 than after a positive SARS-CoV-2 test (97 [95% CI, 91–99] versus 16 [95% CI, 12–18])” [6].
+Karlstad et al.
+also find significantly higher risk of hospitalized myocarditis after dose 2: 183.9 per million 2nd mRNA-1273 doses [5, Table 2] vs.
+13.7 per million positive tests [5, eTable 7] among males 16–24.
+Moreover, these myocarditis incidence rates expressed in cases per million positive tests overstate the risk of myocarditis after COVID infection expressed in cases per million COVID infections (because many infections are not confirmed by positive tests, but essentially all infections among  those seeking care for myocarditis are confirmed by positive tests) [75].
+Consistent with [75], the authors of [76] note, “[T]he attributable risk estimates for COVID-19 [associated myocarditis] used laboratory confirmed cases as the denominator and will be affected by the proportion of all SARS-CoV-2 infections captured by testing, precluding a direct comparison with vaccine-associated attributable risks.”
+The strength of prior-infection protection
+As noted in Section 2.2, the FDA’s assumption that prior infection would provide no protection against Omicron is contrary to immunological expectations and studies (e.g., [15, 16, 17, 19, 26, 27]) available at the time of the FDA’s mRNA-1273 analysis suggesting that prior-infection protection would likely equal or exceed vaccine protection among the COVID-naïve.
+As noted in Table 5 footnotes, we have assumed that vaccination’s effectiveness against hospitalization is 77% while that provided by prior infection is 88.9%.
+A systematic review [24] involving 21 vaccine-effectiveness studies from the period 12/3/2021 – 4/7/2022 reports,
 In contrast to vaccine effectiveness against delta severe disease, the majority of vaccine effectiveness estimates for omicron severe disease were below 75%; for example, thirteen (81%) of sixteen vaccine effectiveness estimates within three months of vaccination with the primary series were below 75% (Figure).
 Moreover, 13 (42%) vaccine-specific estimates fell below 50% at some point in time after vaccination.
-In contrast, the meta-analysis [22] estimates prior-infection protection against severe Omicron to be over 80% through 6 months after infection, while the meta-analysis [23, Table S2] estimates protection to be over 88% through 45 weeks after infection, with the former analysis based on studies from the period 1/1/22 – 6/1/22 and the latter, on studies published up to 9/31/22.
-In early January of 2022, the minimum reasonable assumption the FDA might have made is that prior-infection protection equals vaccine protection (as in our Scenario E); instead, in all Omicron-based scenarios of [9], the agency ascribed to vaccination 72% effective protection against hospitalization while ascribing 0% effective protection to prior infection.
-The FDA’s reliance on the UK Health Security Agency for its estimates of vaccine effectiveness against Omicron [34], instead of relying on the CDC, is surprising.
-The CDC might have had precisely the data the FDA needed for its mRNA-1273 assessment.
-For example, if the CDC extended its May–November 2021 study [18] through the end of December 2021, then the CDC might have had estimates of vaccine as well as prior-infection effectiveness against Omicron at the time of the FDA’s mRNA-1273 analysis.
-The FDA’s assessments of BNT162b2 [8] and mRNA-1273 [9] suggest that the FDA does not necessarily request from the CDC data needed to conduct rigorous analyses.
-For both assessments, the FDA failed to obtain data on incidental COVID hospitalizations.
-For both assessments, the FDA claims that age/sex specific VAM/P information in the CDC’s VSD system is not available; e.g., Funk et al.
-claim “[W]e have no access to VSD data for smaller age/sex subgroups” [8, Section 2.3.2.1].
-Finally, for its mRNA-1273 assessment, the FDA failed to request that the CDC disaggregate COVID hospitalization data for the age range 18–45 into data for the subranges 18–25, 26–35, and 35–45.
-Also troubling is the withholding of and underreporting of ICU-admissions data for VAM/P in, respectively, [8] and [9] (see Section 2.4) as well as willingness to use international data [34] for the benefit analysis of [9] while eschewing international data relating to risks (e.g., VAM/P data from Public Health Ontario [3, 42] and postvaccination myocarditis data from England ([6, 43]).
-6.2. Summary
-We have shown that extending the FDA’s model to account for prior-infection protection as well as incidental COVID hospitalizations and making evidence-based model-input assumptions is sufficient to invert the calculus of seemingly overwhelming benefit of mRNA-1273 vaccination towards excessive risk of VAM/P in the general population of 18–25-year-old males.
-The failure of the FDA author team of [9] to perform a rigorous benefit-risk analysis led to COVID-vaccination recommendations that both failed reflect “The health condition [and prior-infection status] of individuals may remarkably impact the B-R profile and its evaluation” ([9, Section 4]) and that supported vaccination mandates enforced by many institutions placing certain subgroups at a greater risk of hospitalization for a serious adverse event owing to COVID vaccination than for hospitalization owing to COVID infection—while on an absolute basis this risk difference is small, institutional responsibility for adverse events owing to vaccination is not.
-7. Recommendations
-Our modeling provides a call to action to compel our policy-forming health agencies to employ multi-parameter models that can better stratify vaccination recommendations.
-To protect vaccine acceptance and promote public trust, [^25] A survey from the Robert Wood Johnson Foundation and Harvard School of Public Health released May of 2021 found only 37% of respondents had a great deal of trust in the FDA, while 38% somewhat trusted the agency, and 24% described their level of trust “Not very much/Not at all” [74].
-the FDA must adopt a rigorous framework for benefit-risk assessment of new vaccines and revisit recommendations as new data becomes available.
+In contrast, the meta-analysis [21] estimates prior-infection’s effectiveness against severe Omicron to be over 80% through 6 months after infection, while the meta-analysis [22, Table S2] estimates its effectiveness to be over 88% through 45 weeks after infection, with the former estimate based on studies from 1/1/22–6/1/22 and the latter, on studies published by 9/31/22.
+In mid-January of 2022, the minimum reasonable assumption the FDA might have made is that prior-infection protection would equal vaccine protection (as in our Scenario E); instead, in all Omicron-based scenarios of [11], the agency ascribed to vaccination 72% effective protection against hospitalization while assuming prior infection offers no protection.
+5.1. Recommendations
+Our modeling provides a call to action to policy-forming health agencies to employ multi-variable models like ours that can better stratify vaccination recommendations.
+To protect vaccine acceptance and promote public trust, the FDA must adopt a rigorous framework for benefit-risk assessment of new vaccines and revisit recommendations as new data becomes available.
 In particular, the framework must incorporate consideration of the protection that prior infection affords against development of severe disease.
-A re-alignment of priorities of the public-health agencies of the United States is necessary.
-Two years after the first cases SARS-CoV-2 were identified in the U.S., the FDA decided to “not investigate the benefits and risks [of mRNA-1273 vaccination] for specific subpopulations, such as those with comorbidities, due to limited information for these populations” [9, Section 4].
-Also, public-health agencies should have prioritized understanding the effectiveness of prior-infection protection against reinfection and hospitalization.
-That said, in Pfizer’s phase-3 trial for BNT162b2, 567 participants in the placebo group and 526 in the treatment arm had evidence of prior COVID-19 infection.
-In each of these cohorts, there was only 1 reinfection (or ¡0.2% for both) according to the primary endpoint definition [75, p.
+We also recommend United States public-health agencies reassess their priorities.
+Two years after the first cases of COVID were identified in the U.S., the FDA decided to “not investigate the benefits and risks [of mRNA-1273 vaccination] for specific subpopulations, such as those with comorbidities, due to limited information for these populations” [11, Section 4].
+Also, public-health agencies had reason to prioritize understanding the effectiveness of prior-infection protection and had evidence that it was likely to be high.
+For example, in Pfizer’s phase-3 trial for BNT162b2, 567 participants in the placebo group and 526 in the treatment arm had evidence of prior COVID-19 infection.
+In each of these cohorts, there was only 1 reinfection (or <0.2<0.2% for both) [77, p.
 27].
-This alone should have prompted health agencies to aggressively pursue further data on protection from prior infection to stratify risk and recommendations.
-We note that by late fall of 2022, the percentage of the U.S.
-population having experienced at least one COVID infection was estimated to be 95% by an IHME study [76] and 94% by a Harvard study [77].
-Given that nearly all persons in the United States now have significant protection from severe COVID conferred by prior-infection, there is an urgent need to reassess whether the following current blanket CDC recommendations [78] for COVID-19 vaccination advance public health: “ Everyone ages 6 months and older should get a 2024–2025 COVID-19 vaccine.” Compare the CDC’s policy to that of the UK: “All those aged 65 or over, all those in care homes for older people, and those aged 6 months or over who are in a clinical risk group will be eligible for vaccination this autumn” [79].
-The urgency of the need for reassessment is reinforced by the fact that Pfizer’s and Moderna’s reformulated COVID-19 vaccines are available only via emergency-use authorization for individuals 6 months through 11 years of age (despite the expiration of the COVID-19 public health emergency 11 May 2023).
-8. Limitations
-We have assumed that young males are hospitalized for COVID at a higher rate than young females, as did the FDA in its benefit-risk assessments [8] and [9].
+5.2. Limitations
+Our principal aim has been to illustrate how prior-infection protection, as well as comorbidity status, can be integrated into benefit-risk modeling of a two-dose vaccination series.
+In applying our model to extend and improve the FDA’s mRNA-1273 assessment [11] relative to its hospitalizations endpoint, we have assumed that young males are hospitalized for COVID at a lower rate than young females, as did the FDA in [10] and [11].
 Data from [10] and [54] suggests this rate difference is substantial; however, the rates of [10] and [54] include incidental hospitalizations and incidental rates for males may differ from those for females.
-We have assumed, as did the FDA, that protection provided by vaccination and prior infection against hospitalization and infection is essentially equivalent for males and females, and this may not be the case.
+We have assumed, as did the FDA, that vaccination and prior infection offer essentially equivalent protection for males and females.
 Also, we have based our assessment of the difference in hospitalization rates between young males and females on data in Table 2 of [10] representing mostly pre-Delta cases.
@@
 However, the rate differences provided by Table 1 of [54], based on two years of data including that of the initial Omicron wave, are consistent with those of Table 2 of [10].
-COVID-hospitalizations data from Public Health Ontario (see Supplement 2) and from the State of Connecticut (see the appendix of Supplement 1) suggests that that our Omicron IHR value $0.0014$ for 18–25-year-old males is a reasonable estimate for the Omicron IHR for all 18–25 year-olds (general population mix of males and females).
+COVID-hospitalizations data from Ontario [Supplement 2] and from Connecticut [Supplement 1, Appendix 1] suggests that our Omicron IHR value $0.0014$ for 18–25-year-old males is a reasonable estimate for the Omicron IHR for all 18–25 year-olds (general population mix of males and females).
 Thus, even if we assume males and females in this age range are hospitalized for COVID at roughly equivalent rates, our IHR value of 0.14% for 18–25-year-old males appears to be a good estimate.
-As noted in Section 3.3, based on data available before 1/22/22, we arrived at an Omicron IHR estimate for all 18–25 year-olds of 0.22%, not 0.14%.
+In Section 2 of Supplement 1, based on data available before 1/22/22, we obtain an Omicron IHR estimate for 18–25 year-olds (general-population mix of males and females) of 0.22%, not 0.14%.
 One possible explanation is that we may have underestimated the incidental COVID-hospitalization rate for 18–25 year-olds.
-Our IHR estimate for 18–25 year-olds of 0.22% results from assuming an incidental COVID-hospitalization rate of 40%, which is the rate suggested by [29] and [31] for hospitalizations with COVID among patients of all ages.
-However, there is strong evidence that incidental COVID-hospitalization rates are higher among the young than among the elderly; see, e.g., [30], [80], and Supplement 1.
-Figure 2 of [30] suggests an incidental rate of over 60% among 0–18 year-olds from April of 2020 through May of 2021.
-COVID-hospitalization data from the CDC and the state of Connecticut, discussed in the appendix of Supplement 1, suggests an incidental rate of about 72% among 15–24 year-olds in Connecticut during calendar year 2020.
+Our IHR estimate for 18–25 year-olds of 0.22% results from assuming an incidental Omicron-hospitalization rate of 40%, a rate suggested by [28] and [30] for hospitalizations with COVID among patients of all ages.
+However, as noted in Section 2.2 (and discussed in detail in Supplement 1), there is strong evidence that incidental COVID-hospitalization rates are higher among the young than among the elderly.
+In particular, COVID-hospitalization data discussed in Appendix 1 of Supplement 1 suggests an incidental rate of over 70% among 15–24 year-olds in Connecticut during calendar year 2020.
 Assuming an incidental rate of about 62.2% (instead of 40%) transforms our IHR estimate of 0.22% into the estimate 0.14%.
-In terms of applying our model to an emerging pathogen and vaccine campaign in the future, our functions $U$ and $V$ require reasonably accurate inputs to produce reasonably accurate projected hospitalizations as outputs.
-In particular, if public-health agencies do not provide accurate rates of hospitalizations (requiring treatment for the infection) based on age, vaccination-status, sex, and prior-infection status, then obtaining accurate infection-hospitalization rates and rate-reduction factors becomes difficult.
-In applying our improvement of the FDA’s model for the general population of 18–25-year-old males, we constrained ourselves to using only information available at the time of the FDA’s benefit-risk analysis of mRNA-1273 .
-Our assessment of risks associated with vaccination, like the FDA’s, ignores known rare but potentially severe vaccine-related adverse events other than myo/pericarditis that may require hospitalization, for example, anaphylaxis [58], Multisystem Inflammatory Syndrome [59, 60], Guillain Barré Syndrome [81], posterior orthostatic tachycardia syndrome (POTS) [82], retinal vascular occlusion [83], and cerebral venous sinus thrombosis (CVST) [64].
-Our risk assessment as well as the FDA’s ignores the possibility that COVID-19 vaccination has long-term health effects; for instance, the FDA’s package-insert information [84, Section 13] for Moderna’s mRNA-1273 begins, “SPIKEVAX has not been evaluated for carcinogenic, mutagenic potential, or impairment of male fertility in animals.”
-Finally, we have carried out a retrospective analysis illustrating how the FDA might have conducted its mRNA-1273 benefit-risk assessment in a more rigorous way.
-Data we rely on for assessing benefit and risks of vaccination pertains to the original, monovalent formulation of Moderna’s COVID vaccine, having a two-dose vaccination schedule.
-Acknowledgments: We thank Retsef Levi for providing comments that improved both the structure and content of this paper.
-Links to Supplements:
+For a vaccine having high effectiveness against infection that does not quickly wane, vaccination can have an important benefit not captured by our model—providing an enduring reduction in the level of infection in the population that may allow some members to avoid infection entirely through the contribution of vaccination to herd immunity.
+Our risk assessment, like the FDA’s, ignores known, rare vaccine-related adverse events other than myo/pericarditis that may require hospitalization.
+Finally, our risk assessment, like the FDA’s, ignores the possibility that COVID-19 vaccination may have long-term health effects; for instance, the FDA’s package-insert information for Moderna’s mRNA-1273 begins, “SPIKEVAX has not been evaluated for carcinogenic, mutagenic potential, or impairment of male fertility in animals” [78, Section 13].
+Acknowledgment: We thank Retsef Levi for providing comments that improved both the structure and content of this paper.
+Supplements: Supplement 1, Supplement 2

### v4

@@
-A reanalysis of the FDA’s benefit-risk assessment of Moderna’s mRNA-1273 COVID vaccine stratified not only based on age and sex but also on prior-infection and comorbidity status
+A Reanalysis of the FDA’s Benefit–Risk Assessment of Moderna’s mRNA-1273 COVID Vaccine Based on a Model Incorporating Benefits Derived from Prior COVID Infection
 Abstract.
-We demonstrate that the outcome of a benefit-risk assessment for vaccination may be dramatically impacted by accounting for the benefits derived from prior infection by the disease against which the vaccine is targeted.
-The United States Food and Drug Administration (FDA) conducted a benefit-risk assessment for Moderna’s COVID vaccine mRNA-1273 prior to its full approval, announced 1/31/2022.
-The FDA’s assessment focused on males of ages 18–64 years because the agency’s risk analysis was limited to vaccine-attributable myocarditis/pericarditis (VAM/P) given the excess risk among males.
-The FDA’s analysis concluded that vaccine benefits clearly outweighed risks, even for 18–25-year-old males (those at highest VAM/P risk).
-We reanalyze the FDA’s benefit-risk assessment using information available through the third week of January 2022 and focusing on 18–25-year-old males.
-We use the FDA’s framework but extend its model by accounting for protection derived from prior COVID infection, finer age-stratification in COVID-hospitalization rates, and incidental hospitalizations (those of patients who test positive for COVID but are being treated for something else).
-We also use more realistic projections of Omicron-infection rates and more accurate rates of VAM/P.
-With hospitalizations as the principal endpoint of the analysis (those prevented by vaccination vs.
+Background: The U.S.
+Food and Drug Administration (FDA) conducted a benefit–risk assessment for Moderna’s COVID vaccine mRNA-1273 prior to its full approval, announced 31 January 2022.
+The FDA’s assessment focused on males 18–64 years old because its risk analysis was limited to vaccine-attributable myocarditis/pericarditis (VAM/P), given the excess risk among males.
+The FDA’s analysis concluded that vaccine benefits outweighed risks, even for 18–25-year-old males (those at highest VAM/P risk).
+We reanalyze the FDA’s benefit–risk assessment using information available through the third week of January 2022 and focusing on 18–25-year-old males.
+Methods: We develop a benefit–risk model, extending the FDA’s, that can stratify benefits and risks of vaccination by prior-infection and comorbidity status.
+We use the FDA’s framework but apply our model to account for benefits derived from prior COVID infection, while also accounting for finer age stratification in COVID-hospitalization rates, incidental hospitalizations (those of patients who test positive for COVID but receive treatment for something else), more realistic projections of Omicron-infection rates, and more accurate VAM/P rates.
+Results: With hospitalizations as the principal endpoint of the analysis (those prevented by vaccination vs.
 those caused by VAM/P), our model finds vaccine risks outweighed benefits for 18–25-year-old males, except in scenarios projecting implausibly high Omicron-infection prevalence.
-Our assessment suggests that mRNA-1273 vaccination of 18–25-year-old males generated between 16% and 63% more hospitalizations from vaccine-attributable myocarditis/pericarditis alone compared to COVID hospitalizations prevented (over a five-month period of vaccine protection assumed by the FDA).
-The preceding assessment derives from model inputs based on data available at the time of the FDA’s mRNA-1273 assessment.
+Our assessment suggests that mRNA-1273 vaccination of 18–25-year-old males generated between 8% and 52% more hospitalizations for VAM/P compared to COVID hospitalizations prevented (over a five-month period of vaccine protection assumed by the FDA).
+The preceding assessment uses model inputs based on data available at the time of the FDA’s mRNA-1273 assessment.
 Moreover, these inputs as well as model outputs are validated by subsequently available data.
+Conclusions: The outcome of a vaccine benefit–risk assessment may be dramatically impacted by accounting for the benefits derived from prior infection by the vaccine-targeted disease.
+To increase public confidence in vaccines and thereby reduce vaccine hesitancy, public-health agencies should employ benefit–risk models capable of supporting stratification of vaccination recommendations not only based on age and sex but also on prior-infection and comorbidity status.
 1. Introduction
-Because vaccines can be given to large populations of healthy persons and can be introduced as mandatory, robust benefit-risk analyses must be conducted to ensure that vaccines have highly favorable benefit-risk profiles [1].
+Because vaccines can be given to large populations of healthy persons and can be introduced as mandatory, robust benefit–risk analyses must be conducted to ensure that vaccines have highly favorable benefit–risk profiles [1].
 The Food and Drug Administration (FDA) issued emergency-use authorizations for mRNA COVID-19 vaccines from Pfizer (BNT162b2) and Moderna (mRNA-1273) in December 2020 [2].
-Once the mRNA COVID-19 vaccines were given to large populations, it became evident that they could cause myocarditis or pericarditis (hereafter myo/pericarditis), inflammation of the heart or of the membrane surrounding the heart, respectively, with studies indicating the highest risk among males younger than 40 years [3, 4, 5, 6, 7, 8, 9].
-The FDA has published two benefit-risk analyses of primary series mRNA vaccination, the first for BNT162b2 [10] and the second for mRNA-1273 [11].
-The FDA presented its analyses through scenarios involving different modeling assumptions to account for uncertainties in pandemic dynamics as well as in the benefits and risks of vaccination.
+Once the mRNA COVID-19 vaccines were given to large populations, it became evident that they could cause myocarditis or pericarditis (hereafter myo/pericarditis), inflammation of the heart or of the membrane surrounding the heart, respectively, with studies indicating the highest risk among males under 40 years old [3, 4, 5, 6, 7, 8, 9].
+Possible mechanisms by which mRNA COVID vaccination causes myocarditis are explored in  [10, 11].
+The FDA published two benefit–risk analyses of primary series mRNA vaccination, the first for BNT162b2 [12] and the second for mRNA-1273 [13].
 To assess vaccine benefits, the FDA estimated vaccine-preventable COVID-19 cases, hospitalizations, intensive-care unit (ICU) admissions, and deaths [12, 13].
 To assess vaccine risks, the FDA estimated vaccine-attributable myo/pericarditis (VAM/P) cases, hospitalizations, ICU admissions, and deaths [12, 13].
-The FDA concluded that the benefits of the mRNA vaccines outweighed VAM/P risk even for males in its highest-risk age-groups: 16–17 years for BNT162b2 [10] and 18–25 years for mRNA-1273 [11].
-However, these FDA benefit-risk assessments omitted important inputs in the analysis of both the hospitalizations-prevented benefit of vaccination and VAM/P risk.
-We focus mainly on the FDA’s analysis for Moderna’s mRNA-1273 to simplify our exposition—there are significant differences in modeling assumptions in the FDA’s Pfizer and Moderna assessments.
+The FDA concluded that the benefits of the mRNA vaccines outweighed VAM/P risk even for males in its highest-risk age groups: 16–17 years for BNT162b2 [12] and 18–25 years for mRNA-1273 [13].
+To simplify our exposition, we focus mainly on the FDA’s analysis for Moderna’s mRNA-1273—there are significant differences in modeling assumptions in the FDA’s Pfizer and Moderna assessments.
 Also, VAM/P risk is substantially greater for mRNA-1273 than for BNT162b2 [3, 4, 5, 6, 7, 8, 9].
-For the FDA’s assessment of mRNA-1273, benefits of vaccination were overestimated.
-First, the FDA did not account for the protection conferred by prior infection; and, at the time of the FDA’s Moderna analysis (mid-January 2022), likely over 70% of 18–25 year-olds in the U.S.
-had been infected by COVID-19—see Supplement 1.
-Second, the FDA did not account for incidental COVID-19 hospitalizations.
-Third, the FDA overestimated benefits for young males by assuming the COVID-hospitalization rate for 18–25 year-olds is equivalent to that for 26–35 year-olds and 36–45 year-olds.
-Finally, the FDA assumed in its most likely scenario that case and hospitalization rates would remain constant at peak 2021 Omicron levels for 5 months.
-By contrast, the FDA underestimated mRNA-1273 VAM/P risks; see, e.g.,Table 3 of Section 2.3.
-Here, our primary aim is to demonstrate how to incorporate the protection derived from prior infection into a population-level quantitative benefit-risk model for a two-dose vaccination series.
+Our benefit–risk modeling differs from the FDA’s primarily by its accounting for the protection conferred by prior infection, and, at the time of the FDA’s Moderna analysis (mid-January 2022), likely over 70% of 18–25-year-olds in the U.S.
+had been infected by COVID-19—see Supplement S1.
+On the risk side, we use data available at the time of the FDA’s mRNA-1273 assessment, including that from [14, 15, 16, 12, 17], suggesting VAM/P risk for mRNA-1273 exceeded FDA estimates; see Section 1.4 below.
+Here, our principal aim is to demonstrate how to incorporate the protection derived from prior infection into a population-level quantitative benefit–risk model for a two-dose vaccination series.
 Our model has hospitalizations as its endpoint, comparing the hospitalizations-prevented benefit of vaccination to the hospitalization risk attributable to vaccination.
@@
 In addition to prior-infection protection, our modeling accounts for incidental hospitalizations, age-specific infection-hospitalization rates, and a more accurate assessment of VAM/P risk.
-We apply our model using data available to the FDA at the time of its benefit-risk analysis for mRNA-1273, assumed to be the 3rd week of January 2022 for reasons described in Section 2.1.
+We apply our model using data available to the FDA at the time of its benefit–risk analysis for mRNA-1273, assumed to be the 3rd week of January 2022 (see Section 1.2).
 We present evidence, using real-world data, that our model’s estimates of the hospitalizations-prevented benefit of mRNA-1273 vaccination are significantly more accurate than those the FDA obtained in [13].
 We also analyze benefits and risks of vaccination for specific subpopulations such as males 18–25 having prior-infection protection and no comorbidities.
-Our model can be adapted to produce for any vaccine designed for deployment on a population-wide scale a benefit-risk assessment stratified not only based on age and sex but also on prior-infection and comorbidity status.
-2. The FDA’s mRNA-1273 Benefit-Risk Assessment
-In [11], the FDA describes its benefit-risk modeling of Moderna’s mRNA-1273 for males ages 18 to 64 years.
+Our model can be adapted to produce a benefit–risk assessment for any vaccine designed for deployment on a population-wide scale, stratified not only based on age and sex but also on prior-infection and comorbidity status.
+We emphasize that the application of our model to reanalyze the FDA’s benefit–risk assessment for mRNA-1273 is for demonstration purposes and does not constitute a comprehensive re-evaluation, which would address all of the FDA’s benefit–risk endpoints (cases, hospitalizations, ICU admissions, deaths).
+1.1. Framework of the FDA’s mRNA-1273 Benefit–Risk Assessment
+In [13], the FDA describes its benefit–risk modeling of Moderna’s mRNA-1273 for males ages 18 to 64 years.
 Outcomes are provided via six scenarios reflecting uncertainties in three major model inputs: COVID-19 incidence, vaccine effectiveness against cases ($VE$) as well as hospitalization ($VEH$), and vaccine-attributable myo/pericarditis (VAM/P) rate.
 At the time of the FDA’s mRNA-1273 analysis, the CDC was recommending a 5-month interval between primary-series completion and a booster dose.
-Consequently, the FDA evaluated benefits and risks over a 5-month period, which we take to be 1/1/22 – 5/31/22.
+Consequently, the FDA evaluated benefits and risks over a 5-month period, which we take to be 1 January 2022–31 May 2022.
 The FDA assumed that all vaccine effectiveness estimates remained constant over this 5-month evaluation period.
 The FDA’s major modeling assumptions are summarized in Table 1 below.
-Table 1. The FDA’s mRNA-1273 model-input assumptions for 18–25-year-old males [11, Table 1].
-Scenario COVID-19 Incidence Vaccine Effectiveness VAM/P Rate Per Million 2nd doses [11, Table 2]
+Table 1. The FDA’s mRNA-1273 model-input assumptions for 18–25-year-old males [13, Table 1].
+Scenario COVID-19 Incidence Vaccine Effectiveness VAM/P Rate Per Million 2nd doses [13, Table 2]
 11 Highest 2021 Incidence, 12/25–12/31 Omicron Dominant; VE = 30%, VEH=72% 128 cases, 110 hospitalizations
@@
 6 Same as Scenario 1 Same as Scenario 1 241 cases, 207 hospitalizations
-Using inputs from first row of Table 1 and incidence data from [11, Table 3], the FDA concluded for Scenario 1, its “most likely scenario,” that vaccinating one million males ages 18–25 years with two doses would prevent, over 5 months, 82,484 COVID-19 cases, 4,766 COVID-19 hospitalizations, 1,144 ICU admissions, and 51 deaths due to COVID-19, while risking 128 VAM/P cases, 110 VAM/P hospitalizations, 0 ICU admissions and 0 deaths from VAM/P [11, Table 2].
-.
+Using inputs from the first row of Table 1 and incidence data from [13, Table 3], the FDA concluded for Scenario 1, its “most likely scenario,” that vaccinating one million males ages 18–25 years with two doses would prevent, over 5 months, 82,484 COVID-19 cases, 4766 COVID-19 hospitalizations, 1144 ICU admissions, and 51 deaths due to COVID-19, while risking 128 VAM/P cases, 110 VAM/P hospitalizations, 0 ICU admissions and 0 deaths from VAM/P [13, Table 2].
+Observe that the FDA does include a high VAM/P-risk scenario (6) as well as lower-benefit scenarios (2 and 3).
 1.2. The Time of the FDA’s mRNA-1273 Analysis
-Although the FDA does not define in [11] the time of its benefit-risk analysis for mRNA-1273, a natural choice is the 3rd week of January 2022, which concluded 10 days before the FDA’s announcement of the full approval of mRNA-1273 on 1/31/2022 [12].
-This choice is consistent with the FDA’s time-of-analysis definition for its BNT162b2 assessment: the 2nd week of August 2021 [10, Section 2.1], which concluded nine days before the FDA’s announcement of the full approval of BNT162b2 on 8/23/2021 [13].
-Consequently, our reanalysis of the FDA’s benefit-risk assessment [11] is based on model inputs derived from information available before 1/22/2022.
+Although the FDA does not define in [13] the time of its benefit–risk analysis for mRNA-1273, a natural choice is the 3rd week of January 2022, which concluded 10 days before the FDA’s announcement of the full approval of mRNA-1273 on 31 January 2022 [18].
+This choice is consistent with the FDA’s time-of-analysis definition for its BNT162b2 assessment: the 2nd week of August 2021 [12, Section 2.1], which concluded nine days before the FDA’s announcement of the full approval of BNT162b2 on 23 August 2021 [19].
+Consequently, our reanalysis of the FDA’s benefit–risk assessment [13] is based on model inputs derived from information available before 22 January 2022.
 Subsequent information validating our input choices will often be provided but only via footnotes.
-Starting in Section 4.2, we discontinue basing our modeling on information only available before 1/22/2022; in Section 4.2, our intent is to show how our model can be modified to incorporate a comorbidity (obesity) as a risk modifier.
-2.2. The FDA’s mRNA-1273 Benefits Assessment
+Starting in Section 3.2, we discontinue basing our modeling on information available only before 22 January 2022; in Section 3.2, our intent is to show how our model can be modified to incorporate a comorbidity (obesity) as a risk modifier.
+1.3. The FDA’s mRNA-1273 Benefit Assessment
 The FDA’s computation in [13] of the cases-prevented and hospitalizations-prevented benefits of vaccination is straightforward.
@@
 $C\cdot VE\ \text{cases prevented and}\ H\cdot VEH\ \text{hospitalizations prevented},$
-where $VE$ and $VEH$ are, respectively, the effectiveness of full mRNA-1273 vaccination against cases and against hospitalization; the FDA assumed that for Omicron $VE=0.30$ and $VEH=0.72$, as noted in Table  1 above.
-For Scenario 1, its most likely scenario, and two others, the FDA assumed that peak 2021 Omicron-case rates (late December 2021) would be sustained for 5 months, resulting in an exceptionally high projection of COVID-19 cases among the unvaccinated; e.g., $C=274,947$ for 18–25-year-old males.
-Similarly, for these scenarios, the FDA assumed that the CDC’s reported late-December 2021 Omicron hospitalization rate for unvaccinated 18–45-year-old males would be sustained for 5 months at approximately 43.54 cases per million daily [11, Table 3] and assumed that the same rate would apply to the younger 18–25 bracket, arriving at $H=6,619$ hospitalizations.
-Thus, for these scenarios, the FDA predicted that full mRNA-1273 vaccination of 1 million 18–25-year-old males would prevent $274,947\cdot 0.30\approx 82,484$ COVID cases and $6619\cdot 0.72\approx 4,766$ COVID hospitalizations over the 5-month evaluation period [11, Table 2].
-The following factors contributed to the FDA’s overestimation of vaccine benefits:
-$\bullet$
-The FDA did not account for the protection conferred by prior infection.
-The FDA’s disregarding prior-infection protection in its mRNA-1273 assessment is difficult to justify.
-Dan et al.
+where $VE$ and $VEH$ are, respectively, the effectiveness of full mRNA-1273 vaccination against cases and against hospitalization; the FDA assumed that for Omicron $VE=0.30$ and $VEH=0.72$, as noted in Table 1 above.
+[…]
+How our benefit assessment differs from the FDA’s:
+$\bullet$ We account for the protection conferred by prior infection.
+At the time of the FDA’s mRNA-1273 assessment, there was substantial evidence that prior infection would provide significant protection against COVID infection and hospitalization.
+In a study published 5 January 2021, Dan et al.
 found “About 95% of subjects retained immune memory at $\sim$6 months after [COVID] infection” [20].
-A meta-analysis based on studies published December 2020 through August 2021 demonstrated that protection from prior COVID infection “is, at least, equivalent to the protection afforded by complete vaccination of COVID-naïve populations” [15].
-The CDC acknowledged in a science brief released 10/29/21 that “The immunity provided by vaccine and prior infection are both high” [16], while another CDC study found that during the Delta wave unvaccinated persons with prior COVID infection had significantly greater protection against infection and hospitalization than did fully vaccinated, COVID-naïve persons [17].
-Moreover, in December 2021, the FDA acknowledged, “The omicron variant has significantly more mutations than previous SARS-CoV-2 variants, particularly in its S-gene, the gene that encodes the virus’s spike protein”[18], suggesting that prior infection, entailing exposure to all SARS-CoV-2 proteins (such as the nucleoprotein), would likely yield greater protection against Omicron than would mRNA-1273 vaccination, which entails exposure only to the spike protein of the ancestral strain of SARS-CoV-2.
+A meta-analysis of studies published December 2020 through August 2021 demonstrated that protection from prior COVID infection “is, at least, equivalent to the protection afforded by complete vaccination of COVID-naïve populations” [21].
+The CDC acknowledged in a 29 October 2021 science brief that “The immunity provided by vaccine and prior infection are both high” [22], while another CDC study found that during the Delta wave, unvaccinated persons with prior COVID infection had significantly greater protection against infection and hospitalization than did fully vaccinated, COVID-naïve persons [23].
+Moreover, in December 2021, the FDA acknowledged, “The omicron variant has significantly more mutations than previous SARS-CoV-2 variants, particularly in its S-gene, the gene that encodes the virus’s spike protein”[24], suggesting that prior infection, entailing exposure to all SARS-CoV-2 proteins (e.g., the nucleoprotein), would likely yield greater protection against Omicron than would mRNA-1273 vaccination, which entails exposure only to the spike protein of the ancestral strain of SARS-CoV-2.
 Among the findings of a study of SARS-CoV-2 specific B and T cell memory is “CD8+ T cell responses preferentially target the nucleoprotein, highlighting the potential importance of including the nucleoprotein in future vaccines” [25].
-[^1] Consistent with the evidence presented in this paragraph, a Qatari study found prior infection provided 52.2% effective protection against symptomatic Omicron infection and 91% effective protection against “severe, critical, or fatal” Omicron, with corresponding effectiveness estimates for 2 doses of mRNA-1273 being 2.2% and 66.3%, respectively [20, Figure 2C, 2D] (study period 12/23/21 – 2/21/22).
-Findings from the meta-analyses and systematic reviews [21, 22, 23, 24] provide convincing evidence that prior infection provided greater protection against Omicron during the evaluation period than did two-dose vaccination among the COVID-naïve [Supplement 1, Section 7].
-Further evidence of the need to account for prior-infection protection includes (i) a CDC estimate that 54.9% of 18–49 year-olds had been infected by COVID-19 by 9/30/2021 [25], (ii) a South African study [26] released 12/14/21 suggesting overall effectiveness of prior-infection protection against Omicron to be about 47% [Supplement 1, Section 4], and (iii) data from England [27] released 12/22/21 suggested that prior infection among the unvaccinated provides significant protection from Omicron hospitalization [Supplement 1, Section 5].
-$\bullet$ The FDA did not account for incidental COVID-19 hospitalizations.
+[^1] Consistent with the preceding paragraph’s evidence, a Qatari study found prior infection provided 52.2% effective protection against symptomatic Omicron infection and 91% effective protection against “severe, critical, or fatal” Omicron, with corresponding effectiveness estimates for two doses of mRNA-1273 being 2.2% and 66.3%, respectively [26, Figure 2C, 2D] (study period 23 December 2021–21 February 2022).
+Findings from the meta-analyses and systematic reviews [27, 28, 29, 30] provide convincing evidence that prior infection provided greater protection against Omicron during the evaluation period than did two-dose vaccination among the COVID-naïve [Supplement S1, Section S7].
+Further evidence of the need to account for prior-infection protection includes (i) a CDC estimate that 54.9% of 18–49 year-olds had been infected by COVID-19 by 30 September 2021 [31], (ii) a South African study [32] released 14 December 2021 suggesting overall effectiveness of prior- infection protection against Omicron to be about 47% [Supplement S1, Section S4], and (iii) data from England [33] released 22 December 2021 suggesting that prior infection among the unvaccinated provides significant protection from Omicron hospitalization [Supplement S1, Section S5].
+$\bullet$ We account for incidental COVID-19 hospitalizations.
 At the time of its mRNA-1273 assessment, incidental-hospitalization rates were well known to be substantial.
-On 1/9/22, the director of the CDC stated, “[U]p to 40 percent of the patients who are coming in with COVID-19 are coming in …with something else” [28].
+On 9 January 2022, the director of the CDC stated, “[U]p to 40 percent of the patients who are coming in with COVID-19 are coming in $\ldots$ with something else” [34].
 [^2] We searched unsuccessfully for published CDC incidental-COVID-hospitalizations data.
-Figure 2 of a large U.S.
-study suggests an incidental rate for 0–18 year-olds typically well above 50% during its study period 3/1/2020–1/31/22 and for those 19–64 an incidental rate above 40% for 1/1/2022–1/31/2022 [29].
-On 1/10/22, Ontario began publishing incidental rates, with rates for January 8–15 exceeding 45% [30].
+A large U.S.
+study suggests an incidental rate for 0–18 year-olds typically well above 50% during its study period 1 March 2020–31 January 2022 and for those 19–64 an incidental rate above 40% for 1 January 2022–31 January 2022 [35, Figure2].
+On 10 January 2022, Ontario began publishing incidental rates, with rates for January 8–15 exceeding 45% [36].
 There is evidence that incidental rates are especially high for the young.
-For example, a major study [31] found, “As of May 2021, the cumulative rate for hospitalization associated with SARS-CoV-2 infection was 35.9 per 10,000 children” and “When limiting the analysis to include only patients with COVID-19 who required therapeutic interventions, the hospitalization rate decreased 5.5-fold to 6.5 per 10,000 children,” which suggests an incidental rate of over 80% “for children” (persons 0-17-years-old).
-See Supplement 1 for additional incidental-hospitalization data.
-$\bullet$ The FDA overestimated the proportion of projected COVID cases requiring hospitalization.
-All of the FDA’s projected case-hospitalization rates, which are described in Table 2 below, are over four times the Omicron case-hospitalization rate ($0.58\%\approx 56/9,585$) for the unvaccinated of all ages based on Omicron cases in England tested between 1 December 2021 and 14 December 202 [27, Table 3]; moreover, all are well over twice the case-hospitalization rate (0.92%) for the unvaccinated of ages 18–29 in Ontario 12/14/20 – 12/12/21 [32, Table A1a].
-[^3] The FDA case-hospitalization rates in Table 2 are all well over 5 times the Omicron rate 0.42% for the unvaccinated of all ages reported in Table S10 of [33] (data from 12/15/21 – 1/17/22).
-Based on information available at the time of the FDA’s mRNA-1273 assessment, we expect the COVID-case hospitalization rate for the period 12/14/20 – 12/12/21 to be higher than that for Omicron; e.g., “The risk of hospital admission from emergency departments with Omicron was approximately one-third of that for Delta (Hazard Ratio 0.33, 95% CI: 0.30 to 0.37)” [34, p.
-4]; see Supplement 1 for more examples.
-Furthermore, we expect the hospitalization rate for those 18–25 to be lower than that for 18–29 [35].
-Factors contributing to the FDA’s overestimating case-hospitalization rates for the 18–25 age range include (i) assuming hospital-admission rates were uniform for ages 18–45, contradicting CDC models [36], and (ii) including incidental COVID hospitalizations.
-Our source for the 0.58% rate mitigated the incidental-hospitalization issue [27, p.
-10], while our source [32] for the 0.92% rate did not.
-Table 2. The FDA’s case-hospitalization-rate projections for unvaccinated 18–25-year-old males over the evaluation period 1/1/22 – 5/31/22 for the Omicron-based scenarios of [11, Table 2].*
-Scenario Projected Cases Projected Hospitalizations Projected Case-Hospitalization Rate
-11   274,947   6,619 $\frac{6619}{274,947}\approx 2.41\%$
-2 89,017 2,900 $\frac{2900}{89,017}\approx 3.26\%$
-3 13,010 882 $\frac{882}{13,010}\approx 6.78\%$
-55 274,947 6,619 $\frac{6619}{274,947}\approx 2.41\%$
-66 274,947 6,619 $\frac{6619}{274,947}\approx 2.41\%$
-*Case and hospitalization projections are obtained from Table 2 of [11] by dividing, respectively, cases-prevented estimates by 0.30 and hospitalizations-prevented estimates by 0.72.
-$\bullet$ The FDA assumed in its most likely scenario (as well as Scenarios 5, and 6) that case and hospitalization rates would remain constant at peak 2021 Omicron levels for 5 months.
+For example, a major study [37] found, “As of May 2021, the cumulative rate for hospitalization associated with SARS-CoV-2 infection was 35.9 per 10,000 children” and “When limiting the analysis to include only patients with COVID-19 who required therapeutic interventions, the hospitalization rate decreased 5.5-fold to 6.5 per 10,000 children,” which suggests an incidental rate of over 80% “for children” (persons 0–17-years-old).
+See Supplement S1 for additional incidental-hospitalization data.
+$\bullet$ We account for finer age stratification in COVID-hospitalization rates.
+The FDA assumed that hospital admission rates were uniform for males of ages 18–45, contradicting CDC models indicating that hospitalization rates for those 30–49 are two times the rates for 18–29-year-olds [38].
+In contrast, we employ an Omicron infection-hospitalization rate estimate for precisely the group of interest—males, 18–25 years old.
+Our estimate, based on a model developed in [39], is consistent with COVID-hospitalization data for the U.S.
+military [Supplement S1, Section S2].
+$\bullet$ We use a CDC case-to-infections multiplier to inform our projections of Omicron-infection prevalence.
+The FDA assumed in its most likely scenario (as well as Scenarios 5 and 6) that case and hospitalization rates would remain constant at peak 2021 Omicron levels for 5 months.
 Throughout the pandemic, the number of COVID infections has always been significantly larger than the corresponding number of reported cases.
 For the period February 2020 through September 2021, the CDC estimated the number of infections to be four times the corresponding number of reported cases [31].
-Moreover, the case-to-infection multiplier likely increased (e.g.
-owing to increased home testing) during the Omicron-dominant period in the U.S., which started mid-December 2021.
+Moreover, the case-to-infection multiplier likely increased (e.g., owing to increased home testing [^3] See [40]) during the Omicron-dominant period in the U.S., which started mid-December 2021.
 If we conservatively assume that the CDC’s 4:1 infection:case ratio holds over the evaluation period, we find the FDA’s evaluation-period case-count projection of 274,947 per 1 million unvaccinated 18–25-year-old males (Scenarios 1, 5, 6) corresponds to 4 · 274,947 = 1,099,788 infections, well over 1 million.
 Rounding down, we deem the FDA’s case-total assumption to be equivalent to the assumption that among one million unvaccinated males, there would be, over 5 months, 1,000,000 infections.
-Having well over one million infections in a group of 1 million 18–25 year-olds over 5 months is implausible.
-Multiple infections over a five-month period would be unlikely given the high protection from reinfection resulting from recovery from a recent infection, and any reinfections would be offset by individuals who did not become infected at all (e.g., owing to prior-infection protection, limited exposure to infectious persons, or, possibly, genetic factors [37]).
-[^4] The preprint [37], posted 6/10/21, was later published [38].
+Having well over one million infections in a group of 1 million 18–25-year-olds over 5 months is implausible.
+Multiple infections over a 5-month period would be unlikely given the high protection from reinfection resulting from recovery from a recent infection, and any reinfections would be offset by individuals who did not become infected at all (e.g., owing to prior-infection protection, limited exposure to infectious persons, or, possibly, genetic factors [41]).
+[^4] The preprint [41], posted 10 June 2021, was later published [42].
 Even direct inoculation with SARS-CoV-2 does not necessarily produce infection: in the challenge study [43] of the 36 inoculated, 16 “remained uninfected.” Thus, for 3 of the FDA’s 5 Omicron-based scenarios, including its most likely scenario, the FDA essentially assumed that every unvaccinated male 18–25 would be infected by COVID over the evaluation period.
-2.3. The FDA’s mRNA-1273 Risks Assessment
-The FDA relied on its Biologics Effectiveness and Safety (BEST) System for the VAM/P case rates appearing in Table 1 above and used the CDC’s Vaccine Safety Datalink (VSD) system to estimate the percentage of VAM/P cases requiring hospitalization [11, Section 2.3.2].
-VAM/P-case rates vary considerably depending on the case definition employed and the extent of stratification of risk, based on age, sex, vaccine manufacturer and dose, with most cases occurring postdose 2 [9].
-For its mRNA-1273 assessment, the FDA four BEST-System data partners reported, with wide confidence intervals, the following myo/pericarditis incidence rates for 18–25-year-old males during the 7-day interval following dose 2: approximately 35, 70, 146, and 256 cases per million second doses [40, p, 12].
-The FDA’s BEST System uses health-claims data not validated by a complete review of patients’ medical charts [11, Section 4], and has “inherent limitations, such as small sample sizes and imperfect sensitivity of ICD-10 codes to identify these rare outcomes” [10, Section 4].
+In contrast, for our most likely scenario, we project COVID incidence over the evaluation period to be twice that of the 2nd COVID wave in the U.S.
+(1 October 2020–28 February 2021), arriving at an estimate that 45.6% of unvaccinated, COVID-naïve 18–25 year-old males in the U.S.
+would be infected by Omicron over the evaluation period—see Section S1 of Supplement S1.555Based on seroprevalence and modeling information, we estimate the actual evaluation-period infected rate was less than 50% [Supplement S2, Section S2.3].
+1.4. The FDA’s mRNA-1273 Risk Assessment
+The FDA relied on its Biologics Effectiveness and Safety (BEST) System for the VAM/P-case rates appearing in Table 1 above and used the CDC’s Vaccine Safety Datalink (VSD) system to estimate the percentage of cases requiring hospitalization [13, Section 2.3.2].
+VAM/P-case rates vary considerably depending on the case definition employed and the extent of stratification of risk, based on age, sex, vaccine manufacturer, and dose, with most cases occurring postdose 2 [9].
+For its mRNA-1273 assessment, the FDA’s four BEST-System data partners reported, with wide confidence intervals, the following myo/pericarditis incidence rates for 18–25-year-old males during the 7-day interval following dose 2: approximately 35, 70, 146, and 256 cases per million 2nd doses [44, p, 12].
+The FDA’s BEST System uses health-claims data not validated by a complete review of patients’ medical charts [13, Section 4] and has “inherent limitations, such as small sample sizes and imperfect sensitivity of ICD-10 codes to identify these rare outcomes” [12, Section 4].
 We rely on four sources of data to estimate the VAM/P rate for dose 2 among 18–25-year-old males: (i) Buchan et al.
-[41] who apply the Brighton Collaboration’s case definition, (ii) Patone et al.
-[42] who consider only hospitalized cases of myocarditis, (iii) Sharff et al.
-[43] who apply the CDC’s case definition [44], and (iv) the FDA (with VSD-based adjustments).
-The FDA underestimated mRNA-1273 VAM/P risk:
-$\bullet$ The FDA omitted dose-one VAM/P events.
+[14], who apply the Brighton Collaboration’s case definition, (ii) Patone et al.
+[15], who consider only hospitalized cases of myocarditis, (iii) Sharff et al.
+[16], who apply the CDC’s case definition [45], and (iv) the FDA (with VSD-based adjustments).
+How our mRNA-1273 VAM/P-risk assessment differs from the FDA’s:
+$\bullet$ We include dose-one VAM/P events.
 For its mRNA-1273 analysis [13], the FDA computed the evaluation-period benefits of two doses of mRNA-1273 without accounting for the VAM/P risk associated with dose 1, potentially biasing the FDA’s analysis in favor of vaccination.
-In contrast, for BNT162b2-related VAM/P, the FDA added excess myo/pericarditis cases postdose 1 and postdose 2 to approximate the population-level risk of completing 1 million two-dose vaccination series [10, Section 2.3.1].
+In contrast, for BNT162b2-related VAM/P, the FDA added excess myo/pericarditis cases postdose 1 and postdose 2 to approximate the population-level risk of completing 1 million two-dose vaccination series [12, Section 2.3.1].
 Hence, for BNT162b2, the FDA did account for 1st-dose VAM/P risk but did not assess the benefits of dose 1 alone during the interval starting with dose 1 and ending at the point of full vaccination, potentially biasing the analysis against vaccination.
-Because adverse events of vaccination other than VAM/P were disregarded by the FDA and mRNA vaccination might have unknown long-term adverse health impacts, we believe that even in a benefit-risk analysis assessing benefits only postdose 2, the choice to include 1st-dose VAM/P risk is appropriate, being consistent with the first-do-no-harm principle, the FDA’s choice in its BNT162b2 assessment, and the FDA’s choice in the benefit-risk analysis of its BLA (Biologics License Application) Clinical Review Memorandum for mRNA-1273—see Section 2.4.
-Thus, we include 1st-dose VAM/P risk for mRNA-1273 in our main analysis.
-In Supplement 2, we conduct a sensitivity analysis showing that accounting for dose 1 benefits does not substantially change our benefit-risk conclusions.
-$\bullet$ The FDA disregarded evidence that VAM/P risk for mRNA-1273 was over twice that for BNT162b2.
-In a “Review Memorandum Addendum” [45], dated 11/19/21, the FDA describes multiple sources of evidence that VAM/P risk for mRNA-1273 is significantly greater than that for BNT162b2, including reports from regulatory agencies of five countries, data from Ontario later analyzed by Buchan et al.
+Because we, like the FDA, do not account for adverse events of vaccination other than VAM/P, and mRNA vaccination might have unknown long-term adverse health impacts, we believe that even in a benefit–risk analysis assessing benefits only postdose 2, the choice to include 1st-dose VAM/P risk is appropriate, being consistent with the first-do-no-harm principle, the FDA’s choice in its BNT162b2 assessment, and the FDA’s choice in the benefit–risk analysis of its BLA (Biologics License Application) Clinical Review Memorandum for mRNA-1273—see Section 1.5 below.
+Thus, we include the 1st-dose VAM/P risk for mRNA-1273 in our main analysis.
+See Section S2 of Supplement S2 for a sensitivity analysis showing that accounting for dose 1’s benefits does not substantially change our benefit–risk conclusions.
+$\bullet$ We incorporate evidence that VAM/P risk for mRNA-1273 was over twice that for BNT162b2.
+In a “Review Memorandum Addendum” [46], dated 19 November 2021, the FDA describes evidence that VAM/P risk for mRNA-1273 is significantly greater than that for BNT162b2, including regulatory-agency reports from five countries, data from Ontario later analyzed by Buchan et al.
 [14], and data from the CDC: “[T]he most recent data presented by the Center for Disease Control and Prevention (CDC) analyzed from the Vaccine Safety Datalink (VSD) indicates a higher rate of myocarditis with the Moderna COVID-19 Vaccine than the Pfizer-BioNTech COVID-19 Vaccine ….”
@@
 Adjusted-rate ratios for 18–39-year-old males from the table on page 27 of [17] suggest the 2nd-dose VAM/P rate for mRNA-1273 is 2.26 times that of BNT162b2.
-Data from other sources suggests higher rate ratios e.g,, a 2nd-dose rate ratio of 5.1 was reported by Buchan et al.
-for males 18–24 [41, p.
+Data from other sources suggests higher rate ratios, e.g., a 2nd-dose rate ratio of 5.1 was reported by Buchan et al.
+for males 18–24 [14, p.
 5] and of 5.2 was reported by Husby et al.
@@
 A study by Patone et al.
-of myocarditis cases requiring hospitalization within a 28-day risk interval after a vaccine dose found among males under 40 a rate of 101 excess myocarditis hospitalizations per million 2nd doses of mRNA-1273 and 12 per million 1st doses, with corresponding BNT162b2 rates being 12 per million 2nd doses and 3 per million 1st doses [42, p.
+of myocarditis cases requiring hospitalization within a 28-day risk interval after a vaccine dose found among males under 40 a rate of 101 excess myocarditis hospitalizations per million 2nd doses of mRNA-1273 and 12 per million 1st doses, with corresponding BNT162b2 rates being 12 per million 2nd doses and 3 per million 1st doses [15, p.
 4].
-Based on “preliminary findings” [45, p.
-2], having “large uncertainty” owing to “small numbers of observed events” [40, p.
-16], as well as omitting dose-1 cases, the FDA suggests that for 18–25-year-old males VAM/P risk for mRNA-1273 is comparable to that for BNT162b2: 128 cases per million 2nd doses of mRNA-1273 [11, Table 4] vs.
-131 cases per million full vaccinations of BNT162b2 (among 18–24-year-old males) [10, Table 3], with corresponding hospitalization projections of 110 per million 2nd doses of mRNA-1273 and 131 per million full BNT162b2 vaccinations.
+Based on “preliminary findings” [46, p.
+2], having “large uncertainty” owing to “small numbers of observed events” [44, p.
+16], as well as omitting dose-1 cases, the FDA suggests that for 18–25-year-old males, VAM/P risk for mRNA-1273 is comparable to that for BNT162b2: 128 cases per million 2nd doses of mRNA-1273 [13, Table 4] vs.
+131 cases per million full vaccinations of BNT162b2 (among 18–24-year-old males) [12, Table 3], with corresponding hospitalization projections of 110 per million 2nd doses of mRNA-1273 and 131 per million full BNT162b2 vaccinations.
 Throughout this paper, we assume that VAM/P rates for males 18–24 well approximate those for males 18–25.
-Using the rate ratios and BNT162b2 VAM/P-risk assessment discussed above, we compute in Section 6.1 of Supplement 1 a VAM/P hospitalization rate of 224 per million 2nd doses of mRNA-1273 for males 18–25 years old.
+Using the rate ratios and BNT162b2 VAM/P-risk assessment discussed above, we compute a VAM/P-hospitalization rate of 224 per million 2nd doses of mRNA-1273 for males 18–25 years old [Supplement S1, Section S6.1].
 The FDA’s BEST System also provided evidence that VAM/P risk for mRNA-1273 might be over twice that for BNT162b2.
-An FDA briefing document for a meeting held 10/14/2021 indicates that one of the agency’s BEST datapartners reported a postvaccination myo/pericarditis incidence rate of 283.7 (95% CI 145.2-573.5) cases per million full mRNA-1273 vaccinations (7 day risk window after each dose) [48, Section 2.3.3].
-Applying a background rate of 4 cases per million per week for 18–25-year-old males [Supplement 1, Section 6.3] to each dose, we find the VAM/P case rate corresponding to 283.7 is 275.7 per million full vaccinations.
+An FDA briefing document for a meeting held 14 October 2021 indicates that one of the agency’s BEST data partners reported a postvaccination myo/pericarditis incidence rate of 283.7 (95% CI 145.2–573.5) cases per million full mRNA-1273 vaccinations (7 day risk window after each dose) [48, Section 2.3.3].
+Applying a background rate of four cases per million per week for 18–25-year-old males for each dose [Supplement S1, Section S6.3], we find the VAM/P-case rate corresponding to 283.7 is 275.7 per million full vaccinations.
 Assuming a case-hospitalization rate of 100%, as in FDA’s BNT162b assessment, we obtain 275.7 VAM/P hospitalizations per million full mRNA-1273 vaccinations vs.
 131 per million full BNT162b vaccinations.
-$\bullet$ The FDA’s meta-analysis of its BEST-System data [11, Section 2.3.1] yielded VAM/P-rate estimates substantially lower than those of major international studies by Buchan et al.
-[41] and Patone et al.
-[42], as well as a U.S.
-study by Sharff et al.
-[43] [^5] The conclusion of [43] includes “We identified additional valid cases of myopericarditis following an mRNA vaccination that would be missed by the VSD’s search algorithm, which depends on select hospital discharge diagnosis codes.
-The true incidence of myopericarditis is markedly higher than the incidence reported to US advisory committees.” .
+As we have indicated, we rely on four sources of data for our VAM/P-rate estimates for dose 2 of mRNA-1273: the FDA with VSD-based adjustments, as well as studies by Buchan et al.
+[14], Patone et al.
+[15], and Sharff et al.
+[16] (from, respectively, Ontario, England, and the U.S.).
+The conclusion of [16], which includes the following, is noteworthy: [^6] The preprint [16] and the later publication [4] share the same conclusion.
+We identified additional valid cases of myopericarditis following an mRNA vaccination that would be missed by the VSD’s search algorithm $\ldots$.
+The true incidence of myopericarditis is markedly higher than the incidence reported to US advisory committees.
 In Table 2 below, we compare the FDA’s projected mRNA-1273 VAM/P case and hospitalization rates to those of the other sources discussed above.
-(We have excluded VAM/P-rate sources, such as [49], depending on strictly passive surveillance—see Section 6.8 of Supplement 1.)
+We exclude VAM/P-rate sources, such as [49], depending on strictly passive surveillance—see Section S6.8 of Supplement S1.
 Table 2. Estimated VAM/P Events Per Million 2nd Doses of mRNA-1273 Among Males 18–24 or 18–25 years old.a Based on Data Available Before 1 January 2022
@@
 (Scenarios 5 & 6)
-[11, Table 2]
+[13, Table 2]
 Buchan et al.b
@@
 21
-[40, p. 14]
+[44, p. 14]
 55
-[41, Table 4]
+[14, Table 4]
   23
 7
-[43, Table 1]
+[16, Table 1]
   22
@@
 107
-a For discussions of the data in columns 3–6, see Section 6 of Supplement 1; in particular, data in column 4 are estimates based on findings from [42] for males under 40, adjusted using an age-based myocarditis-incidence ratio as well as pericarditis data from [41].
+a For discussions of the data in columns 3–6, see Section S6 of Supplement S1; in particular, data in column 4 are estimates based on findings from [15] for males under 40, adjusted using an age-based myocarditis-incidence ratio as well as pericarditis data from [14].
 To obtain the column-5 VAM/P case rate of 525 per million we applied the background myo/pericarditis rate suggested by [14].
 b Buchan et al.
-applied the Brighton Collaboration’s case definition for myo/pericarditisx, which reduced the reported 417 post-mRNA-vaccination study-period cases of myo/pericarditis to 297 [41, p.
+applied the Brighton Collaboration’s case definition for myo/pericarditisx, which reduced the reported 417 post-mRNA-vaccination study-period cases of myo/pericarditis to 297 [14, p.
 5].
@@
 For homologous dosing (1st dose also mRNA-1273), Table 3 of [14] indicates the rate rises to at least 320 per million when the 2nd dose is taken less than 8 weeks after the 1st (with 4–6 weeks being the recommended interval [50]).
-d For its BNT162b2 assessment, the FDA assumed all VAM/P cases were hospitalized [10, Section 2.3.2.2].
-For its mRNA-1273 assessment the FDA assumed VAM/P case-hospitalization rates decrease with age, starting with 86% for males 18–25 and ending with 77% for males 36 and over [11, Section 2.3.2.2].
+d For its BNT162b2 assessment, the FDA assumed all VAM/P cases were hospitalized [12, Section 2.3.2.2].
+For its mRNA-1273 assessment the FDA assumed VAM/P case-hospitalization rates decrease with age, starting with 86% for males 18–25 and ending with 77% for males 36 and over [13, Section 2.3.2.2].
 We apply the 86% rate in columns 4–6 (which, for column 5, matches the case hospitalization rate: 6 of the 7 cases were hospitalized).
 For column 3, Buchan et al.
-report a case-hospitalization rate postdose 2 of an mRNA vaccine (all ages, males & females) of 79.2% ($\approx 164/207\times 100$%) [41, Table 1]; we expect the rate to be higher for the age range 18–24, but, conservatively, choose a hospitalization rate of 80%.
+report a case-hospitalization rate postdose 2 of an mRNA vaccine (all ages, males & females) of 79.2% ($\approx 164/207\times 100$%) [14, Table 1]; we expect the rate to be higher for the age range 18–24, but, conservatively, choose a hospitalization rate of 80%.
 $301\approx 302\cdot\frac{55}{107}.+269\cdot\frac{23}{107}+525\cdot\frac{7}{107}+260\cdot\frac{22}{107}$           $\ddagger$ $250\approx 242\cdot\frac{55}{107}.+231\cdot\frac{23}{107}+452\cdot\frac{7}{107}+224\cdot\frac{22}{107}$
 1.5. A Seventh Benefit Risk Scenario for mRNA-1273
-The FDA’s BLA Clinical Review Memorandum [51] for mRNA-1273, completed 1/28/2022, describes a benefit-risk scenario not matching any scenario of the FDA’s published benefit-risk assessment of mRNA-1273 [11].
-The projections of the BLA-Memo Scenario [51, Section 4.7] are described in the initial row of data of Table 4 below and compared to the FDA’s scenarios of [11, Table 2] in particular to Scenario 1, the most likely scenario.
+The FDA’s BLA Clinical Review Memorandum [51] for mRNA-1273, completed 28 January 2022, describes a benefit–risk scenario not matching any scenario of the FDA’s published benefit–risk assessment of mRNA-1273 [13].
+The projections of the BLA-Memo Scenario [51, Section 4.7] are described in the initial row of data of Table 3 below and compared to the FDA’s scenarios of [13, Table 2] in particular to Scenario 1, the most likely scenario.
 Differences include the BLA-Memo Scenario’s accounting for 1st-dose VAM/P risk and projecting 1 million full vaccinations would cause 47 ICU admissions (vs.
 0 admissions).
-[^6] Subsequent analyses provide further evidence that the FDA may underestimate VAM/P severity in [11]; e.g., a meta-analysis [52] of 23 studies including 854 patients ages 12–20 years found a 92.6% case-hospitalization rate for postvaccination myo/pericarditis and a 23.2% ICU-admission rate.
+[^7] Subsequent analyses provide further evidence that the scenarios of [13] may underestimate VAM/P severity; e.g., a meta-analysis [52] of 23 studies including 854 patients ages 12–20 years found a 92.6% case-hospitalization rate for postvaccination myo/pericarditis and a 23.2% ICU-admission rate.
 Table 3. FDA scenarios projecting benefit-risk outcomes per million 18–25-year-old males vaccinated with two primary series doses of mRNA-1273.*
@@
 6 82,484 4,766 1,144 51 247 207 0 0
-* For Scenarios 1–6, the FDA estimated VAM/P risk only for dose 2. Projections for the FDA’a Scenarios 1–6 are from [11, Table 2]. BLA-Memo-scenario projections are from [51, Section 4.7].”
-Consistent with the BLA Memo Scenario, the FDA’s 1/31/22 press release [12] announcing the full approval of the mRNA-1273 vaccine indicates some mRNA-1273 VAM/P cases do require ICU care.
-Both the BLA Memo [51, Section 4.7] and the FDA’s mRNA-1273 benefit-risk assessment [11, Section 2.3.2.2] attribute their contradictory ICU risk estimates to VSD data.
-The BLA Memo lacks a specific reference, while the FDA in [11] cites a source of VSD data [53] with collection cutoff point 8/21/2021, over four months before completion of the FDA’s mRNA-1273 assessment.
+* For Scenarios 1–6, the FDA estimated VAM/P risk only for dose 2. Projections for the FDA’a Scenarios 1–6 are from [13, Table 2]. BLA-Memo-scenario projections are from [51, Section 4.7].
+Consistent with the BLA-Memo Scenario, the FDA’s 31 January 2022 press release [18] announcing the full approval of mRNA-1273 indicates some mRNA-1273 VAM/P cases require ICU care.
+To support its estimates of VAM/P-hospitalization and ICU-admission risk, the FDA in [13] cites a source of VSD data [53] with a collection-cutoff point of 21 August 2021, over four months before completion of its mRNA-1273 assessment.
 The VSD data from [53] cited by the FDA [13] is scant, reporting only 21 postvaccination myo/pericarditis cases for those 18–29 [53, p.
 16].
-A comparison of Buchan et al.’s myo/pericarditis data from Ontario [41, Table 4, collected 6/1/21– 9/4/21] to VSD data through 10/9/21 [46, pp.
-8, 16] shows that, among males and females ages 18–39, the number of myo/pericarditis cases per million second doses of mRNA-1273 detected by Public Health Ontario is 3.76 times the number detected by the CDC’s VSD system [Supplement 2, Appendix 3].
-3. Reanalysis: Methods
-Recall that for its case and hospitalization endpoints, the FDA’s benefit-risk model has inputs $C$ and $H$ (the numbers of evaluation-period COVID cases and hospitalizations, respectively, in a test population of one million unvaccinated males) as well as VE and VEH (the fractions of these events prevented by full vaccination).
+A comparison of Buchan et al.’s myo/pericarditis data from Ontario [14, Table 4, collected 1 June 2021–4 September 2021] to VSD data through 19 October 2021 [17, pp.
+8, 16] shows that, among males and females ages 18–39, the number of myo/pericarditis cases per million 2nd doses of mRNA-1273 detected by Public Health Ontario is 3.76 times the number detected by the CDC’s VSD system [Supplement S2, Appendix S3].
+2. Methods
+Recall that for its cases and hospitalizations endpoints, the FDA’s benefit–risk model has inputs $C$ and $H$ (the numbers of anticipated evaluation-period COVID cases and hospitalizations, respectively, in a test population of one million unvaccinated males) as well as $VE$ and $VEH$ (the fractions of these events prevented by full vaccination).
 These model inputs are combined to project the cases-prevented and hospitalizations-prevented benefits of vaccinating the test population, and these benefits are compared to the model’s corresponding risk inputs, the numbers of VAM/P cases and hospitalizations projected for the 2nd mRNA-1273 dose.
@@
 Our model extends the FDA’s by not only incorporating prior-infection protection but also by being essentially equivalent to the FDA’s model if prior infection is assumed to provide no protection.
-Because the FDA relied on CDC data to supply projected values for $C$ and $H$, the FDA’s model does not explicitly link cases to hospitalizations and similarly doesn’t explicitly link VE and VEH.
+Because the FDA relied on CDC data to supply projected values for $C$ and $H$, the FDA’s model does not explicitly link cases to hospitalizations and similarly does not explicitly link $VE$ and $VEH$.
 In contrast, our model does provide these links.
-For example, it links infections, a key model input replacing reported cases, to hospitalizations via another key input—the infection hospitalization rate IHR.
-Additional input variables are described in Section 3.1 below and summarized in Table 5.
+For example, it links infections, a key model input replacing reported cases, to hospitalizations via another key input—the infection-hospitalization rate IHR.
+Additional input variables are described below.
 Model inputs are combined to project the hospitalizations-prevented benefit of vaccinating a test population, and this projected benefit is compared to the model’s risk input, the projected number of VAM/P hospitalizations expected to occur in the course of achieving a fully vaccinated test population.
@@
 Test population.
-Our model is based on a hypothetical test population of 1 million males 18–25.
-We assume that the test population is representative, as of 1/1/2022, of all males in this age group in terms of: distribution of ages within the 18–25 range, comorbidities (e.g., obesity, cardiovascular disease, diabetes), and the percentage having had a prior COVID-19 infection.
-We project the number of hospitalizations for Omicron infection that will occur in this test population over the evaluation period 1/1/22 – 5/31/22 (a) assuming all members remain unvaccinated throughout the period, and (b) assuming all members of the population are vaccinated throughout the period (all fully vaccinated, 14 days past dose 2, on 1/1/22).
+Our model is based on a hypothetical test population of 1 million males aged 18–25.
+We assume that the test population is representative, as of 1 January 2022 of all males in this age group in terms of: distribution of ages within the 18–25 range, comorbidities (e.g., obesity, cardiovascular disease, diabetes), and the percentage having had a prior COVID-19 infection.
+We project the number of hospitalizations for Omicron infection that will occur in this test population over the evaluation period 1 January 2022–31 May 2022 (a) assuming all members remain unvaccinated throughout the period, and (b) assuming all members of the population are vaccinated throughout the period (all fully vaccinated, 14 days past dose 2, on 1 January 2022).
 The difference—the projection of (a) minus that of (b)—is the hospitalizations-prevented benefit of vaccinating the test population.
-We call a test population consisting of infection naïve unvaccinated persons a baseline test population.
+We call a test population consisting of infection-naïve unvaccinated persons a baseline test population.
 2.1. Model Inputs
@@
 $H_{r}$ is the fraction of a population of unvaccinated individuals experiencing a primary COVID infection that will progress to severe disease requiring hospitalization.
-We emphasize that $H_{r}$ is a sex-specific rate.
-The FDA provides data in Table 2 of [10] indicating that for the 18–24 age-range $H_{r}$ for males is significantly less that that for females.
-[^7] The data of Table 1 of [54] is consistent with that of [10, Table 2].
-Prior-infection fraction $F_{pi}$. This is the fraction of the test population having had a COVID-19 infection prior to the evaluation period.
+Prior-infection fraction $F_{pi}$. This is the fraction of the test population having had a COVID-19 infection before the start of the evaluation period.
 Vaccine effectiveness against infection $E_{v}$.
@@
 These are the fractions by which the baseline infection rate $I_{r}$ is reduced due, respectively, to the protection afforded by prior infection, or to both vaccination and prior infection (hybrid protection).
-Hospitalization risk reduction $HRR_{v}$, $HRR_{pi}$, and $HRR_{h}$.
-These are the fractions by which the hospitalization rate $H_{r}$ is reduced for populations that are protected respectively by vaccination, prior infection, or both (hybrid protection).
+Hospitalization-risk reduction $HRR_{v}$, $HRR_{pi}$, and $HRR_{h}$.
+These are the fractions by which the hospitalization rate $H_{r}$ is reduced for populations that are protected, respectively, by vaccination, prior infection, or both (hybrid protection).
 VAM/P hospitalizations $H_{\rm VAM/P}$.
 Our final model input $H_{\rm VAM/P}$ is the projected number of VAM/P hospitalizations expected to occur in attaining a test population of 1 million fully mRNA-1273 vaccinated 18–25-year-old males.
-To vaccinate fully 1 million individuals with an mRNA COVID-vaccine, more than 2 million doses must be administered (assuming, e.g., that those who experience myo/pericarditis related to dose 1 do not take dose 2), and the adverse-events associated with each of the over 2 million doses reflect the cost to the population of obtaining the benefits of 1 million full vaccinations.
-In Section 2.3, based on data available before 1/1/22, we obtained an estimated VAM/P hospitalization rate of 250 per million 2nd doses of mRNA-1273 given to 18–25-year-old males.
-In Section 6.7 of Supplement 1, we discuss VAM/P-hospitalization risk among 18–25-year-old males for dose 1 of mRNA-1273, justifying a choice of 18 VAM/P hospitalizations per million 1st doses, which is the 1st-dose rate of the FDA’s BLA-Memo Scenario—compare the first two rows of Table 4 above.
-Thus, for all our scenarios, we set $H_{\rm VAM/P}=268$.
-[^8] In Supplement 1 (Appendix 2, Section 2.4), we show that our VAM/P hospitalization-rate estimate of 268 per million full mRNA-1273 vaccinations among 18–25-year-old males is consistent with VAM/P data of three major studies [5, 7, 8] appearing in 2022 but available only after 1/22/22, showing the data suggests a VAM/P hospitalization rate between 240 and 277 per million full, homologous mRNA-1273 vaccinations among males 18–25.
+To fully vaccinate 1 million individuals with an mRNA COVID vaccine, more than 2 million doses must be administered (assuming, e.g., that those who experience myo/pericarditis related to dose 1 do not take dose 2), and the adverse events associated with each of the over 2 million doses reflect the cost to the population of obtaining the benefits of 1 million full vaccinations.
 2.2. Combining Model Inputs to Produce Outputs
-Recall that we are considering a test population of 1 million males 18–25 over the evaluation period 1/1/22 – 5/31/22.
+Recall that we are considering a test population of 1 million males 18–25 over the evaluation period 1 January 2022–31 May 2022.
 Let $U$ denote the expected number of Omicron hospitalizations occurring in this test population over the 5-month evaluation period assuming its members remain unvaccinated throughout the period.
 Let $V$ denote the corresponding number of expected hospitalizations assuming the population is vaccinated throughout the period.
-The difference UVU-V is the hospitalizations-prevented benefit of vaccination for the test population, and the outputs $U$ and $V$ can be written in terms of the model inputs defined in the preceding section (thus, our model is deterministic).
+The difference UVU-V is the hospitalizations-prevented benefit of vaccination for the test population, and the outputs $U$ and $V$ can be written in terms of the model inputs defined in the preceding section (thus, like the FDA’s, our model is deterministic).
 In our modeling, we consider only the possibility of hospitalizations associated with test-population members’ first infections during the evaluation period.
@@
 Then, in this baseline population, we expect $(1{\rm m})\cdot I_{r}$ infections to develop over 5 months, resulting in $(1{\rm m})\cdot I_{r}\cdot H_{r}$ hospitalizations, where $1{\rm m}=$ 1 million.
-Next suppose a fraction $F_{pi}$ of the unvaccinated test population has had a COVID-19 infection prior to the start of the evaluation period.
+Next, suppose a fraction $F_{pi}$ of the unvaccinated test population has had a COVID-19 infection prior to the start of the evaluation period.
 We must account for the protection against infection and hospitalization afforded by prior infection.
@@
 Among the remaining $F_{pi}\cdot(1{\rm m})$ members of the test population with prior infection, we expect $F_{pi}\cdot(1{\rm m})\cdot I_{r}\cdot(1-E_{pi})$ infections resulting in $F_{pi}\cdot(1{\rm m})\cdot I_{r}\cdot(1-E_{pi})\cdot H_{r}\cdot(1-HRR_{pi})$ hospitalizations.
-Thus the total number of hospitalizations expected in the unvaccinated test population is
+Thus, the total number of hospitalizations expected in the unvaccinated test population is
 (2.2.1) $U=(1-F_{pi})(1{\rm m})I_{r}H_{r}+F_{pi}(1{\rm m})I_{r}(1-E_{pi})H_{r}(1-HRR_{pi}).$
 Deriving $V$.
-Next we assume that every individual in the test population is fully vaccinated.
-Since all members of this test population are vaccinated, those with prior infection have hybrid protection.
-Once again we reason in two stages, first accounting for the protection against infection, then for the protection against hospitalization for infected individuals.
+Next, we assume that every individual in the test population is fully vaccinated.
+Since all members are vaccinated, those with prior infection have hybrid protection.
+Once again, we reason in two stages, first accounting for the protection against infection, then for the protection against hospitalization for infected individuals.
 Once again, we assume a fraction $F_{pi}$ of the test population has had a prior COVID-19 infection.
 Then, in the fully vaccinated test population, the expected number of hospitalizations is $(1-F_{pi})\cdot(1{\rm m})\cdot I_{r}\cdot(1-E_{v})\cdot H_{r}\cdot(1-HRR_{v})$ for those without prior infection and $F_{pi}\cdot(1{\rm m})\cdot I_{r}\cdot(1-E_{h})\cdot H_{r}\cdot(1-HRR_{h})$ for those with hybrid protection.
-So the total number of hospitalizations in the population is
+So, the total number of hospitalizations in the population is
 (2.2.2) $V=(1-F_{pi})(1{\rm m})I_{r}(1-E_{v})H_{r}(1-HRR_{v})+F_{pi}(1{\rm m})I_{r}(1-E_{h})H_{r}(1-HRR_{h}).$
@@
 […]
-We remark that disregarding the protection afforded by prior infection, as the FDA did, is equivalent in our model to assuming that all persons in the test population are infection-naïve, i.e.
-that $F_{pi}=0$.
+We remark that disregarding the protection afforded by prior infection, as the FDA did, is equivalent in our model to assuming that all persons in the test population are infection-naïve, i.e., that $F_{pi}=0$.
 With $F_{pi}=0$ in Equation (2.2.3), its right-hand side is
 (2.2.4) $\left[\text{Expected Baseline Hospitalizations}\right]\left(1-(1-E_{v})(1-HRR_{v})\right),$
-[…]
+which is essentially the FDA’s model $H\cdot VEH$ for hospitalizations prevented.
+The quantity $1-(1-E_{v})\cdot(1-HRR_{v})$ of (4) corresponds to the $VEH$ formula on page 10 of [54], which is the reference the FDA used in [13] for its $VEH$ estimate of 0.72.
+Thus, our model for hospitalizations-prevented UVU-V represents a natural extension of the FDA’s model.
 2.3. Estimating Values for Model Inputs
 Like the FDA, we account for uncertainties in model inputs by adopting a scenario-based assessment.
-The FDA’s model-input assumptions for its six scenarios of [11] are summarized in Table 1 of Section 2.
-Observe that the first three scenarios differ only in their COVID-incidence assumptions, with Scenario 1 representing the FDA’s most likely scenario.
-Our first three scenarios, Scenarios A–C, are reanalyses, respectively, of the FDA’s Scenarios 1–3 because our COVID incidence assumptions are derived from the FDA’s corresponding assumptions.
-Based on information available at the time of the FDA’s mRNA-1273 analysis, we consider the COVID incidence assumption of our Scenario D more likely to be accurate than those of Scenarios A-C; also, for Scenario D our values for other inputs represent our best estimates.
+The FDA’s model-input assumptions for its six scenarios of [13] are summarized in Table 1 above.
+Observe that the first three scenarios differ only in their COVID incidence assumptions, with Scenario 1 representing the FDA’s most likely scenario.
+Our first three scenarios, Scenarios A–C, are reanalyses, respectively, of the FDA’s Scenarios 1–3 because our COVID incidence assumptions are derived from the FDA’s.
+Based on information available at the time of the FDA’s mRNA-1273 analysis, we consider the COVID incidence assumption of our Scenario D more likely to be accurate than those of Scenarios A–C; also, for Scenario D, our values for other inputs represent our best estimates.
 Thus, Scenario D is our most likely scenario.
-In Scenario E, our final scenario, we assume that prior infection and mRNA-1273 vaccination provide equivalent protection against Omicron, which, for reasons discussed in Section 2.2, we believe to be the minimum reasonable assumption of prior-infection’s protectiveness relative to vaccination’s protectiveness.
-Table 5 below summarizes model inputs with footnotes providing some comments on the estimation of their values—see Supplement 1 for details.
+In Scenario E, our final scenario, we assume that prior infection and mRNA-1273 vaccination provide equivalent protection against Omicron, which, for reasons discussed Section 1.3, we believe to be the minimum reasonable assumption of prior infection’s protectiveness relative to vaccination’s protectiveness.
+Model inputs and outputs for our scenarios are described in the next section, while the methods we use to estimate input values are described below and summarized in Table 4.
+Our estimates are based on data available to the FDA before 22 January 2022.
+Estimating $I_{r}$.
+As we noted in Section 1.3, the FDA’s COVID-case-rate assumption for its most likely scenario, Scenario 1, corresponds to the assumption that essentially all unprotected (i.e., unvaccinated) 18–25-year-old males become infected over the 5-month evaluation period.
+Thus, in Scenario A, our reanalysis of Scenario 1, we too assume all unprotected persons in the test population become infected.
+However, in our modeling, the unprotected are the infection-naïve unvaccinated.
+Assuming all such persons become infected over 5 months is equivalent to setting $I_{r}=1$, our Scenario-A value for $I_{r}$.
+As discussed in Section S1 of Supplement S1, for our Scenarios B and C, values of $I_{r}$ are determined by the FDA’s COVID-incidence assumptions for its Scenarios 2 and 3, respectively, while for our Scenarios D and E, $I_{r}$ is derived from the assumption that over the evaluation period the increase in COVID-infection level among 18–25-year-old males is twice that of the general population during the second COVID wave in the U.S.
+(1 October 2020–28 February 2021).
+Estimating $H_{r}$.
+We estimate $H_{r}$ using an IHR model [55] developed via a seroprevalence-based meta-analysis anchored to the ancestral strain of SARS-CoV-2 and available as a preprint in October 2021 [39].
+Applying the model, we obtain a “with COVID” IHR of 0.37% for the infection-naïve unvaccinated of ages 18–25 and reduce this rate assuming 40% of hospitalizations attributed to COVID are incidental (see Section 1.3), arriving at an IHR of $0.22\%\approx 0.6\cdot 0.37\%$ (for infection-naïve unvaccinated 18–25-year-olds).
+The 0.22% rate for the general population of 18–25-year-olds in the U.S.
+(males and females) is consistent with U.S.
+military COVID-hospitalization data [56].
+See Supplement S1 for further details concerning our derivation of the IHR value $0.22\%$.
+In its benefit–risk assessments [12, 13] for, respectively, Pfizer’s and Moderna’s mRNA vaccines, the FDA relied on sex-specific CDC hospitalizations data.
+The case and hospitalization rates per 100,000 persons in Table 2 of [12] suggest that for the 18–24 age range, females are 2.53 times as likely as males to be hospitalized with COVID-19 and 1.21 times as likely to have a case of COVID-19.888These relationships are consistent with those of [57, Table 1] based on data from 1 March 2020 to 1 March 2022.
+In Section S2 of Supplement S1, we use hospitalization and case data from Table 2 of [12] to estimate the IHR for unvaccinated, infection-naïve males 18–25 to be $0.14\%$; thus, we set $H_{r}=0.0014$.
+This value agrees with a natural IHR estimate derived from the FDA’s BLA-Memo scenario (discussed in Section 1.5).
+For the BLA-Memo Scenario, vaccination prevents 76,362 cases and 1755 hospitalizations by providing 30% effectiveness against cases and 72% effectiveness against hospitalizations.
+Thus, the scenario assumes that $76,362/0.3=254,540$ cases among the unvaccinated would lead to $1755/0.72\approx 2438$ hospitalizations, yielding a case hospitalization rate of $\approx$0.958%.
+Applying the CDC case-to-infection multiplier of 4 and assuming a 40% incidental COVID-hospitalization rate, we arrive at the estimate $H_{r}=0.6\cdot(2438)/(4\cdot$ 254,540) $\approx$ 0.0014.
+Estimating $F_{pi}$.
+For our Scenarios A–C, we use the value $F_{pi}=0.62$, which is the average of (a) the CDC’s estimate, mentioned earlier, that 54.9% of those 18–49 years old had been infected by COVID-19 before 1 October 2021 [31] with (b) an estimate derived in Supplement S1 that 69% of unvaccinated male 18–25-year-olds had been infected by 1 January 2022.
+For Scenarios D and E, we use $F_{pi}=0.69$.
+These values of $F_{pi}$ may be underestimates because they rely on the CDC’s infected rate of 54.9% by 1 October 2021 for the general population of 18–49 year-olds, and we would expect a somewhat higher infected rate among the unvaccinated.
+Estimating $E_{v}$, $HRR_{v}$.
+Recall that the FDA used $VE=0.30$ and $VEH=0.72$ for all its Omicron-based scenarios [13, Table 1].
+For reasons discussed in Supplement S1, for all our scenarios, we use $E_{v}=0.30$ and $HRR_{v}=0.67$.
+Hence, we are assuming a $VEH$ of $1-(1-0.30)(1-0.67)\approx 0.77$, which exceeds the $VEH$ used by FDA.
+Estimating $E_{pi}$ and $E_{h}$.
+Our values $E_{pi}=0.45$ and $E_{h}=0.57$ are derived (see Supplement S1) from a December 2021 study [32], based on data from South Africa.
+[^9] As discussed in Supplement S1, our effectiveness estimates involving prior infection are consistent with those of meta-analyses [27] and [28].
+Estimating $HRR_{pi}$ and $HRR_{h}$.
+In Supplement S1, we use data from [33] (based on Omicron cases in England for the period 1 December 2021–14 December 2021) to estimate $HRR_{pi}=0.79$ and $HRR_{h}=0.88$.
+[^10] The estimates $HRR_{pi}=0.79$ and $HRR_{h}=0.88$ combined with our estimates $E_{pi}=0.45$ and $E_{h}=0.57$ yield overall estimates for protection against hospitalization—88.5% for prior infection and 94% for hybrid protection—consistent with those of meta-analyses [27] and [28].
+We note that hybrid protection’s exceeding vaccine-only protection is consistent with studies such as [58, 59].
+.
+Estimating $H_{\rm VAM/P}$.
+In Section 1.4, we combined four assessments of VAM/P risk based on data available before 1 January 2022 to obtain an estimate of 250 VAM/P hospitalizations per million 2nd doses of mRNA-1273 given to 18–25-year-old males.
+In Section S6.7 of Supplement S1, we discuss VAM/P-hospitalization risk among 18–25-year-old males for dose 1 of mRNA-1273, justifying a choice of 18 VAM/P hospitalizations per million 1st doses, which is the 1st-dose risk the FDA uses in its BLA-Memo Scenario—compare the first two rows of Table 3 above.
+Thus, for all our scenarios, we set $H_{\rm VAM/P}=268$.
+[^11] In Supplement S1 (Appendix S2, Section S2.4), we show that our VAM/P hospitalization-rate estimate of 268 per million full mRNA-1273 vaccinations among 18–25-year-old males is consistent with VAM/P data of three major studies [5, 7, 8] appearing in 2022 but available only after 22 January 2022, showing the data suggests a VAM/P-hospitalization rate between 240 and 277 per million full, homologous mRNA-1273 vaccinations among males 18–25.
 Table 4. Summary of Input Variables*
@@
 $F_{pi}$: Fraction of a test population of 1 million unvaccinated males 18–25 having had a COVID-19 infection before the start of the 5-month evaluation period $0\leq F_{pi}\leq 1$ 0.62
-$E_{pi}$: Effectiveness of prior-infection protection against COVID-19 infection $0\leq E_{pi}\leq 1$ 0.47
+$E_{pi}$: Effectiveness of prior-infection protection against COVID-19 infection $0\leq E_{pi}\leq 1$ 0.45
 $E_{v}$: Effectiveness of vaccination against COVID-19 infection $0\leq E_{v}\leq 1$ $0.3$
@@
 * Remarks.
-(i) Estimates are based on data available before 1/22/22.
-(ii) $E_{pi}$, $E_{v}$, and $E_{h}$ are relative to infection-naïve, unvaccinated persons.
-(iii) $HRR_{pi}$, $HRR_{v}$, and $HRR_{h}$ are relative to unvaccinated persons experiencing primary infections.
-(iv) Our Scenario-A value $I_{r}=1$ reflects the assumption that all infection-naïve unvaccinated males 18–25 become infected over the evaluation period, consistent with FDA Scenario-1 assumptions (see Section 2.2).
-(v) We estimate $H_{r}$ using an IHR model [55] developed via a seroprevalence-based meta-analysis anchored to the ancestral strain of SARS-CoV-2 and available as a preprint in October 2021 [56].
-We assume a 40% incidental-hospitalization rate and make sex-based adjustments based on [10, Table 2].
-(See [Supplement 1, Section 2]) (vi) $F_{pi}=0.62$ is the average of (a) the CDC’s estimate that 54.9% of those 18–49 years old had been infected by COVID-19 before 10/1/21 [25] and (b) an estimate derived in Supplement 1 that 69% of unvaccinated male 18–25 year-olds had been infected by 1/1/2022.
-(vii) For reasons discussed in Supplement 1, for all scenarios, we use $E_{v}=0.3$ and $HRR_{v}=0.67$; thus, we are assuming a VEH of $1-(1-0.30)(1-0.67)\approx 0.77$, which exceeds the FDA’s VEH (viii) Estimates for $E_{pi},E_{h},HRR_{pi}$, and $HRR_{h}$, are based on data from [26] and [27].
-$HRR_{pi}=0.79$ and $HRR_{h}=0.88$ combine with $E_{pi}=0.47$ and $E_{h}=0.57$ to yield overall estimates for protection against hospitalization—88.9% for prior infection and 94% for hybrid protection—consistent with those of meta-analyses [21] and [22].
+All measures of effectiveness of protection against infection ($E_{pi}$, $E_{v}$, and $E_{h}$) are relative to infection-naïve, unvaccinated persons.
+All hospitalization-risk reductions ($HRR_{pi}$, $HRR_{v}$, and $HRR_{h}$) are relative to unvaccinated persons experiencing primary infections.
 3. Reanalysis: Results
@@
 $I_{r}=0.456$, $H_{r}=0.0014$, $F_{pi}=0.69$, $E_{pi}=0.30$, $E_{v}=0.30$, $E_{h}=0.57$, $HRR_{pi}$$HRR_{v}=0.67$, $HRR_{h}=0.88$ 300 68 232 268 0.87
-* For discussions of “Assumptions,” see Section 3.3 as well as Supplement 1.
-We illustrate our model’s application by carrying out Scenario-A computations, Using Scenario A assumptions from Table 6 and our definitions (3.2.1) and (3.2.2) of $U$ and $V$, we obtain
-$U=\left(1-0.62\right)\cdot\left(1{\rm m}\right)\cdot 1\cdot 0.0014+0.62\cdot\left(1{\rm m}\right)\cdot 1\cdot\left(1-0.47\right)\cdot 0.0014\cdot\left(1-0.79\right)\approx 629,{\rm and}\\$ $V=\left(1-0.62\right)\cdot\left(1{\rm m}\right)\cdot 1\cdot\left(1-0.30\right)\cdot 0.0014\cdot\left(1-0.67\right)+0.62\cdot 1\cdot\left(1-0.57\right)\cdot 0.0014\cdot\left(1-0.88\right)\approx 168.$
-Thus, for Scenario A, the hospitalizations-prevented benefit of vaccination is projected to be UV= 629168=461U-V=\ 629-168=461 per million full mRNA-1273 vaccinations, roughly one-tenth of the FDA’s Scenario-1 estimate 4,766 [11].
-Although Scenario A has a benefit-risk ratio above 1, its assumption that $I_{r}=1$ makes this scenario highly unlikely.
-[^9] Based on seroprevalence and modeling information, we estimate $I_{r}\leq 0.5$ [Supplement 2, Section 2.3].
-For the more likely scenarios B, D, and E, our modeling suggests that mRNA-1273 vaccination of 18–25-year-old males generated between 16% and 63% more hospitalizations from VAM/P alone compared to COVID hospitalizations prevented.
-Comparing the projections of FDA’s “most likely scenario” to ours, we find that the FDA’s benefit-risk ratio, 43.33, is over 60 times more favorable than ours, 0.67.
-Though Scenario D is our most likely scenario, Scenario E is equally important, being a “best-case scenario” for vaccination incorporating the minimum reasonable assumption of prior-infection protection vs.
+* For discussions of “Assumptions,” see Section 2.3 as well as Supplement S1.
+We illustrate our model’s application by carrying out Scenario-A computations. Using Scenario A assumptions from Table 5 and our definitions (1) and (2) of $U$ and $V$, we obtain
+$U=\left(1-0.62\right)\cdot\left(1{\rm m}\right)\cdot 1\cdot 0.0014+0.62\cdot\left(1{\rm m}\right)\cdot 1\cdot\left(1-0.45\right)\cdot 0.0014\cdot\left(1-0.79\right)\approx 632,{\rm and}\\$ $V=\left(1-0.62\right)\cdot\left(1{\rm m}\right)\cdot 1\cdot\left(1-0.30\right)\cdot 0.0014\cdot\left(1-0.67\right)+0.62\cdot 1\cdot\left(1-0.57\right)\cdot 0.0014\cdot\left(1-0.88\right)\approx 168.$
+Thus, for Scenario A, the hospitalizations-prevented benefit of vaccination is projected to be UV= 632168=464U-V=\ 632-168=464 per million full mRNA-1273 vaccinations, roughly one-tenth of the FDA’s Scenario-1 estimate 4766 [13].
+Although Scenario A has a benefit–risk ratio above 1, its assumption that $I_{r}=1$ makes this scenario highly unlikely.
+[^12] Based on seroprevalence and modeling information, we estimate $I_{r}\leq 0.5$ [Supplement S2, Section S2.3].
+For the more likely scenarios B, D, and E, our modeling suggests that mRNA-1273 vaccination of 18–25-year-old males generated between 16% and 62% more hospitalizations for VAM/P alone compared to COVID hospitalizations prevented.
+To be conservative, we adjust these estimates downward.
+As discussed in Section 1.4, for our main analysis (whose results are presented in Table 5 above), we have included 1st dose VAM/P risk without accounting for 1st dose benefits (which is the choice the FDA made in its BNT162b2 assessment [12] and its BLA-Memo assessment of mRNA-1273 [51]).
+This may be viewed as a conservative choice, being consistent with the principle that vaccines should have “highly favorable benefit risk profiles” [1].
+Nevertheless, as we indicated in Section 1.4, including 1st dose VAM/P risk while not including 1st dose benefits potentially biases an assessment against vaccination.
+Thus, we adjust our results to remove this bias by comparing the hospitalizations-prevented benefit of two doses to the VAM/P risk attributable to dose two only.
+Although this adjustment potentially biases our outcomes analysis in favor of vaccination, we, conservatively, characterize the outcomes of our more likely scenarios in the following way: comparing our estimated 2nd-dose VAM/P risk at 250 hospitalizations per million to the hospitalizations-prevented benefit of two doses—165 per million in Scenarios B, 181 per million in Scenario D, and 232 per million in Scenario E—we conclude that mRNA-1273 vaccination of 18–25-year-old males generated between $8\%$ and $52\%$ more hospitalizations for VAM/P compared to COVID hospitalizations prevented, with an increase of $38\%$ for Scenario D.
+In particular, relative to hospitalizations, risks exceed benefits in both Scenario D, our most likely scenario, and in Scenario E, a “best-case scenario” for vaccination incorporating the minimum reasonable assumption of prior-infection protection vs.
 vaccine protection.
-In Section 2 of Supplement 2, we carry out sensitivity analyses exploring how modifying model inputs impacts the benefit-risk picture.
-For example, if we fix our model-input values, except for $I_{r}$, at their Scenario-D levels, then benefit-risk equipoise is reached when $I_{r}\approx 0.685$.
-We also illustrate how slider-plots can provide insights concerning equipoise questions involving variation of more than one model input.
+An ideal benefit–risk assessment of a two-dose vaccination series would account for the benefits and risks of both doses.
+We present such an assessment in Section S2 of Supplement S2 and provide some commentary on it in Section 3.3 below, where we discuss sensitivity analyses exploring how modifying model inputs impacts the benefit–risk picture.
 3.1. Subpopulation Analyses Based on Prior-Infection Status
-There is a compelling case for conducting a benefit-risk assessment for the subpopulation of unvaccinated 18–25-year-old males having prior-infection protection because prior infection not only decreases the hospitalizations-prevented benefit of vaccination but also increases the risk of severe adverse events101010e.g., those described in [57, 58, 59, 60, 61, 62, 63] after vaccination [64, 65], with [64] reporting prior infection is “associated with an increased risk of severe side effects leading to hospital care (1.56 [risk ratio] (95%CI: 1.14–2.12)).” Moreover, prior infection may increase the risk of VAM/P [42]—see Appendix 4 of Supplement 2 for details.
-[^11] Survey results reported in [66] also suggest that prior infection may increase the risk of VAM/P; e.g., in Table 4, we see that 2.53% of those having a confirmed SARS-CoV-2 infection before dose 1 of mRNA-1273 reported “chest tightness/pain” during the 7-day period after the dose while only 0.54% of those not having a confirmed infection before dose 1 reported these symptoms; corresponding percentages for dose 2 (Table 5) are 1.78% and 0.98%.
-For the subpopulation of 18–25-year-old unvaccinated males having a COVID-infection prior to the evaluation period 1/1/22 – 5/31/22, we can use our model to compute the expected number of COVID hospitalizations in various scenarios simply by setting $F_{pi}=1$ while leaving all other inputs unchanged.
-For example, with $F_{pi}=1$ in Scenario A, our model predicts $U\approx 155.82$ hospitalizations per million and $V\approx 72.24$, yielding a hospitalizations-prevented benefit of 83.58 per million vaccinated, far below VAM/P hospitalization risk ($\approx$ 268 per million vaccinated).
-As we discuss in Appendix 3 of Supplement 1, the preceding paragraph’s analysis may be reinterpreted in terms of evaluation-period hospitalization risk for a typical unvaccinated 18–25-year-old male, COVID-infected prior to the evaluation period.
+There is a compelling case for conducting a benefit–risk assessment for the subpopulation of unvaccinated 18–25-year-old males having prior-infection protection because prior infection not only decreases the hospitalizations-prevented benefit of vaccination but also increases the risk of severe adverse events (e.g., those described in [60, 61, 62, 63, 64, 65, 66]) after vaccination [67, 68], with [67] reporting prior infection is “associated with an increased risk of severe side effects leading to hospital care (1.56 [risk ratio] (95%CI: 1.14–2.12)).” Moreover, prior infection may increase the risk of VAM/P [15]—see Appendix S4 of Supplement S2 for details.
+[^13] Survey results in [69] also suggest that prior infection may increase VAM/P risk; e.g., 2.53% of those having a confirmed SARS-CoV-2 infection before dose 1 of mRNA-1273 reported “chest tightness/pain” during the 7-day postdose period [69, Table 4] while only 0.54% of those not having a confirmed infection before dose 1 reported these symptoms; corresponding percentages for dose 2 are 1.78% and 0.98% [69, Table 5].
+For the subpopulation of 18–25-year-old unvaccinated males having a COVID infection prior to the evaluation period 1 January 2022–31 May 2022, we can use our model to compute the expected number of COVID hospitalizations in various scenarios simply by setting $F_{pi}=1$ while leaving all other inputs unchanged.
+For example, with $F_{pi}=1$ in Scenario A, our model predicts $U\approx 161.7$ hospitalizations per million and $V\approx 72.24$, yielding a hospitalizations-prevented benefit of 89.46 per million vaccinated, far below VAM/P-hospitalization risk ($\approx$268 per million vaccinated).
+As we discuss in Appendix S3 of Supplement S1, the preceding paragraph’s analysis may be reinterpreted in terms of evaluation-period hospitalization risk for a typical unvaccinated 18–25-year-old male COVID-infected prior to the evaluation period.
 For such a male, the total hospitalization risk of the choice to vaccinate is essentially the sum of his probability of VAM/P hospitalization and his probability of having a breakthrough-case COVID hospitalization during the evaluation period.
-Our analysis is summarized in Tables 7 and 8 below in terms of hospitalization-risk ratios, with ratios <1<1 indicating the choice to vaccinate decreases risk and >1>1 indicating it increases risk.
+Our analysis is summarized in Tables 6 and 7 below in terms of hospitalization-risk ratios, with ratios <<1 indicating the choice to vaccinate decreases risk and >>1 indicating it increases risk.
 Table 6.
@@
 VAM/P Risk 268 in 1 million
-11 155.82 in 1 million 72.24 in 1 million $2.18\approx\frac{268+72.24}{155.82}$
-$0.5$ 77.91 in 1 million 36.12 in 1 million $3.90\approx\frac{268+36.12}{77.91}$
-$0.25$ 38.96 in 1 million 18.06 in 1 million $7.34\approx\frac{268+18.06}{38.955}$
+11 161.70 in 1 million 72.24 in 1 million $2.10\approx\frac{268+72.24}{161.70}$
+$0.5$ 80.85 in 1 million 36.12 in 1 million $3.76\approx\frac{268+36.12}{80.85}$
+$0.25$ 40.43 in 1 million 18.06 in 1 million $7.08\approx\frac{268+18.06}{40.43}$
 *COVID hospitalization risk is considered over the 5-month evaluation period
-$I_{r}(1-E_{pi})$ is the probability the “typical male” (18–25 with prior-infection protection) will become COVID infected over the evaluation period if unvaccinated; e.g., 0.53 if $I_{r}=1$
-According to Table 7, our modeling suggests that for a typical unvaccinated 18–25-year-old male with prior-infection, the choice to vaccinate created a hospitalization risk over twice that associated with the choice not to vaccinate even if $I_{r}=1$ (evaluation-period infection is certain for an infection-naïve male 18–25).
+$I_{r}(1-E_{pi})$ is the probability the “typical male” (18–25 with prior-infection protection) will become COVID infected over the evaluation period if unvaccinated; e.g., 0.55 if $I_{r}=1$
+According to Table 6, our modeling suggests that for a typical unvaccinated 18–25-year-old male with prior infection, the choice to vaccinate created a hospitalization risk over twice that associated with the choice not to vaccinate even if $I_{r}=1$ (evaluation-period infection is certain for an infection-naïve male 18–25).
 Because a “typical male” has an “average comorbidities level,” the information in Table 6 is difficult to apply.
-However, if our typical unvaccinated male of Table 7 also has no comorbidities, then his hospitalization-risk ratios should exceed those of Table 7, [^12] See, e.g., the BMI 23–24 column of Table 9 because risk-ratio denominators should decrease given no comorbidities and there is no evidence that being free of comorbidities reduces VAM/P risk [67].
-[^13] See also [68, p.
+However, if our typical unvaccinated male of Table 6 also has no comorbidities, then his hospitalization risk ratios should exceed those of Table 6, [^14] Compare, e.g., the BMI 23–24 column of Table 8 to the final column in Table 6.
+because risk-ratio denominators should decrease given no comorbidities and there is no evidence that being free of comorbidities reduces VAM/P risk [70].
+[^15] See also [71, p.
 383].
@@
 Table 7.
-Hospitalization-risk ratios for a typical male 18—25 without prior-infection protection: risk of hospitalization after choosing mRNA-1273 vaccination (VAM/P or severe COVID-breakthrough infection) divided by risk of COVID hospitalization after choosing not to vaccinate.
+Hospitalization-risk ratios for a typical male 18—25 without prior-infection protection: risk of hospitalization after choosing mRNA-1273 vaccination (VAM/P or severe COVID-breakthrough infection*) divided by risk of COVID hospitalization after choosing not to vaccinate.
 Probability
@@
 $0.25$ 350 in 1 million 80.85 in 1 million $1.00\approx\frac{268+80.85}{350}$
-4.2. Subpopulation Analyses Based on Prior-Infection Status and BMI
+*COVID hospitalization risk is considered over the 5-month evaluation period.
+3.2. Subpopulation Analyses Based on Prior Infection Status and BMI
 At this point, we discontinue confining to footnotes discussions and citations of data unavailable at the time of the FDA’s benefit-risk analysis of mRNA-1273.
-The FDA acknowledges, “The health condition of individuals may remarkably impact the B-R profile and its evaluation” [11, Section 4].
+The FDA acknowledges, “The health condition of individuals may remarkably impact the B-R profile and its evaluation” [13, Section 4].
 Comorbidities such as obesity may triple COVID-hospitalization risk [72].
-Our model can produce a benefit-risk analysis based on comorbidity status by appropriately adjusting its infection-hospitalization-rate input $H_{r}$.
-We illustrate the process in Appendix 3 of Supplement 1, focusing on obesity because evidence suggests that Body Mass Index (BMI) is the single most potent modifier of COVID-19 risk for those under 60 years old [70, 71].
+Our model can produce a benefit–risk analysis based on comorbidity status by appropriately adjusting its infection-hospitalization-rate input $H_{r}$.
+We illustrate the process in Appendix S3 of Supplement S1, focusing on obesity because evidence suggests that Body Mass Index (BMI) is the single most potent modifier of COVID-19 risk for those under 60 years old [73, 74].
 We use an assessment of COVID-hospitalization risk based on BMI that finds risk is minimized when BMI $\approx 23$ and that for $23\leq\text{BMI}\leq 44$ risk is multiplied by a factor of approximately 1.09 (1.08, 1.10) for each unit increase in BMI [75].
@@
 BMI 23–24 BMI 25–29 BMI 30–34 BMI 35–39 BMI 40–44
-11 2.66, 0.39 1.86, 0.30 1.29 , 0.23 0.92, 0.19 0.68, 0.17
-$0.5$ 5.08, 0.66 3.49, 0.48 2.35, 0.35 1.61, 0.27 1.13, 0.22
-$0.25$ 9.92 ,1.20 6.74 , 0.84 4.47 , 0.59 2.99, 0.42 2.03, 0.32
+11 2.56, 0.39 1.79, 0.30 1.25 , 0.23 0.89, 0.19 0.66, 0.17
+$0.5$ 4.90, 0.66 3.36, 0.48 2.26, 0.35 1.55, 0.27 1.09, 0.22
+$0.25$ 9.56 ,1.20 6.49 , 0.84 4.30 , 0.59 2.85, 0.42 1.85, 0.32
 *With prior-infection protection, $I_{r}(1-E_{pi})$ is the probability of infection over the evaluation period.
-Consistent with Tables 7 and 8, Table 9 indicates that for those having BMI in the intermediate range 25–34, prior-infection status determines whether vaccination is beneficial.
-However, for the those with high BMI (35–44) in a high-infection-risk scenario Table 9 indicates that vaccination becomes beneficial even for those with prior-infection protection.
-4.3. Validating the Model
-In Supplement 2, we use Ontario COVID-hospitalizations data from the evaluation period 1/1/22– 5/31/22 to show that our model’s projections of Table 6 for the hospitalizations-prevented benefit of vaccination for 18–25-year-old males are much more accurate than those of the FDA [11].
-[^14] We use Ontario data for many reasons; e.g., unlike the CDC, Public Health Ontario chose to publish incidental COVID-hospitalization rates (starting 1/8/2022) [30]; see [Supplement 2, Section 3] for other reasons.
-We test our model in two ways: (i) showing that it fits Ontario COVID-hospitalization data with inputs at plausible levels for Ontario over the evaluation period [Supplement 2 Section 8], and (ii) using Ontario COVID-hospitalization data to predict corresponding data for the U.S.
-[Supplement 2, Section 7].
-For (i), we fit our model to Ontario data for the general-population mix of 18–25-year-old male-and-female Ontarians by using Mathematica to solve a sequence of least-squares problems depending on the value of $H_{r}$.
-For $H_{r}=0.0011$ and $0.0010$, we obtain model inputs in plausible ranges for Ontario that yield a perfect modeling fit and note that these values of $H_{r}$ are consistent with COVID hospitalization data from the state of Connecticut.
-A brief discussion of (ii) is provided in the next paragraph.
-Based on COVID hospitalization counts from Public Health Ontario as well as corresponding incidental hospitalization rates, we estimate that vaccinating one million 18–25-year-old Ontarians prevented 152 COVID hospitalizations over the evaluation period—see Table S2.5.
-Making various adjustments e.g., accounting for the better health of Ontarians vs.
-Americans, we obtain an upper-bound estimate that 214 evaluation-period hospitalizations were prevented by one million full COVID vaccinations of 18–25–year-old males in the U.S.—see [Supplement 2, Table S2.7] and the discussion that follows it.
-Observe that 214 is an upper bound for our projections of Scenarios B–E (see Table 6) but not an upper bound for any of the FDA’s projections.
-Ontario data also supports the definitiveness of the results of Table 7 above given the male has no comorbidities, suggesting that 151 in 1 million is an overestimate of his evaluation-period COVID hospitalization risk [Supplement 2, Section 2.3].
+Tables 6 and 7 of the preceding section establish the expectation that for a “typical male” 18–25 (having an “average comorbidities level”), prior-infection status determines whether vaccination is beneficial.
+Consistent with this expectation, Table 8 above indicates that for those having BMI in the intermediate range 25–34, prior-infection status does determine whether vaccination is beneficial.
+However, for those with high BMI (35–44) in a high-infection-risk scenario, Table 8 indicates that vaccination becomes beneficial even for those with prior-infection protection.
+3.3. Sensitivity Analyses
+Perhaps the most important of our sensitivity analyses assesses how the inclusion of 1st-dose benefits in our Scenarios B–E changes the corresponding benefit–risk ratios.
+This analysis, carried out in Section S2 of Supplement S2, yields the following benefit–risk ratios for our more likely scenarios: 0.68 for B, 0.73 for D, and 0.96 for E.
+Estimating inputs for 1st-dose benefit modeling was especially challenging given the extremely limited data on 1st-dose benefits for Omicron available before 22 January 2022.
+Our having lower confidence in input estimates for 1st-dose modeling contributed to our decision to present the modeling only as a sensitivity analysis.
+Also, in Section S2 of Supplement S2, we provide a systematic sensitivity analysis related to Scenario D, our most likely scenario.
+For each input, we hold all others constant at their Scenario D values, determining the extent of possible variation in the chosen input before equipoise is reached.
+For example, if we fix our model input values, except for $I_{r}$, at their Scenario D levels, then benefit–risk equipoise is reached when $I_{r}\approx 0.678$, a value nearly 50% greater than the Scenario D value of $0.456$.
+We also point out that the risk-exceeds-benefit conclusion of Scenario D is not sensitive to the value of $E_{pi}$: even if $E_{pi}=0$, the benefit–risk ratio of this scenario is $\approx 0.83$.
+The insensitivity of the risk-exceeds-benefit outcome of Scenario D to significant changes in inputs from their Scenario D values increases confidence in the outcome.
+3.4. Model Validation
+In Supplement S2, we test our model using Ontario COVID-hospitalization data from the evaluation period of 1 January 2022–31 May 2022, showing its projections of the hospitalizations-prevented benefit of vaccination are more accurate than those of the FDA [13].
+We use Ontario data for many reasons; e.g., unlike the CDC, Public Health Ontario chose to publish incidental COVID-hospitalization rates (starting 8 January 2022) [36]; see [Supplement S2, Section S3] for other reasons.
+We test our model in two ways: (i) showing that it fits Ontario COVID-hospitalization data with inputs at plausible levels for Ontario over the evaluation period [Supplement S2, Section S8], and (ii) using Ontario COVID-hospitalization data to predict corresponding data for the U.S.
+[Supplement S2, Section S7].
+Based on COVID-hospitalization counts from Public Health Ontario as well as corresponding incidental-hospitalization rates, we estimate that vaccinating one million 18–25-year-old Ontarians prevented 152 COVID hospitalizations over the evaluation period—see Table S2.5 of Supplement S2.
+Making various adjustments, e.g., accounting for the better health of Ontarians vs.
+Americans, we obtain an upper-bound estimate that 214 evaluation-period hospitalizations were prevented by one million full COVID vaccinations of 18–25–year-old males in the U.S.—see Table S2.7 of Supplement S2 and the discussion that follows it.
+Observe that 214 is an upper bound for our projections of Scenarios B–D (see Table 5) and a little below our projection 232 of Scenario E, but all of the FDA’s projections, ranging from 635 to 4766 for Omicron-based scenarios, substantially exceed 214.
+Ontario data also supports the definitiveness of the results of Table 6 above given the male has no comorbidities, suggesting that 151 in 1 million is an overestimate of his evaluation-period COVID hospitalization risk [Supplement S2, Section S2.3].
 4. Discussion
 By accounting for prior-infection protection and using evidence-based model inputs, we find in our Scenarios B–E of Table 5 that risks of mRNA-1273 vaccination outweighed benefits for the general population of 18–25-year-old males, assuming hospitalizations are the principal endpoint of analysis.
-In particular, for our scenarios predicting plausible infection-level increases over the evaluation period (Scenarios B, D, and E), our model suggests that vaccination of 18–25-year-old males generated between 16% and 63% more hospitalizations from vaccine-attributable myo/pericarditis (VAM/P) alone compared to COVID hospitalizations prevented.
-[^15] The range 16%–63% is consistent with our model-validation work discussed in Section 4.3: recall we estimate that at most 214 evaluation-period hospitalizations were prevented by one million mRNA-1273 vaccinations of 18–25-year-old males in the U.S.
-[Supplement 2, Table S2.7], and note that our corresponding VAM/P hospitalization projection of 268 is a little over 25% higher than 214.
-Recalling that the FDA found, “[T]he benefits of vaccination with the Moderna COVID-19 vaccine clearly outweigh the risks, even among the male adolescent population” [11, Section 4], we have demonstrated that the conclusions of a benefit-risk assessment for vaccination may be dramatically impacted by accounting for the benefits conferred by prior infection.
-We have also demonstrated in Tables 79 how the benefit-risk picture for vaccination may vary based on an individual’s prior-infection and comorbidity status.
-Consistent with our model’s finding that for 18–25-year-old males in the U.S.
-the risks of mRNA-1273 vaccination outweighed the benefits relative to hospitalizations, we note that France suspended mRNA-1273 use on 10/15/2021, making it available again on 11/8/21 but at half-dose and only for those over 30 years [73, p.
+In particular, for our scenarios predicting plausible infection-level increases over the evaluation period (Scenarios B, D, and E), our model suggests that vaccination of 18–25-year-old males generated between 16% and 62% more hospitalizations for vaccine-attributable myo/pericarditis (VAM/P) alone compared to COVID hospitalizations prevented, or, discounting 1st-dose VAM/P risk, between 8% and 52% more hospitalizations.
+Recalling that the FDA found, “[T]he benefits of vaccination with the Moderna COVID-19 vaccine clearly outweigh the risks, even among the male adolescent population” [13, Section 4], we have demonstrated that the conclusions of a benefit–risk assessment for vaccination may be dramatically impacted by accounting for the benefits conferred by prior infection.
+We have also demonstrated in Tables 68 how the benefit–risk picture for vaccination may vary based on an individual’s prior-infection and comorbidity status.
+Finally, we have demonstrated how a benefit–risk assessment for a two-dose vaccination series may acknowledge potential biases relating to the inclusion or exclusion of 1st dose benefits or risks.
+Consistent with our model’s finding that for 18–25-year-old males in the U.S., risks of mRNA-1273 vaccination outweighed benefits relative to hospitalizations, we note that France suspended mRNA-1273 use on 15 October 2021, making it available again on 8 November 2021 but at half-dose and only for those over 30 years [76, p.
 4].
+Also observe our model’s finding that vaccination of 18–25-year-old males generated between 8% and 52% more hospitalizations for vaccine-attributable myo/pericarditis (VAM/P) compared to COVID hospitalizations prevented is consistent with our model-validation work discussed in Section 3.4: recall that we estimate that at most 214 evaluation-period hospitalizations were prevented by one million mRNA-1273 vaccinations of 18–25-year-old males in the U.S.
+[Supplement 2, Table S2.7], and note that our corresponding VAM/P-hospitalization projection of 268 is a little over 25% higher than 214.
 We emphasize that the cardiac risk of COVID infection is incorporated into our modeling because our estimates of COVID hospitalizations include those for COVID-related cardiac issues.
@@
 Contrary to the editorial’s assertion, Patone et al.
-find that “In men younger than 40 years old, the number of excess myocarditis events [leading to hospitalization] per million people was higher after a second dose of mRNA-1273 than after a positive SARS-CoV-2 test (97 [95% CI, 91–99] versus 16 [95% CI, 12–18])” [6].
+find that, “In men younger than 40 years old, the number of excess myocarditis events [leading to hospitalization] per million people was higher after a 2nd dose of mRNA-1273 than after a positive SARS-CoV-2 test (97 [95% CI, 91–99] versus 16 [95% CI, 12–18])” [6].
 Karlstad et al.
@@
 13.7 per million positive tests [5, eTable 7] among males 16–24.
-Moreover, these myocarditis incidence rates expressed in cases per million positive tests overstate the risk of myocarditis after COVID infection expressed in cases per million COVID infections (because many infections are not confirmed by positive tests, but essentially all infections among  those seeking care for myocarditis are confirmed by positive tests) [75].
-Consistent with [75], the authors of [76] note, “[T]he attributable risk estimates for COVID-19 [associated myocarditis] used laboratory confirmed cases as the denominator and will be affected by the proportion of all SARS-CoV-2 infections captured by testing, precluding a direct comparison with vaccine-associated attributable risks.”
-The strength of prior-infection protection
-As noted in Section 2.2, the FDA’s assumption that prior infection would provide no protection against Omicron is contrary to immunological expectations and studies (e.g., [15, 16, 17, 19, 26, 27]) available at the time of the FDA’s mRNA-1273 analysis suggesting that prior-infection protection would likely equal or exceed vaccine protection among the COVID-naïve.
-As noted in Table 5 footnotes, we have assumed that vaccination’s effectiveness against hospitalization is 77% while that provided by prior infection is 88.9%.
-A systematic review [24] involving 21 vaccine-effectiveness studies from the period 12/3/2021 – 4/7/2022 reports,
+Moreover, these myocarditis-incidence rates expressed in cases per million positive tests overstate the risk of myocarditis after COVID infection expressed in cases per million COVID infections (because many infections are not confirmed by positive tests, but essentially all infections among  those seeking care for myocarditis are confirmed by positive tests) [78].
+Consistent with [78], the authors of [79] note, “[T]he attributable risk estimates for COVID-19 [associated myocarditis] used laboratory confirmed cases as the denominator and will be affected by the proportion of all SARS-CoV-2 infections captured by testing, precluding a direct comparison with vaccine-associated attributable risks”.
+4.1. The Strength of Prior-Infection Protection
+As noted in Section 1.3, at the time of the FDA’s mRNA-1273 assessment, immunological expectations and studies such as [21, 22, 23, 25, 32, 33] suggested that prior-infection protection would likely equal or exceed vaccine protection among the COVID-naïve.
+In our modeling, we have assumed that vaccination’s effectiveness against hospitalization is 77% in all scenarios, while that of prior infection is 88.5% except in Scenario E.
+A systematic review [30] involving 21 vaccine-effectiveness studies from the period 3 December 2021–7 April 2022 reports,
 In contrast to vaccine effectiveness against delta severe disease, the majority of vaccine effectiveness estimates for omicron severe disease were below 75%; for example, thirteen (81%) of sixteen vaccine effectiveness estimates within three months of vaccination with the primary series were below 75% (Figure).
 Moreover, 13 (42%) vaccine-specific estimates fell below 50% at some point in time after vaccination.
-In contrast, the meta-analysis [21] estimates prior-infection’s effectiveness against severe Omicron to be over 80% through 6 months after infection, while the meta-analysis [22, Table S2] estimates its effectiveness to be over 88% through 45 weeks after infection, with the former estimate based on studies from 1/1/22–6/1/22 and the latter, on studies published by 9/31/22.
-In mid-January of 2022, the minimum reasonable assumption the FDA might have made is that prior-infection protection would equal vaccine protection (as in our Scenario E); instead, in all Omicron-based scenarios of [11], the agency ascribed to vaccination 72% effective protection against hospitalization while assuming prior infection offers no protection.
-5.1. Recommendations
-Our modeling provides a call to action to policy-forming health agencies to employ multi-variable models like ours that can better stratify vaccination recommendations.
-To protect vaccine acceptance and promote public trust, the FDA must adopt a rigorous framework for benefit-risk assessment of new vaccines and revisit recommendations as new data becomes available.
-In particular, the framework must incorporate consideration of the protection that prior infection affords against development of severe disease.
-We also recommend United States public-health agencies reassess their priorities.
-Two years after the first cases of COVID were identified in the U.S., the FDA decided to “not investigate the benefits and risks [of mRNA-1273 vaccination] for specific subpopulations, such as those with comorbidities, due to limited information for these populations” [11, Section 4].
-Also, public-health agencies had reason to prioritize understanding the effectiveness of prior-infection protection and had evidence that it was likely to be high.
-For example, in Pfizer’s phase-3 trial for BNT162b2, 567 participants in the placebo group and 526 in the treatment arm had evidence of prior COVID-19 infection.
-In each of these cohorts, there was only 1 reinfection (or <0.2<0.2% for both) [77, p.
-27].
+In contrast, the meta-analysis [27] estimates prior infection’s effectiveness against severe Omicron to be over 80% through 6 months after infection, while the meta-analysis [28, Table S2] estimates its effectiveness to be over 88% through 45 weeks after infection, with the former estimate based on studies from 1 January 2022–1 June 2022 and the latter, on studies published by 31 September 2022.
+In mid-January of 2022, the minimum reasonable evidence-based assumption would have been that prior-infection protection would equal vaccine protection (as in our Scenario E).
+4.2. Assessing VAM/P Risk for mRNA-1273
+Contributing to our finding mRNA-1273 vaccination’s risks exceeded benefits for 18–25-year-old males relative to hospitalizations is our assessment that VAM/P risk for mRNA-1273 is substantially greater than that suggested by the FDA’s BEST System.
+As detailed in Sections 1.4 and 2.3, our mRNA-1273 VAM/P-risk estimate of 268 hospitalizations per million full vaccinations is based on data, available to the FDA at the time of its mRNA-1273 assessment, from Ontario [14], England [15], and the United States [16, 12, 44, 51].
+In Appendix S2 of Supplement S1, we pool data, unavailable at the time of the FDA’s mRNA-1273 assessment, from Scandinavia [5], France [7], and British Columbia [8] to obtain a VAM/P-hospitalization rate between 240 and 277 per million full, homologous mRNA-1273 vaccinations among males 18–25.
+A possible explanation for the FDA’s underestimating VAM/P risk for mRNA-1273 is that its BEST System has “inherent limitations, such as small sample sizes and imperfect sensitivity of ICD-10 codes to identify these rare outcomes” [12, Section 4].
+Recall that there is evidence that the CDC’s VSD system also underascertained VAM/P risk for mRNA-1273.
+As noted in Section 1.5, during the period 1 June 2021–4 September 2021 of “enhanced passive surveillance” among males and females 18–39, Public Health Ontario’s myo/pericarditis detection rate per million 2nd doses of mRNA-1273 was 3.76 times that of the CDC’s VSD system based on its data through 9 October 2021 [Supplement S2, Appendix S3].
+The study [4] offers at least a partial explanation for the VSD’s lower detection rate by revealing shortcomings in the VSD’s search algorithm.
 4.3. Limitations
-Our principal aim has been to illustrate how prior-infection protection, as well as comorbidity status, can be integrated into benefit-risk modeling of a two-dose vaccination series.
-In applying our model to extend and improve the FDA’s mRNA-1273 assessment [11] relative to its hospitalizations endpoint, we have assumed that young males are hospitalized for COVID at a lower rate than young females, as did the FDA in [10] and [11].
-Data from [10] and [54] suggests this rate difference is substantial; however, the rates of [10] and [54] include incidental hospitalizations and incidental rates for males may differ from those for females.
+Our principal aim has been to illustrate how prior-infection protection, as well as comorbidity status, can be integrated into a deterministic benefit–risk model that extends, in a natural way, the one the FDA employed in its assessment of mRNA-1273 vaccination.
+We have developed and presented our model in accordance with BRIVAC (Benefit–RIsk models applied to VACcines) guidelines, in particular, its item 17 “Sensitivity/scenario analyses” [80].
+We acknowledge that it is desirable to employ more sophisticated approaches to benefit–risk modeling and to accounting for uncertainty in such modeling; for instance, Monte Carlo simulation modeling with Beta/Poisson distributions for model inputs, guiding input choices for simulations.
+In applying our model to extend and improve the FDA’s mRNA-1273 assessment [13] relative to its hospitalizations endpoint, we have assumed that young males are hospitalized for COVID at a lower rate than young females, as did the FDA in [12, 13].
+Data from [12, 57] suggests this rate difference is substantial; however, the rates of [12, 57] include incidental hospitalizations, and incidental rates for males may differ from those for females.
 We have assumed, as did the FDA, that vaccination and prior infection offer essentially equivalent protection for males and females.
@@
 However, the rate differences provided by Table 1 of [57], based on two years of data including that of the initial Omicron wave, are consistent with those of Table 2 of [12].
-COVID-hospitalizations data from Ontario [Supplement 2] and from Connecticut [Supplement 1, Appendix 1] suggests that our Omicron IHR value $0.0014$ for 18–25-year-old males is a reasonable estimate for the Omicron IHR for all 18–25 year-olds (general population mix of males and females).
+COVID-hospitalization data from Ontario [Supplement S2] and from Connecticut [Supplement S1, Appendix S1] suggests that our Omicron IHR value $0.0014$ for 18–25-year-old males is a reasonable estimate for the Omicron IHR for all 18–25-year-olds (general population mix of males and females).
 Thus, even if we assume males and females in this age range are hospitalized for COVID at roughly equivalent rates, our IHR value of 0.14% for 18–25-year-old males appears to be a good estimate.
-In Section 2 of Supplement 1, based on data available before 1/22/22, we obtain an Omicron IHR estimate for 18–25 year-olds (general-population mix of males and females) of 0.22%, not 0.14%.
+In Section S2 of Supplement S1, based on data available before 22 January 2022, we obtain an Omicron IHR estimate for 18–25 year-olds (general-population mix of males and females) of 0.22%, not 0.14%.
 One possible explanation is that we may have underestimated the incidental COVID-hospitalization rate for 18–25 year-olds.
-Our IHR estimate for 18–25 year-olds of 0.22% results from assuming an incidental Omicron-hospitalization rate of 40%, a rate suggested by [28] and [30] for hospitalizations with COVID among patients of all ages.
-However, as noted in Section 2.2 (and discussed in detail in Supplement 1), there is strong evidence that incidental COVID-hospitalization rates are higher among the young than among the elderly.
-In particular, COVID-hospitalization data discussed in Appendix 1 of Supplement 1 suggests an incidental rate of over 70% among 15–24 year-olds in Connecticut during calendar year 2020.
+Our IHR estimate for 18–25 year-olds of 0.22% results from assuming an incidental Omicron-hospitalization rate of 40%, a rate suggested by [34, 36] for hospitalizations with COVID among patients of all ages.
+However, as noted in Section 1.3 (and discussed in detail in Supplement S1), there is strong evidence that incidental COVID-hospitalization rates are higher among the young than among the elderly.
+In particular, COVID-hospitalization data discussed in Appendix S1 of Supplement S1 suggests an incidental rate of over 70% among 15–24-year-olds in Connecticut during calendar year 2020.
 Assuming an incidental rate of about 62.2% (instead of 40%) transforms our IHR estimate of 0.22% into the estimate 0.14%.
@@
 Our risk assessment, like the FDA’s, ignores known, rare vaccine-related adverse events other than myo/pericarditis that may require hospitalization.
-Finally, our risk assessment, like the FDA’s, ignores the possibility that COVID-19 vaccination may have long-term health effects; for instance, the FDA’s package-insert information for Moderna’s mRNA-1273 begins, “SPIKEVAX has not been evaluated for carcinogenic, mutagenic potential, or impairment of male fertility in animals” [78, Section 13].
+Finally, our risk assessment, like the FDA’s, ignores the possibility that COVID-19 vaccination may have long-term health effects; for instance, the FDA’s package-insert information for Moderna’s mRNA-1273 begins, “SPIKEVAX has not been evaluated for carcinogenic, mutagenic potential, or impairment of male fertility in animals” [81, Section 13].
+5. Conclusions
+Our model illustrates the importance of nuanced vaccine recommendations based on careful benefit–risk modeling that takes into account health status and prior infection(s), two covariates that were not included in the FDA’s models [12, 13].
+For healthy young males with prior infection, our model suggests that risks outweighed benefits (relative to hospitalizations) based on data available prior to licensure of the mRNA-1273 product.
+However, for higher-risk young males with obesity, and especially those with no prior infection, vaccination was found to confer benefit.
+Risks of infection and vaccination, such as long COVID and VAM/P, must be carefully weighed to offer informed consent.
+Current long-COVID prevalence estimates among adults with confirmed SARS-CoV-2 infection are most reliably estimated in the range of 7–15% when using standardized definitions and representative sampling [82, 83], and when using a conservative definition consistent with other post-viral syndromes, the WHO finds 2.8% among those younger than 20 [84].
+With respect to VAM/P, more recent formulations of Moderna’s COVID vaccine have a reduced mRNA dose and VAM/P-incidence rates have declined [85].
+Long-term VAM/P prognosis remains somewhat unclear.
+A CDC follow-up at 90 days found that although the majority (81%) of VAM/P cases recovered, 54% had ongoing abnormal cardiac findings and 26% were on daily cardiac medication; only 68% were cleared for all physical activity [86].
+Clinical guidelines for return-to-sport following myocarditis call for 3–6 months’ activity restriction [87, 88].
+In addition, there is the unknown prognosis for complications like late gadolinium enhancement (LGE) [89].
+The question of boosting the general population, including healthy young males, is beyond the scope of this paper and requires tremendous nuance, a fact widely recognized by fall 2021 [90].
+The FDA indicated in May 2025 [91] that it anticipates needing data from RCTs to evaluate clinical outcomes prior to authorizing COVID-19 vaccines for healthy individuals ages 6 months to 64 years, while immunogenicity standards will likely suffice for those with at least one chronic condition and those over age 65.
+The agency’s move thus operationalizes the risk-stratified approach supported by our deterministic modeling and will generate data that can be used in more sophisticated modeling.
 Acknowledgment: We thank Retsef Levi for providing comments that improved both the structure and content of this paper.
-Supplements: Supplement 1, Supplement 2